### v3

# Efficient and Transferable Agentic Knowledge Graph RAG via Reinforcement Learning

Abstract
Knowledge-graph retrieval-augmented generation (KG-RAG) couples large language models (LLMs) with structured, verifiable knowledge graphs (KGs) to reduce hallucinations and expose reasoning traces. However, many KG-RAG systems compose multiple LLM modules (e.g planning, reasoning, and responding), inflating inference cost and binding behavior to a specific target KG. To address this, we introduce KG-R1, an agentic KG retrieval-augmented generation (KG-RAG) framework through reinforcement learning (RL). KG-R1 utilizes a single agent that interacts with KGs as its environment, learning to retrieve at each step and incorporating the retrieved information into its reasoning and generation. The process is optimized through end-to-end RL. In controlled experiments across Knowledge-Graph Question Answering(KGQA) benchmarks, our method demonstrates both efficiency and transferability: Using Qwen-2.5-3B, KG-R1 improves answer accuracy with fewer generation tokens than prior multi-module workflow methods that use larger foundation or fine-tuned models. Furthermore, KG-R1 enables plug and play: after training, it maintains strong accuracy on new KGs without modification. These properties make KG-R1 a promising KG-RAG framework for real-world deployment. Our code is publicly available at  https://github.com/Jinyeop3110/KG-R1.
1 Introduction
22footnotetext: Work done during an internship at MIT-IBM Watson AI Lab, IBM Research.
Figure 1: Overview of KG-R1, a multi-turn agentic framework for KG-RAG trained with reinforcement learning. The framework enables cost-efficient inference and demonstrates strong cross-KG transferability.
Retrieval-augmented generation (RAG) has gained popularity as a way to enhance large language models (LLMs) with access to external knowledge, thereby reducing hallucinations and improving accuracy in knowledge-intensive tasks. Recent research has extended this idea to knowledge graph retrieval-augmented generation (KG-RAG), where knowledge graphs (KGs) are leveraged as the retrieval source. A KG is a structured representation of knowledge in the form of entities (nodes) and their relationships (edges) that encodes factual knowledge in a graph format. Augmenting LLMs with KGs has proven effective not only in mitigating the knowledge bottleneck, but also in improving reasoning performance over complex multi-hop relations and enhancing adaptation to continually evolving real-world information (Sun et al., 2024; Luo et al., 2024; Wang et al., 2025c). These properties make KG-RAG especially promising in high-stakes domains, such as medical consultation and legal analysis (Xiong et al., 2023; Cui et al., 2024).
As shown in Figure 1, typical KG-RAG adopts a modular workflow consisting of four primary subtasks: Retrieval to query facts from KGs, Reasoning to process the retrieved information, Reviewing to verify logical consistency, and Responding to synthesize the final answer (Baek et al., 2023; Sun et al., 2024; Luo et al., 2024; Wang et al., 2025c). Each distinct subtask is handled by specialized LLM modules through two main methods: (i) prompt-based with task-specific instructions, often including in-context demonstrations (Baek et al., 2023); and (ii) fine-tuned modules tailored to particular tasks (e.g., SPARQL generation (D’Abramo et al., 2025; Lee & Shin, 2024) or relation extraction (Yao et al., 2019)) on specific KGs.
Figure 2: Prior multi-module methods are costly and do not transfer well across KGs. Left: mean end-to-end generated tokens per query on WebQSP (Yih et al., 2016). Right: average $\mathrm{Hit@1}$ over five out-of-training KGQA datasets (See Sec. 4.2). KG-R1 achieves both low token cost and strong cross-KG transferability.
Despite improved reasoning accuracy, the real-world deployment of such workflows faces two key challenges, high computational cost and lack of generalization to new or updated KGs. First, prompt-based methods that repeatedly call large foundation LLMs accumulate inference across stages and drive up latency, token usage, and compute (e.g., ToG and ReKnoS; see left panel of Fig. 2). Second, these prompted or fine-tuned modules are typically tuned to a particular KG’s domain and schema (entity types, relations, constraints), often via curated in-context examples or KG-specific fine-tuning. As a result, performance does not transfer reliably when the domain shifts, the schema changes, or the system is deployed on a new KG (e.g., RoG; see right panel of Fig. 2).
To tackle these challenges, we introduce KG-R1, an agentic KG-RAG system that employs end-to-end multi-turn reinforcement learning (Jin et al., 2025; Zeng et al., 2025; DeepSeek-AI, 2025). As shown in Figure 1, the architecture of KG-R1 has two components: a single LLM agent and a KG retrieval server (as an environment). The KG retrieval server hosts the knowledge graph along with a set of retrieval actions. The LLM agent iteratively performs cycles of short reasoning followed by retrieval actions over multiple turns, with each decision informed by knowledge obtained from the KG retrieval server, and generates a final answer. Figure 2 demonstrates that KG-R1 achieves both high efficiency and strong cross-KG transferability simultaneously using a 3B-parameter model, outperforming prior methods. Our key contributions are summarized as follows:
1. KG-R1 framework. We introduce an agentic KG-RAG system (Section 3) that replaces multi-module pipelines with a single agent for KG-RAG, running against a lightweight KG server. The agent alternates between reasoning and retrieval actions over multiple turns, with the end-to-end trajectory optimized by RL using both turn-wise and outcome-based reward signals. Turn-wise rewards evaluate individual action effectiveness and adherence to formatting, while global rewards measure answer quality and retrieval relevance.
2. Efficient inference. By consolidating reasoning and retrieval into a single-agent, near-continuous workflow, KG-R1 achieves competitive reasoning accuracy with a small-parameter model while reducing token usage. This lowers latency and computational cost, making deployment feasible under tight budgets. Experiments demonstrate improved performance and efficiency compared to traditional multi-module workflows (Section 4.1).
3. Plug-and-play transferability across diverse KGs. KG-R1 easily transfers to diverse KGs and maintains strong KG-RAG performance (Section 4.2). The trained KG-R1 agent generalizes to new KGs without modification—backend KGs can be swapped without changing prompts or hyper-parameters, and without fine-tuning. This enables zero-shot transfer to unseen knowledge graphs.
2 Related Work
KG-RAG. Knowledge Graph Retrieval-Augmented Generation (KG-RAG) augments LLMs with structured knowledge graphs (KGs) to improve factuality and compositional reasoning. Early work grounds LLMs’ generation by translating natural-language questions into executable graph queries (e.g., SPARQL/Cypher), retrieving relevant subgraphs or answers, and feeding them back to the model (Ouyang et al., 2022; Izacard et al., 2023; Lee & Shin, 2024). More recent approaches adopt a modular LLM pipeline over KGs, interleaving natural-language reasoning with multi-stage planning, path search, and execution, where each stage uses a prompted or fine-tuned LLM (Luo et al., 2024; Sun et al., 2024; Wang et al., 2025c). Despite these advances, most systems rely on hand-engineered modules or prompt heuristics tied to a specific KG schema, which induce cost inefficiency and limit generalization. These challenges motivate our KG-R1 framework: a single-agent KG-RAG approach that improves efficiency and transfers well to new KGs.
Multi-turn RL for LLM. Reinforcement learning (RL) has become central to equipping LLMs with step-by-step (chain-of-thought) reasoning behavior (OpenAI et al., 2024; DeepSeek-AI, 2025). RL-enhanced models yield substantial gains in math and coding (Le et al., 2022; Chervonyi et al., 2025), and broader complex reasoning tasks. More recently, RL has been applied to agentic LLMs that invoke external tools (e.g., bash terminals and APIs) or interact with knowledge bases, improving tool use (Qin et al., 2024) and retrieval-augmented generation (RAG) (Jin et al., 2025; Wang et al., 2025a) to facilitate the tool use or RAG. Building on these advances, our work, KG-R1 adopts end-to-end RL as the core optimization algorithm for agentic KG-RAG framework.
3 KG-R1: An Agentic KG-RAG Framework
3.1 Problem Definition
KG–RAG spans many applications, including conversational assistants (Chaudhuri et al., 2021), recommendation systems (Wang et al., 2025b), and open-domain QA (Zhu et al., 2025). In this work, we instantiate and evaluate our approach on Knowledge Graph Question Answering (KGQA), which provides a grounded testbed for KG-RAG: ground-truth answers are tied to a fixed KG, evaluation is verifiable and intermediate graph reasoning is interpretable. We now formalize knowledge graphs and KGQA tasks.
Knowledge graphs. A knowledge graph (KG) is a graph-structured representation of real-world knowledge that encodes factual information as triples of entities and their relationships. We denote a KG as $G=\{\,\langle e,r,e^{\prime}\rangle\mid e,e^{\prime}\!\in\!\mathcal{E},\ r\!\in\!\mathcal{R}\,\},$ where $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and relations, respectively, and each triple $\langle e,r,e^{\prime}\rangle$ represents a directed edge from entity $e$ to entity $e^{\prime}$ via relation $r$. For example, there is an edge capital_of from entity Springfield to entity Illinois.
Knowledge Graph Question Answering (KGQA) is a reasoning task based on KGs. Consider a dataset $D=\{(q,G,A_{q})\}$, where $q$ is a natural language question, $G$ is a KG, and $A_{q}\subseteq\mathcal{E}$ is the ground-truth answer set paired with the question. For each gold answer $e^{*}\in A_{q}$, there exist one or more reasoning paths in $G$ that connect anchor entities mentioned in $q$ to $e^{*}$. A reasoning path is a sequence of labeled edges $r_{1},\dots,r_{\ell}\in\mathcal{R}$ instantiated by entities $e_{0},\dots,e_{\ell}$ such that $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell},$ where $e_{0}$ is an anchor entity mentioned in $q$ and $e_{\ell}=e^{*}\in A_{q}$, and $\mathcal{Z}(q,G,A_{q})$ denote the set of all valid reasoning paths for question $q$ over $G$ with respect to ground-truth answers $A_{q}$. In solving KGQA, given $q$ and $G$, a model attempts to discover a subset of valid reasoning paths from $\mathcal{Z}(q,G,A_{q})$ and predicts $\hat{A}_{q}$ based on the terminal entities of the discovered paths. The model’s performance is evaluated by comparing $\hat{A}_{q}$ with the ground-truth $A_{q}$. As an example, to answer the question “What is the capital of the U.S. state whose largest city is Chicago?”, a valid reasoning path is
$\text{Chicago}\xrightarrow{\texttt{located\_in\_state}}\text{Illinois}\xrightarrow{\texttt{capital}}\text{Springfield}$
This path belongs to $\mathcal{Z}(q,G,A_{q})$ and leads to the correct prediction $\hat{A}_{q}=\{\text{Springfield}\}$.
3.2 KG-R1 Framework
Figure 3: KG-R1 framework: a single LLM agent undergoes multi-turn generation–execution loop with a schema-agnostic KG retrieval server and responds with the final answer.
KG-R1 casts KG-RAG as a multi-turn interaction with a KG interface (KG retrieval server). We prioritize two design principles. First, we design a single-agent architecture that simplifies deployment and enables efficient, low-cost inference. Second, we build a schema-agnostic KG retrieval server that avoids KG-specific assumptions and remains portable across varied KGs. Given a KGQA dataset $D=\{(q,G,A_{q})\}$, we set up a KG retrieval server and a KG-R1 agent.
KG retrieval server. The server hosts the knowledge graph $G$ and provides a set of retrieval actions $a\in\mathcal{U}_{\text{ret}}$ that enable graph traversal through 1-hop operations:
[…]
$\mathtt{get\_tail\_relations}(e)$ $:=\{\,r\in\mathcal{R}\mid\exists e^{\prime}\in\mathcal{E}:\ (e,r,e^{\prime})\in G\,\},$
$\mathtt{get\_head\_relations}(e^{\prime})$ $:=\{\,r\in\mathcal{R}\mid\exists e\in\mathcal{E}:\ (e,r,e^{\prime})\in G\,\},$
$\mathtt{get\_tail\_entities}(e,r)$ $:=\{\,e^{\prime}\in\mathcal{E}\mid(e,r,e^{\prime})\in G\,\},$
$\mathtt{get\_head\_entities}(r,e^{\prime})$ $:=\{\,e\in\mathcal{E}\mid(e,r,e^{\prime})\in G\,\}.$
The retrieval action set $\mathcal{U}_{\text{ret}}$ is sufficient for realizing any reasoning path $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell}$ as a finite action sequence that arrives at $e_{\ell}\in A_{q}$ (i.e., the terminal entity is in the answer set). Forward traversal along $Z$ is implemented by $\mathtt{get\_tail\_entities}(e_{i-1},r_{i})$ for $i=1,\dots,\ell$ and backward traversal is implemented by $\mathtt{get\_head\_entities}(r_{i},e_{i})$. Notably, this choice of the retrieval action set guarantees completeness and schema-free transferability, as formalized by propositions 3.1 and 3.2, whose proofs are provided in Appendix A.1.
Proposition 3.1 (Retrieval Action Set Completeness).
For any reasoning path $Z:\ e_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{\ell}}e_{\ell}$ in $G$, there exists an action sequence in $\mathcal{U}_{\text{ret}}$ of length at most $\ell{+}1$ whose output includes $e_{\ell}$.
Proposition 3.2 (Schema-Free Transferability).
The semantics of $\mathcal{U}_{\text{ret}}$ depend only on directed triples $(e,r,e^{\prime})$, so any fixed client policy transfers across heterogeneous directed KGs by replacing $G$ without redefining $\mathcal{U}_{\text{ret}}$.
KG-R1 Agent. We model a single LLM agent that interacts with a KG retrieval server in a multi-turn setting. The agent undergoes initialization followed by a loop of generation and execution (see Figure 3). At initialization, a lightweight base LLM is configured with an instruction prompt $p$ (see the box below for the prompt template) containing general reasoning instructions, the user question $q$, and the KG retrieval server instructions (Table D.2). At each turn $t\leq H$, where $H$ is the maximum turn limit, the agent first undergoes the generation phase where it produces a response $m_{t}$ comprising two parts: (1) an internal reasoning wrapped in <think>...</think>, and (2) an action, which is either a KG retrieval action wrapped in <kg-query>...</kg-query>, or a final answer wrapped in <answer>...</answer>.
Prompt template for KG-R1 You are a helpful assistant. Answer the given question. You can query from knowledge base provided to you to answer the question. You can query knowledge up to [H] times. You must first conduct reasoning inside <<think>><</think>>. If you need to query knowledge, you can set a query statement between <<kg-query>><</kg-query>> to query from knowledge base after <<think>><</think>>. When you have the final answer, you can output the answer inside <<answer>><</answer>>. KG Query Server Instruction : [KG_query_server_instruction] Question: [question]. Assistant:
In the following execution phase, we parse $m_{t}$ into an action $a_{t}\in\mathcal{U}_{\text{ret}}\cup\{\operatorname{Answer}\}$ using an exact-match parser $\Psi$ (i.e., $a_{t}=\Psi(m_{t})$). If $a_{t}\in\mathcal{U}_{\text{ret}}$, executing it yields an observation $o_{t+1}$ (i.e., retrieved entities, relations, or an error message in case the action does not generate a valid retrieval), which is appended to the dialogue context prior to the next turn. If $a_{t}=\operatorname{Answer}$, the interaction terminates: the content inside <answer>...</answer> is extracted and post-processed (normalization, deduplication, and entity resolution) to produce the predicted answer set $\hat{A}_{q}$. Given interaction trajectory $\tau=(p,\,(m_{1},o_{1}),\ldots,(m_{t},o_{t}))$ with $t\leq H$, the KG-R1 agent’s policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ is defined over token sequences $m_{t}$ and governs how textual responses are generated.
3.3 KG-R1 training with RL
Our goal is to find the KG-R1 agent’s policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ that effectively retreive reasoning paths on $G$ via multi-turn interaction and generate the accurate answer $\hat{A}_{q}=A_{q}$. To this end, we optimize the base LLM with reinforcement learning, using a GRPO-style objective (DeepSeek-AI, 2025) in a multi-turn setting (Qian et al., 2025; Jin et al., 2025; Zeng et al., 2025). Our reward function combines action-level and outcome-based signals. The overview of the training procedure refers to Algorithm 1 (Appendix).
Reward Objective. To effectively train the KG-R1 agent, we combine verifiable turn rewards with outcome-level global rewards. Turn rewards ($r^{\text{turn}}_{t}$) provide local signals at each step as the sum of three components: (i) format validity $v_{\mathrm{fmt}}(m_{t})$ checks that $m_{t}$ contains both reasoning and a well-formed action $a_{t}\in\mathcal{A}$; (ii) KG query $v_{\mathrm{kg}}(a_{t},o_{t+1})$ checks that executing $a_{t}$ yields meaningful, schema-valid retrieval in $o_{t+1}$; and (iii) answer $v_{\mathrm{ans}}(m_{T})$ checks the final answer’s format/consistency on the final turn. the turn rewards are computed as follows with weight $w_{fmt},w_{kg},w_{ans}$:
$r^{\mathrm{turn}}_{t}=w_{\mathrm{fmt}}\,v_{\mathrm{fmt}}(m_{t})+w_{\mathrm{kg}}\,v_{\mathrm{kg}}(a_{t},o_{t+1})+w_{\mathrm{ans}}\,v_{\mathrm{ans}}(m_{T}).$ (1)
Global rewards summarize trajectory-level outcomes as the sum of the following: (i) a final-answer F1 score over the predicted set $\hat{A}_{q}$ vs. ground-truth answer set $A_{q}$; and (ii) a retrieval score $v_{ret}$ that is $1$ if any gold entity appears anywhere in the retrieved information along the executed reasoning path, and 0 otherwise.
$R^{\mathrm{global}}=w_{\mathrm{F1}}\cdot\mathrm{F1}\!\bigl(\hat{A}_{q},A_{q}\bigr)\;+\;w_{\mathrm{ret}}\cdot v_{ret}$ (2)
Group-relative turn-level credit assignment and optimization. To convert rewards into token-level credit, we use a group-relative, turn-level credit assignment inspired by Zeng et al. (2025). We collect $N$ rollouts per query $q$. For each episode $\tau^{(n)}$, we compute turn rewards $r^{\mathrm{turn},(n)}_{t}$ and a global reward $R^{\mathrm{global},(n)}$, and then form turn-specific returns with $\lambda$ as the normalization factor:
$G_{t}^{(n)}\;=\;r^{\mathrm{turn},(n)}_{t}\;+\;\lambda\,R^{\mathrm{global},(n)}.$ (3)
Let $\mathcal{S}\!=\!\{(n,t)\!:\!t\leq T^{(n)}\}$ be the set of all turns across the $N$ rollouts. The turn-level advantage $A_{t}^{n}$ is calculated using a single group baseline $\bar{G}$ that averages over $\mathcal{S}$:
$A_{t}^{(n)}=\frac{G_{t}^{(n)}-\bar{G}}{\sigma_{G}+\varepsilon}\quad\text{where}\quad\bar{G}=\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}G_{t}^{(n)},\ \sigma_{G}=\sqrt{\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}\!\big(G_{t}^{(n)}-\bar{G}\big)^{2}}.$
where $\varepsilon$ is a small constant for numerical stability.
RL update. Using turn-level credit $A_{t}^{(n)}$, we optimize the agent’s policy $\pi_{\theta}$ with a GRPO-style objective $\mathcal{J}$ :
$\mathcal{J}(\theta)=\mathbb{E}\!\Bigg[\sum_{n,t,i}\min\!\Big(\rho_{t,i}^{(n)}\,\widetilde{A}_{t,i}^{(n)},\ \operatorname{clip}\big(\rho_{t,i}^{(n)},\,1-\epsilon,\,1+\epsilon\big)\,\widetilde{A}_{t,i}^{(n)}\Big)\;-\;\beta\,\mathrm{KL}\!\big(\pi_{\theta}(\cdot\mid h_{t,i}^{(n)})\,\|\,\pi_{0}(\cdot\mid h_{t,i}^{(n)})\big)\Bigg]$
with $\widetilde{A}_{t,i}^{(n)}=m_{t,i}^{(n)}A_{t}^{(n)}$. $\pi_{\theta}$ is the current policy, and $\pi_{\theta_{\text{old}}}$ is the behavior policy used for sampling; $\pi_{0}$ is a fixed reference policy used only for KL regularization (weight $\beta$). $\rho_{t,i}^{(n)}=\exp\!\big(\log\pi_{\theta}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})-\log\pi_{\theta_{\text{old}}}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})\big)$ is the importance ratio that corrects off-policy sampling at token $i$. We maximize $\mathcal{J}(\theta)$ by gradient ascent. The proposed group-based credit assignment per turn produces stabilized and effective signals in the multi-turn interaction decisions. We provide comprehensive ablation studies of each component of our reward design in Section 4.3.
3.4 Cross-KG Transfer: Plug-and-Play
We propose a plug-and-play approach for KG-R1 that enables KG–RAG to operate across different knowledge graphs without retraining. Let $\pi_{\theta\mid\mathcal{D}}$ denote the policy of the KG-R1 agent trained on a KGQA dataset $\mathcal{D}=\{(q,G,A_{q})\}$. For a new KGQA dataset $\mathcal{D}^{*}=\{(q^{*},G^{*},A_{q^{*}})\}$, we replace the backend KG of the retrieval server with $G^{*}$ (plug) and evaluate $\pi_{\theta\mid\mathcal{D}}$ on $\mathcal{D}^{*}$ without any modification (play). This approach requires no task-specific fine-tuning, data relabeling, or architectural changes.
This plug-and-play capability stems from two design choices in KG-R1. First, the KG server exposes a schema-agnostic action set consisting solely of 1-hop retrieval operations, which are universally compatible with directed KGs. This contrasts with SPARQL-generation approaches in prior work that depend on KG-specific syntax and schema knowledge. Second, the agent’s strategy for retrieving and using KG evidence is learned in a way that is independent of any particular schema, enabling immediate transfer to new domains.
4 Experiments
We assess KG-R1 along two axes. (1) Performance and Efficiency on Trained KGs: We train KG-R1 on the training splits of KGQA benchmarks and measure answer quality and inference cost on their held-out test splits. (2) Cross-KG transferability: We test how well KG-R1 carries over to KGs out of training. In both parts, we compare against established baselines to support the KG-R1 results.
4.1 Performance and efficiency on Trained KGs
Models. We use Qwen2.5-3B-it (Qwen et al., 2025) as our base model. For the other baseline methods, we use GPT-4o-mini, GPT-3.5, and Llama2-7B-it, following the setups used in prior work.
Datasets. We mainly evaluate KG-R1 in-domain on (1) WebQSP (Yih et al., 2016)—natural-language questions over Freebase, mostly 1–2 hop, using the official 3,098/1,639 train/test QA split; and (2) ComplexWebQuestions (CWQ) (Talmor & Berant, 2018)—compositional, multi-hop questions over Freebase with a 24,649/3,531 train/test QA split. For scalability, we extract 2-hop subgraphs for each question as in RoG (Luo et al., 2024). See Appendix B for detailed sources.
Metrics. Following prior work (Luo et al., 2024; Sun et al., 2024) to evaluate KGQA performance, we use F1 score to consider multiple valid predictions over answer sets and Hit@1 (1 if the gold answer appears in the single search, 0 otherwise). For efficiency, we report generated tokens (completion-only), measured end-to-end per query, aggregated across all turns ($1\ldots H$). This serves as a proxy for inference time and compute cost. For fair comparison, all token counts are computed with the Qwen2.5 tokenizer. We also report the number of modules to compare workflow complexity, and we analyze single-query latency and batched throughput on a single GPU node.
Baselines. We compare three classes of approaches: (1) LLM-only methods that do not access an external KG—Vanilla and Chain-of-Thought (CoT) prompting; (2) prior KG-augmented pipelines: RoG (Luo et al., 2024), which uses a fine-tuned modular workflow with LLaMA2-7B-it as the backbone, and ToG (Sun et al., 2024) and ReKnoS (Wang et al., 2025c), which use a prompt-based modular workflow with GPT-3.5; and (3) our method KG-R1. For LLM-only baselines, we employ LLM-as-Judge (Zheng et al., 2023) using gpt-4o-mini to semantically match predicted answers to ground truth. Full details for baselines are provided in Appendix C.
N-run beam search with KG-R1. Prior works often incorporated k-beam search (Lan & Jiang, 2020; Sun et al., 2024; Luo et al., 2024), retrieving k reasoning paths for enhanced search range on KGs and reporting improved performance. Following this idea, we propose KG-R1 N-run beam search : we run $N$ independent generation per question and collect predicted answer sets $\{\hat{A}_{q}^{(i)}\}_{i=1}^{N}$. We then take the union $\hat{A}_{q,\text{N-runs}}\;=\;\bigcup_{i=1}^{N}\hat{A}_{q}^{(i)},$ and compute F1 and Hit@1 on this union. This simple “$N$-beam” setup broadens search without extra orchestration, and cost scales roughly linearly with $N$ since runs can be executed in parallel.
Implementation details. We use VERL (Sheng et al., 2024) and vLLM-backed decoding (Kwon et al., 2023) for implementing the KG-R1 agent. Training uses distributed optimization with gradient checkpointing and offloading. Unless stated, we set $H=5$ turns and N=16N{=}16 rollouts per query. Additional implementation details (e.g., hyper-parameters) are provided in Appendix D.4.
4.1.1 KG-R1 Training Stability and Effectiveness
Figure 4 shows steady F1 improvement that plateaus after convergence during KG-R1 training on WebQSP and CWQ. The validation F1 scores track the training curve, indicating the agent learns generalization. Results are reproducible across three random seeds (Figure 6, Appendix), showing strong convergence with low variance. Overall, KG-R1 training yields stable, reproducible gains in end-to-end KG-RAG performance.
Figure 4: F1 score over KG-R1 training on WebQSP and CWQ for Qwen2.5-3B-it. Training (blue) and validation (red).
4.1.2 Inference Accuracy and Efficiency Results
Table 1: Performance and efficiency comparison of KGQA methods on WebQSP and CWQ datasets. The max turn number set to $H=5$. All scores are percentages.
Method Model Modules WebQSP CWQ Performance Efficiency Performance Efficiency F1 Hit@1 Total Gen F1 Hit@1 Total Gen LLM-only Methods Vanilla Qwen2.5-3B-it 1 29.4 46.6 95 30 16.6 21.1 104 42 Vanilla Qwen2.5-7B-it 1 33.2 50.9 95 67 20.7 25.7 104 92 Vanilla LLaMA2-7B-it 1 37.4 54.5 114 255 20.7 24.8 123 255 COT Qwen2.5-3B-it 1 30.6 47.6 131 192 17.3 21.4 140 216 COT Qwen2.5-7B-it 1 35.3 53.5 131 194 22.5 27.1 140 212 COT LLaMA2-7B-it 1 33.8 51.6 165 255 19.0 23.1 174 255 LLM+KG Methods RoG LLaMA2-7B-it 2 70.8 85.7 1.2K 295 56.2 62.6 1.1K 266 ToG GPT-3.5 4 72.8 76.2 3.5K 529 52.9 57.1 3.7K 520 ToG 2.0 GPT-3.5 5 74.5 77.8 39.K 605 65.8 68.9 3.8K 650 ReKnoS GPT-4o-mini 3 73.7 81.1 3.1K 617 64.7 66.8 4.1K 752 KG-R1 (Our Methods) KG-R1 (1 run) Qwen2.5-3B-it 1 77.5 84.7 3.2K 302 70.9 73.8 3.3K 377 KG-R1 (3 runs) Qwen2.5-3B-it 1 85.8 91.7 9.7K 906 81.0 83.9 10.0K 1.1K
A1. KG-R1 demonstrates strong accuracy on trained KGs. Table 1 reports accuracy and efficiency. On both WebQSP and CWQ, a single run (N=1N{=}1) of KG-R1 achieves competitive F1 and Hit@1 compared with non-KG and prior KG methods, including those using larger foundation models. With three runs with beam search, KG-R1 surpasses prior LLM+KG systems that also use $k$-beam search by a clear margin (WebQSP: F1 $77.5{\to}85.8$, Hit@1 $84.7{\to}91.7$; CWQ: F1 $70.9{\to}81.0$, Hit@1 $73.8{\to}83.9$). Overall, these results show that a lightweight KG-R1 agent delivers strong KG-RAG performance relative to previous baselines.
A2. KG-R1 achieves efficient inference. Table 1 reports efficiency (generation tokens and number of modules). For a single run (N=1N{=}1), KG-R1 produces $\sim$60 generation tokens per turn (300–370 tokens at H=5H{=}5), substantially below prior prompt-based LLM+KG systems (e.g., ToG $\sim$520 and ReKnoS $\sim$600–650) and comparable to the fine-tuned method RoG ($\sim$250–300). This small generation-token budget, coupled with a lightweight model, implies improved throughput and latency. Under N=1N{=}1, the total tokens are comparable to prior methods, indicating a similar memory (KV-cache) footprint. With three-run beam search (N=3N{=}3), token usage scales nearly linearly, exposing a tunable accuracy–efficiency trade-off via $N$.
A3. Latency and throughput analysis. On a single NVIDIA H100 GPU, we measure (i) single-query end-to-end latency and (ii) batched throughput. The single-query latency averages $6.4\pm 1.5$ s per question. Batched throughput reaches $3.7$ samples per second at batch size 64. The results suggest the feasibility of KG-R1’s single-node deployment. See Appendix E.2 for the full results.
4.2 Cross-KG Transfer via Plug-and-Play
In this subsection, we evaluate the plug-and-play approach (Sec. 3.4) for zero-shot generalization of our KG-R1 agent on knowledge graphs outside its training distribution.
Datasets We evaluate the plug-and-play approach on diverse KGQA benchmarks spanning three categories of KGs. (1) Freebase type: SimpleQA (Bordes et al., 2015) and GrailQA (Gu et al., 2021), which share the Freebase KG but differ in question complexity and reasoning path distributions. (2) Wikidata type: T-REx (Elsahar et al., 2018) and QALD-10en (Usbeck et al., 2023) for cross-schema generalization. (3) Temporal reasoning benchmark: MultiTQ (Chen et al., 2023), which requires reasoning over the ICEWS temporal KG (Boschee et al., 2015) at day, month, and year granularities. Appendix B provides a summary of dataset sources, schema, and evaluation splits.
Baselines For LLM-only methods, we evaluate (i) Vanilla and (ii) Chain-of-Thought (CoT) prompting with Qwen2.5–3B-it. For LLM+KG methods, we report results for RoG, ToG, and ReKnoS without modifying their modules (i.e prompts and models). In addition, for each KGQA benchmark, we include one or two recent state-of-the-art KG-specific methods (trained/tuned on the target dataset with task-specific pipelines) and use their published test-set scores. Full baseline details are provided in Appendix C.3.
Table 2: Zero-shot cross-KG transferability of KG-R1. Agents are trained on WebQSP or CWQ and evaluated on new benchmarks by swapping only the KG-backend server (no policy retraining). AVG. is averaged across QA samples. The max turn number set to $H=5$. All scores are percentages.
Freebase-based Wikidata-based Temporal
Method Trained KG SimpleQA GrailQA T-REx QALD-10en MultiTQ AVG.
F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1
LLM-only Methods
Vanilla Qwen 2.5 3B-IT 13.7 / 13.7 15.9 / 15.9 24.4 / 24.4 23.8 / 23.8 2.2 / 5.4 19.4 / 19.8
CoT Qwen 2.5 3B-IT 10.9 / 10.9 18.2 / 18.2 22.0 / 22.0 25.9 / 25.9 1.9 / 3.9 18.0 / 18.2
LLM+KG Methods
KG-specific Baselines
$79.6^{1}/80.6^{1}$
$76.1^{2}/77.1^{2}$
$84.4^{3}/79.1^{3}$
$64.9^{4}/59.6^{4}$
— / $85.1^{5}$
$49.8^{6}/50.8^{6}$
$62.2^{7}/63.2^{7}$
— / $79.7^{8}$
— / $38.0^{9}$
RoG 13.5 / 43.5 6.4 / 23.7 8.1 / 36.4 11.4 / 44.3 3.6 / 8.7 8.1 / 34
ToG 50.2 / 53.6 63.9 / 68.7 79.3 / 76.4 48.7 / 50.2 25.1 / 27.9 66.2 / 65.9
ReKnoS 50.7 / 52.9 65.8 / 71.9 81.4 / 78.9 50.6 / 53.8 28.5 / 31.1 68.2 / 68.2
KG-R1 (Our Methods)
KG-R1 (1 run) WebQSP 59.1 / 59.1 32.8 / 38.5 80.5 / 84.5 51.9 / 53.4 21.6 / 31.4 64.0 / 68.3
KG-R1 (1 run) CWQ 64.6 / 64.7 42.8 / 50.2 81.3 / 85.6 55.9 / 57.7 27.1 / 38.9 67.2 / 72.1
KG-R1 (3 runs) WebQSP 69.1 / 69.1 44.6 / 52.1 84.6 / 88.8 61.3 / 62.8 33.5 / 45.7 70.9 / 75.8
KG-R1 (3 runs) CWQ 73.1 / 73.1 52.8 / 61.0 86.8 / 91.5 63.9 / 65.5 36.2 / 48.4 74.1 / 79.4
1SPARKLE 2GETT-QA 3SG-KBQA 4ARG-KBQA 5ATLAS 6COT-SPARQL 7DFSL-MQ 8Prog-TQA 9ARI ; see Table 7 for references.
A4. KG-R1 transfers across diverse types of KGs. Table 2 reports zero-shot, plug-and-play results. For comparability, we report AVG., the F1/Hits@1 averaged across the five KGQA datasets, weighted by QA counts. KG-R1 (1 run) substantially outperforms LLM-only baselines using the same base model—Vanilla (19.4/19.8 F1/Hits@1) and CoT (18.0/18.2)—reaching 64.0/68.3 when trained on WebQSP and 67.2/72.1 when trained on CWQ. The gains hold not only on Freebase-based QA but also on Wikidata and Temporal benchmarks, indicating that reinforcement-learned KG-RAG transfer across diverse KGs rather than narrow, in-domain improvements.
A5. KG-R1 achieves accurate inference comparable to LLM+KG baselines. Against LLM+KG baselines, KG-R1 (N=1) is slightly better on average: ToG (66.2/65.9) and ReKnoS (68.2/68.2). Notably, KG-R1, with a 3B model, delivers higher Hits@1 than both ToG and ReKnoS that uses strong foundation models. Increasing to N=3 runs boosts performance to 74.1/70.4 (CWQ-trained), making KG-R1 comparable to KG-specific baselines on average. We observe that RoG collapses (0/0 average) revealing brittle, schema-specific planning with fine-tuned model that does not transfer. Overall, KG-R1 attains strong transferability among plug-and-play methods with a lightweight model, supporting its practicality for real-world KG-RAG deployment across diverse KGs.
A6. Training data on transferability. KG-R1 models trained on CWQ consistently outperform WebQSP-trained variants across all transfer targets, averaging 1–5% higher F1 and Hits@1. Given that CWQ provides a larger training set with greater multi-hop complexity than WebQSP, this suggests that exposure to more complex and diverse reasoning patterns yields superior generalization.
4.3 Ablation Studies
We conduct ablation studies for the three core components of KG-R1: reward design (Section 3.3), RL algorithm (Section 2), and KG-server implementation (Section 3) and base LLM size. We train Qwen-2.5-3B-it on WebQSP and CWQ with maximum turn number $H=5$ and report final step performance(F1/Hit@1/Ret.rate in Table 3). The training curves are shown in Figures 8, 9.
Table 3: Ablation studies of KG-R1 components on CWQ and WebQSP datasets.
Method WebQSP CWQ F1 Hit@1 Ret.Rate F1 Hit@1 Ret.Rate KG-R1 Default KG-R1 77.2 84.3 76.8 66.8 69.7 51.4 Reward Ablation w/o Turn Rewards 67.1 (-13.1%) 76.0 (-9.8%) 68.1 51.9 (-22.3%) 52.9 (-24.1%) 44.7 w/o Turnwise Advantage 49.6 (-35.8%) 61.2 (-27.4%) 22.8 63.5 (-4.9%) 64.5 (-7.5%) 49.4 w/o Retrieval Reward 45.8 (-40.6%) 57.7 (-31.5%) 18.6 66.0 (-1.2%) 67.0 (-3.9%) 47.2 Other Ablations w PPO 0.0 (-100.0%) 0.0 (-100.0%) 0.0 0.0 (-100.0%) 0.0 (-100.0%) 0.0 w/o Hierarchical Rel. Retr. 76.8 (-0.6%) 83.1 (-1.5%) 76.0 55.5 (-16.9%) 58.6 (-15.9%) 39.7 w Qwen-2.5-7B-it 79.6 (+3.1%) 86.6 (+2.7%) 83.4 (+8.6%) 65.6 (-1.8%) 68.7 (-1.4%) 58.8 (+14.4%)
PPO training crashed.    Peak values before 7B model training collapse.
A7. Reward function. First, removing turn-level rewards yields the largest drop on both WebQSP (13.1% / 9.9% in F1 score and Hit@1, respectively) and CWQ (22.3% / 24.1%), highlighting the importance of per-turn validity/effectiveness signals. Next, disabling the turn-wise advantage calculation produces a substantial decline on WebQSP (35.8% / 27.4%) but only a moderate drop on CWQ (4.9% / 7.5%), indicating that turn-specific learning signals matter more for certain datasets. Finally, removing the retrieval reward significantly impacts WebQSP (40.7% / 31.5%) but marginally affects CWQ (1.2% / 3.9%), though it substantially reduces the retrieval success rate of gold entities (Ret.Rate) on both datasets, suggesting that it is essential for motivating exploration of the KG.
A8. RL algorithm (GRPO vs. PPO). Replacing GRPO with vanilla PPO (Ouyang et al., 2022) results in learning collapse (0.0 across all metrics). Under PPO, we observe reward hacking (see examples in Appendix E.3.3): the agent fabricates ”retrieved” content matching expected formats to earn reward. Since PPO’s value critic (another LLM) cannot distinguish genuine KG retrievals from hallucinations, its advantage estimates become unstable, driving training collapse. This underscores the importance of relative advantage methods like GRPO for multi-step reasoning.
A9. Retrieval format. In our standard KG-R1 setup, one-hop relations retrieved by $\mathtt{get\_head\_relations}$ and $\mathtt{get\_tail\_relations}$ are appended to the agent’s context as a flat, comma-separated list (e.g., “rel1, rel2, …”). This can lead to excessive token consumption when many relations are retrieved. Because Freebase relations use a compositional format (e.g., “domain.type.property”), we tested a hierarchical relation retrieval that groups relations into a tree to reduce redundancy (see Appendix E.3.2). Our ablation studies show that switching from flat lists to a hierarchical format reduces performance—WebQSP shows 0.5% drop, while CWQ experiences 16.9% drop. These results suggest that despite its higher token usage, the flat retrieval format provides more direct access to relation information, which proves critical for training.
A10. Model scale. The 7B model improves more quickly than the 3B model but exhibits training collapse (around 350 steps on WebQSP and 200 steps on CWQ). This aligns with reports that fast distribution shift during GRPO—amplified by larger models and higher update rates—drives entropy collapse and model divergence (Jin et al., 2025; Liu et al., 2025).
5 Conclusion
We propose KG-R1, a single-agent, multi-turn KG-RAG framework in which one LLM queries a lightweight KG server and is optimized via RL. Across KG-augmented QA benchmarks, KG-R1 attains strong accuracy while using markedly fewer tokens and lower inference cost than prior work. It also supports plug-and-play transfer, retaining robust performance across KGs. Together, these results position KG-R1 as a practical and deployable KG-RAG system for real-world use.
Appendix
Appendix A Theoretical Proofs
A.1 Completeness and Transferability of the KG-Server
Here we provide robust proofs for completeness and transferability of $\mathcal{U}_{ret}$ in our KG retrieval server.
Preliminaries. Let $G=\{(e,r,e^{\prime})\mid e,e^{\prime}\!\in\!\mathcal{E},\,r\!\in\!\mathcal{R}\}$ be a directed KG. Define the KG-R1 action set
$\mathcal{U}_{\text{ret}}=\left\{\begin{aligned} &\mathtt{get\_tail\_relations},\ \mathtt{get\_head\_relations},\\ &\mathtt{get\_tail\_entities},\ \mathtt{get\_head\_entities}\end{aligned}\right\}$
with the following semantics for any $(e,r,e^{\prime})\in G$:
$\mathtt{get\_tail\_relations}(e)$ $:=\{\,r\in\mathcal{R}\mid\exists e^{\prime}\in\mathcal{E}:\ (e,r,e^{\prime})\in G\,\},$
$\mathtt{get\_head\_relations}(e^{\prime})$ $:=\{\,r\in\mathcal{R}\mid\exists e\in\mathcal{E}:\ (e,r,e^{\prime})\in G\,\},$
$\mathtt{get\_tail\_entities}(e,r)$ $:=\{\,e^{\prime}\in\mathcal{E}\mid(e,r,e^{\prime})\in G\,\},$
$\mathtt{get\_head\_entities}(r,e^{\prime})$ $:=\{\,e\in\mathcal{E}\mid(e,r,e^{\prime})\in G\,\}.$
A relation path is $z=(r_{1},\ldots,r_{\ell})$, and a reasoning path (instantiation) is
$Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell},\quad(e_{i-1},r_{i},e_{i})\in G.$
Proposition A.1 (Finite-Horizon Bound).
For any path $Z$ of length $\ell$, there exists an action sequence from $\mathcal{U}_{\text{ret}}$ of length at most $\ell$ whose final output contains $e_{\ell}$.
Proof. Along the path $Z$, for each step $i=1,\ldots,\ell$ the call $\mathtt{get\_tail\_entities}(e_{i-1},r_{i})$ returns a set that contains $e_{i}$. Therefore, after the $\ell$-th step the output contains $e_{\ell}$.
Proposition A.2 (Completeness (Integrity)).
Every finite reasoning path in $G$ can be realized by a finite sequence of actions from $\mathcal{U}_{\text{ret}}$ whose outputs include its terminal entity.
Proof. Starting with $\ell=1$, if $(e_{0},r_{1},e_{1})\in G$ then $e_{1}\in\mathtt{get\_tail\_entities}(e_{0},r_{1})$. For the inductive step, assume the claim holds for length $\ell-1$. By the hypothesis we reach $e_{\ell-1}$; applying $\mathtt{get\_tail\_entities}(e_{\ell-1},r_{\ell})$ then returns a set containing $e_{\ell}$.
Proposition A.3 (Schema-Free Transferability).
The semantics of each $a\in\mathcal{U}_{\text{ret}}$ depend only on membership of triples $(e,r,e^{\prime})$ in $G$. Replacing $G$ with any directed KG $G^{\prime}$ preserves action meaning and allows a fixed client policy to transfer unchanged.
Proof. Each operator is defined as a set comprehension over triples in $G$, independent of schema, datatypes, or ontology rules. Under $G\mapsto G^{\prime}$, the same comprehensions define the actions on $G^{\prime}$, so a learned policy remains well-defined.
Proposition A.4 (Minimality of $\mathcal{U}_{\text{ret}}$).
No proper subset of $\mathcal{U}_{\text{ret}}$ supports both symmetric navigation and local relation discovery; any operator that only filters or composes these four is derivable.
Proof. Removing $\mathtt{get\_tail\_entities}$ (resp. $\mathtt{get\_head\_entities}$) prevents forward (resp. backward) traversal. Removing both $\mathtt{get\_tail\_relations}$ and $\mathtt{get\_head\_relations}$ blocks local relation discovery when the relation set is unknown. Conversely, a conditional/filtered variant can be constructed by enumerating $\mathtt{get\_tail\_relations}(e)$ and intersecting results of $\mathtt{get\_tail\_entities}$; hence no additional primitives are required.
Appendix B Datasets
B.1 KGQA Datasets
These are the KGQA datasets we used in our work.
Freebase
WebQSP (Yih et al., 2016) (full test; 1,639) Open-domain questions with mostly 1–2 hop reasoning over Freebase MIDs and SPARQL annotations.
ComplexWebQuestions (CWQ) (Talmor & Berant, 2018) (full test; 3,531) Compositional multi-hop questions generated from SPARQL templates that stress longer chains and constraint composition; used to probe multi-turn retrieval quality and robustness without dataset-specific rules.
SimpleQuestions (SimpleQA) (Bordes et al., 2015) (1,000 test) Single-relation 1-hop questions over Freebase; serves as a retrieval-fidelity and token-efficiency baseline for KG-R1. We randomly sample 1,000 QA from the original test split (21,687).
GrailQA (Gu et al., 2021) (1,000 test) Diverse compositional questions emphasizing generalization under Freebase; handled with the same minimal action interface and no hand-crafted schemas. We randomly sample 1,000 from the original test split (13,231).
Wikidata
T-REx (Elsahar et al., 2018) (5,000 test) Large-scale slot-filling–style QA grounded in Wikidata triples; used to assess scalability and coverage under a different KG schema. We randomly sample 5,000 from the corpus ($\sim$2.2M).
QALD-10en (Usbeck et al., 2023) (333 test) Manually curated, linguistically varied questions over Wikidata; useful for evaluating precision on a small but challenging set. We evaluate on 333 examples.
Temporal KG
MultiTQ (Chen et al., 2023) (1,000 test) Time-aware QA requiring temporal qualifiers (e.g., during, from–to); evaluates KG-R1 on temporal reasoning with time-scoped entities and relations. We randomly sample 1,000 from the original test split (9,000).
B.2 Dataset Preprocessing
Two-hop Subgraph Extraction Methodology.
Following subgraph preprocessing practice RoG (Luo et al., 2024), we builds a question-specific near subgraph to shrink the search space and suppress spurious matches: starting from the linked anchor entities $e_{q}$ and gold answers $A_{q}$, it performs a breadth-first expansion over the KG $G$ up to the dataset’s maximum 2-hop radius $h$ The processed subgraph for each question is cached and used for KG retrieval server in KG-R1.
B.3 Freebase, Wikidata, and MultiTQ schema comparison
The following tables show the different schemas of Freebase, Wikidata, and MultiTQ.
Table 4: Freebase knowledge graph dataset used for KGQA evaluation. All datasets share the same underlying Freebase knowledge graph structure with 4.9M entities and 663 relations.
Dataset Entities Relations
Freebase - P!nk - Gender - Ice Cube - United States of America - Nemanja Mikic - broadcast.content.artist - people.person.nationality - music.artist genre - location.location.containedby - basketball.basketball.player position
Questions
WebQSP (4,737 test questions): Q: what does jamaican people speak A: [Jamaican English, Jamaican Creole English Language] CWQ (3,531 test questions): Q: Lou Seal is the mascot for the team that last won the World Series when? A: [2014 World Series] SimpleQA (21,687 test questions): Q: where is the madam satan located A: [United States of America] GrailQA (13,231 test questions): Q: which play is produced by the illusion? A: [The Illusion]
Table 5: Wikidata knowledge graph datasets used for KGQA evaluation. All datasets share the same underlying Wikidata knowledge graph structure with 15M entities and 2.3K relations.
Dataset Entities Relations
Wikidata - Barack Obama - Germany - Albert Einstein - Microsoft - Paris - instance of - country - occupation - date of birth - place of birth
Questions
T-Rex (2.2M test questions): Q: What is the occupation of Albert Einstein? A: [Physicist] QALD-10en (250 test questions): Q: Which companies were founded by Bill Gates? A: [Microsoft]
Table 6: Temporal knowledge graph dataset used for KGQA evaluation. MultiTQ focuses on temporal reasoning with time-aware entities and relations.
Dataset Entities Relations
Temporal KG - Barack Obama (2009-2017) - World War II (1939-1945) - Steve Jobs (1955-2011) - Cold War (1947-1991) - Nelson Mandela (1994-1999) - president during - occurred during - CEO from to - started in - ended in
Questions
MultiTQ (9,000 test questions): Q: Who was the president of the United States when the iPhone was first launched? A: [George W. Bush] Q: Which major historical event ended the same year the European Union was established? A: [Cold War] Q: What technology company was founded during World War II? A: [IBM]
Appendix C Baselines
C.1 LLM-only Methods
Setup. We evaluate Vanilla and CoT in a zero-shot setting without access to the KG retrieval server (i.e., no KG augmentation). All baselines run on Qwen-2.5-3B-IT, Qwen-2.5-7B-IT, and LLaMA-2-7B-Chat with temperature=0 and top_p=50..
Prompts. We use the following prompt templates for vanilla and CoT baselines.
Prompte template for Vanilla setup Answer the given question directly and concisely based on your knowledge. Format your answer as: Answers: [”answer1”, ”answer2”, …]. For single answers, use: Answers: [”answer”]. Question: [Question] Answers:
Prompte template for CoT setup Think through the question step by step, then provide the answer. IMPORTANT: Follow this exact format: 1. Start with ”Reasoning:” followed by your step-by-step thinking. 2. End with ”Answers:” followed by your final answer in brackets. 3. Do NOT put ”Answers:” before your reasoning. Format your answer as: Answers: [”answer1”, ”answer2”, …]. For single answers, use: Answers: [”answer”]. Question: [Question] Reasoning:
Evaluation. Baseline outputs (Vanilla and CoT) differ in format from the KGQA gold answer set, so we employ an LLM-as-judge to align them. For each question, gpt-5-mini (OpenAI API) is given the question, the ground-truth answer set $A_{q}$, and parsed predited entities $\hat{A}_{q}$ from the base model’s output with a concise semantic-entity-matching prompt (Box below); it returns only a binary vector indicating, for each gold entity in order, whether the prediction refers to the same real-world entity at the same specificity (1 for exact/equivalent matches, e.g., “Apple Inc.” = “Apple”; 0 for overly general predictions, e.g., “Islam” vs [“Shia Islam”, “Sunni Islam”]). We report Pass@K (K=1,2,3,4), F1, precision, recall, and generation-token usage.
Prompt template for LLM-as-Judge You are an evaluator performing semantic entity matching. Your task is to decide, for each gold entity (in order), whether the model’s prediction refers to the same real-world entity with the same level of specificity. Respond only with a binary vector in the format: $[0,1,0,1]$ Rules: - Exact matches or equivalent references → output 1 (e.g., ”Apple Inc.” = ”Apple”). - Too general or not specific enough when specificity is required → output 0 (e.g., ”Islam” vs [”Shia Islam”, ”Sunni Islam”]). Gold entities: [gold entity list] Predicted entities: [predicted entity list] Your Response:
C.2 LLM+KG baselinse
RoG (Luo et al., 2024) Reasoning on Graphs (RoG) couples LLMs with KGs via a planning–retrieval–reasoning pipeline: the LLM first proposes KG-grounded relation paths as faithful plans, uses them to retrieve valid paths from the KG, and then performs stepwise reasoning to produce an answer. This yields interpretable multi-hop reasoning and reduces hallucinations by constraining reasoning to KG structure.
ToG (Sun et al., 2024) Think-on-Graph (ToG) treats the LLM as an agent that iteratively explores the KG: it performs beam search over entities/relations, expands promising paths, and alternates retrieval with reasoning until a final path/answer is selected. Compared to prompt-only baselines, ToG improves deep, compositional reasoning by explicitly navigating the KG during multi-hop search.
ReKnoS (Wang et al., 2025c) Reasoning with Knowledge Super-relations (ReKnoS) introduces super-relations that summarize and connect multiple relational paths, enabling both forward and backward reasoning while expanding the search space the LLM can traverse. By reasoning over these super-relations (rather than isolated edges), ReKnoS boosts retrieval success and multi-hop accuracy, especially on complex queries.
C.3 Cross-KG generliazation KG-specific Baselines
KG-Specific Baselines We selected one or two recent (2022–) state-of-the-art KG-specific methods, taking the authors’ reported test/dev scores as basline comparison. The below outlines baseline methods reported in Sec. 4.2 by KGQA benchmarks. The table 7 shows the summary.
[…]
Table 7: KG-specific Baselines For rows where Hits@1 is not reported, we show EM when available; where neither is reported, we show ”–”. Numbers are as reported in the cited papers/leaderboards.
[…]
GrailQA. SG-KBQA (Gao et al., 2025) is a schema-guided system that conditions a LLaMA-3.1-8B backbone on schema context and generates executable logical forms; official leaderboard reports overall Test EM/F1. ARG-KBQA (Tian et al., 2024) retrieves logical-form exemplars via an unsupervised ranker, then prompts GPT-3.5-0613 to generate and execute candidate logical forms.
QALD-10-en. COT-SPARQL (D’Abramo et al., 2025) applies chain-of-thought prompting with entity/relation hints to produce SPARQL; it reports both standard F1 and the Macro F1-QALD. DFSL-MQ (D’Abramo et al., 2024) performs dynamic few-shot retrieval with multi-query generation and beam selection (evaluated with CodeLlama-70B); it reports standard F1 on QALD-10-en.
T-REx. ATLAS (Izacard et al., 2023) is a retrieval-augmented seq2seq (FiD) reader paired with a dense retriever (Contriever); it reports Accuracyon T-Rex(treated as comparable to Hits@1).
MultiTQ (temporal KGQA). Prog-TQA (pro, 2024) uses in-context KoPL program generation with linking, execution, and a self-improvement loop; results are reported for LLaMA-13B/Vicuna-13B readers. ARI (Chen et al., 2024) (Abstract Reasoning Induction) separates planning (knowledge-agnostic) from execution (knowledge-based) with GPT-3.5-0613; it reports Accuracy (treated as comparable to Hits@1).
Notes: COT-SPARQL also reports Macro F1-QALD = 0.6387 on QALD-10 via GERBIL, while we list its standard F1 here for consistency across rows. ATLAS reports Accuracy on KILT hidden tests: zsRE = 80.8, T-REx = 85.1 (we do not map these to F1/Hits@1/EM). SG-KBQA numbers (EM=79.140, F1=84.403) are from the official GrailQA leaderboard (Test). ARG-KBQA is reported on Dev set (EM=59.6, F1=64.9). ARI reports Accuracy on MultiTQ (treated as comparable to Hits@1).
Appendix D KG-R1 details
D.1 Training algorithm
The following pseudocode outlines the training procedure for KG-R1.
Input: Dataset $D=\{(q,G,A_{q})\}$, base LLM $\pi_{0}$, horizon $H$, rollouts NN
Output: Trained policy $\pi_{\theta}$
$\pi_{\theta}\leftarrow\pi_{0}$ // Initialize from base LLM
foreach mini-batch of queries from $D$ do
 foreach $q$ in batch do
    Collect NN rollouts $\{\tau^{(n)}\}_{n=1}^{N}$;
    for $n\leftarrow 1$ to NN do
       $\tau^{(n)}\leftarrow(p)$ ;
       // where p is the instruction prompt for $q$
        for $t\leftarrow 1$ to $H$ ;
        // Multi-turn interaction
       do
          $r_{t}\sim\pi_{\theta}(\cdot\mid\tau^{(n)})$ ;
          // Generate response
          $a_{t}\leftarrow\Psi(r_{t})$ where $a_{t}\in\mathcal{A}_{\text{query}}\cup\{\texttt{answer}\}$;
          if $a_{t}\in\mathcal{A}_{\text{query}}$ then
               Execute $a_{t}$, get $o_{t+1}$, and append to $\tau^{(n)}$;
            end if
          else
               Extract $\hat{A}_{q}$ and break;
            end if
         end for
      end for
    Compute rewards for collected rollouts:;
      Turn: $r^{\text{turn},(n)}_{t}=w_{\text{fmt}}v_{\text{fmt}}(r_{t})+w_{\text{kg}}v_{\text{kg}}(a_{t},o_{t+1})+w_{\text{ans}}v_{\text{ans}}(r_{T})$;
      Global: $R^{\text{global},(n)}=w_{\text{F1}}\cdot\text{F1}(\hat{A}_{q},A_{q})+w_{\text{ret}}\cdot v_{\text{ret}}$;
   end foreach
 Credit assignment: $G_{t}^{(n)}=r^{\text{turn},(n)}_{t}+\lambda R^{\text{global},(n)}$;
 Group baseline: $\bar{G}=\frac{1}{|\mathcal{S}|}\sum_{(n,t)\in\mathcal{S}}G_{t}^{(n)}$, where $\mathcal{S}=\{(n,t):t\leq T^{(n)}\}$;
 Advantages: $A_{t}^{(n)}=\frac{G_{t}^{(n)}-\bar{G}}{\text{std}_{(n,t)\in\mathcal{S}}(G)+\epsilon}$;
 Update: $\pi_{\theta}$ via GRPO with $J(\theta)$, entropy $\mathcal{H}$, and KL divergence to $\pi_{0}$;
end foreach
Algorithm 1 KG-R1 Training with RL
D.2 Server Instruction Prompt
Server Instruction template consists of A,B,C. We use the same server instruction for both training and evaluation across experiments.
Server Instruction template for KG-R1 If you encounter a KG-related error, read the error message carefully and correct your query. Use exactly these query functions: -$\mathtt{get\_tail\_relations}$(entity) : Returns relations where the entity is the subject/head. -$\mathtt{get\_head\_relations}$(entity) : Returns relations where the entity is the object/tail. -$\mathtt{get\_tail\_entities}$(entity, relation) : Returns entities connected to the given entity by the specified relation. -$\mathtt{get\_head\_entities}$(entity, relation) : Returns entities from which the given entity is connected by the specified relation.
D.3 Reward weights
Our RL objective combines per-turn and episode-level signals with fixed weights. Per-turn rewards encourage well-structured interaction and effective retrieval: $r_{t}^{\mathrm{turn}}=w_{\mathrm{fmt}}\,s_{\mathrm{fmt}}+w_{\mathrm{kg}}\,s_{\mathrm{kg}}+w_{\mathrm{ans}}\,s_{\mathrm{ans}}$, where $s_{\mathrm{fmt}}$ scores output structure/format validity, $s_{\mathrm{kg}}$ rewards schema-valid, non-empty KG queries, and $s_{\mathrm{ans}}$ checks final-answer formatting/consistency. Episode-level reward emphasizes answer correctness and retrieval coverage: $R^{\mathrm{global}}=w_{\mathrm{F1}}\cdot\mathrm{F1}(\hat{A}_{q},A_{q})+w_{\mathrm{ret}}\cdot v_{\mathrm{ret}}$, with $v_{\mathrm{ret}}\in\{0,1\}$ indicating adequate retrieval coverage. Unless otherwise noted, we use the following Table 8 as defaults.
Table 8: Reward-component weights ($w$) for KG-R1 reward function.
Symbol Name Scope Role (concise) Default
$w_{\mathrm{fmt}}$ Format weight Turn Rewards valid per-turn output structure in $r_{t}^{\text{turn}}$ $0.5$
$w_{\mathrm{kg}}$ KG-query weight Turn Rewards schema-valid, non-empty KG retrieval in $r_{t}^{\text{turn}}$ $0.5$
$w_{\mathrm{ans}}$ Answer-format weight Turn Rewards correct final-answer formatting/consistency in $r_{t}^{\text{turn}}$ $0.5$
$w_{\mathrm{F1}}$ Final-answer weight Episode Weights $\mathrm{F1}(\hat{A}_{q},A_{q})$ in $R^{\text{global}}$ $1.0$
$w_{\mathrm{ret}}$ Retrieval-coverage weight Episode Rewards coverage signal $v_{\mathrm{ret}}\!\in\!\{0,1\}$ in $R^{\text{global}}$ $1.0$
Notes: Only reward-component weights are shown; optimization and rollout hyperparameters are omitted.
D.4 RL Implementation & Hyperparameter
Learning Rates and Optimizer. Both Qwen-2.5-3B-it and Qwen-2.5-7B-it use an identical learning rate of $1\times 10^{-6}$ with no warmup. We use the AdamW optimizer with weight decay $0.01$ applied to all parameters except biases and layer-normalization weights, and gradient clipping by global norm set to $1.0$ to prevent gradient explosion during RL training.
Batch Configuration and Gradient Accumulation. The 3B model uses a training batch size of $128$ with validation batch size of $256$, whereas the 7B model uses a training batch size of $256$ with validation batch size of $128$. During GRPO rollout, we collect $N=16$ trajectories per prompt for the 3B model and $N=8$ for the 7B model to balance exploration with memory constraints. The mini-batch size is $128$ for both models, with dynamic batch sizing enabled to utilize GPU memory efficiently.
RL Coefficients. RL training uses GRPO (Group Relative Policy Optimization) with multi-turn advantage computation enabled. The KL loss coefficient differs by model: $\beta=0.01$ for 3B and $\beta=0.02$ for 7B, using the K3 KL loss to limit divergence from the reference policy. The entropy coefficient is set to 0 for both models, favoring exploitation over exploration for multi-turn KG reasoning.
[…]
Hardware Specifications and Training Duration. Training is conducted on $4$ NVIDIA H100 GPUs for both model sizes. For the 3B model, we allow up to 21,00021{,}000 tokens/GPU for PPO processing and 18,38418{,}384 max batched tokens for VLLM rollout. For the 7B model, we allocate 12,00012{,}000 tokens/GPU for PPO and 12,28812{,}288 max batched tokens for VLLM, using FSDP parameter and optimizer offloading to fit memory constraints. Both configurations use bfloat16 precision with chunked prefill enabled. Training runs for $400$ steps with checkpoints saved every $50$ steps (3B) or every $25$ steps (7B).
D.5 KG-R1 Retreival Server details
Setup. We implemented the KG-R1 retrieval server with FastAPI (Ramírez, 2018) and Uvicorn (Encode, 2018) to support (i) a schema-free 1-hop KG query API, (ii) high-throughput async batch execution, and (iii) robust validation and observability (regex action parsing, standardized <information> wrapping with auto-closure, and final-turn safeguards). After the KG-R1 agent generates a response, the parsed <kg-query> action is sent to the server, which performs exact string matching over the per-question knowledge graph to resolve entities and relations and returns the retrieved information. If the call is invalid—one of: KG_SERVER_ERROR: Invalid Action; KG_FORMAT_ERROR: Missing Required Fields; KG_FORMAT_ERROR: Wrong Argument Count; KG_SAMPLE_NOT_FOUND: Sample Missing; KG_ENTITY_NOT_FOUND: Entity Not in KG; KG_RELATION_NOT_FOUND: Invalid Relation; KG_NO_RESULTS: No Relations Found; KG_NO_RESULTS: No Entities Found—the server responds with a descriptive error message (see box below).
KG-R1 Server Error Examples KG_SERVER_ERROR: Invalid Action <error>Action ”get_entity_info” not available (use: get_head_relations, get_tail_relations, get_head_entities, get_tail_entities)</error>  KG_FORMAT_ERROR: Missing Required Fields <error>Missing required fields for get_tail_entities: relation_name</error>  KG_FORMAT_ERROR: Wrong Argument Count <error>get_tail_relations accepts only one entity argument</error>  KG_SAMPLE_NOT_FOUND: Sample Missing <error>Sample ”sample_12345” not found in KG</error>  KG_ENTITY_NOT_FOUND: Entity Not in KG <error>Entity ”Barack Obamaa” not found in KG</error>  KG_RELATION_NOT_FOUND: Invalid Relation <error>Relation ”location.capital” not found in KG</error>  KG_NO_RESULTS: No Relations Found <error>No tail relations found for entity ”Random_Entity_123” in knowledge graph</error>  KG_NO_RESULTS: No Entities Found <error>No tail entities found for relation ”film.director.film” with head ”Barack Obama” in knowledge graph</error>
Figure 5: KG-R1 error types with actual server error messages.
Appendix E Supplementary Experimental Results
Figure 6 shows Training dynamics of Qwen-2.5b-it are highly consistent across three random seeds, with smooth F1 improvements and minimal variance throughout 400 optimization steps. Both WebQSP (red) and CWQ (blue) curves show rapid early gains and stable convergence, indicating robust optimization and reproducible policy learning.
E.1 Reproducibility of KG-R1 Training
Figure 6: Training dynamics of Qwen 2.5b-it across 3 random seeds demonstrate reproducibility with steady F1 improvement and low variance. WebQSP (red) and CWQ (blue) metrics over 400 steps show stable convergence.
E.2 Latency and Throughput Analysis
[…]
(a) Histogram of end-to-end latency per question.
(b) Cumulative time by agent turn within a query.
[…]
Batched Throughput We evaluate batched inference at batch size $64$ on a single NVIDIA H100 (Table 9). LLM-only baselines (no KG calls) achieve high throughput—$81.8$ (Vanilla) and $70.1$ (CoT) samples/s—driven by short generation (43.0/206.0 tokens per sample). KG-R1 incurs KG-server retrieval, reducing throughput but remaining practical for offline processing: the single-run setting reaches $3.7$ samples/s ($1205.9$ gen tokens/s) with $4.4$ KG calls per query, while the N=4N{=}4 runs setting trades throughput for more KG interaction ($17.5$ calls per query), yielding $2.0$ samples/s ($612.1$ gen tokens/s). Overall, KG-R1 sustains batch processing on one H100 while supporting KG-grounded reasoning.
Table 9: Batched throughput on one NVIDIA H100 (256 queries; batch size 64). Samples: total queries. Batch: batch size. KG Calls: total KG-server requests. Calls/Sample: average KG requests per query. Total (s): end-to-end wall-clock time. Gen Tok./Sample: generated tokens per query. Samples/s: queries per second. Gen Tok./s: generated tokens per second.
Configuration Samples Batch KG Calls Calls/Sample Total (s) Gen Tok./Sample Samples/s Gen Tok./s Vanilla Baseline 256 64 0 0.0 12.4 43.0 81.8 887.7 Vanilla CoT 256 64 0 0.0 14.1 206.0 70.1 3740.1 KG-R1 (single run) 256 64 1127 4.4 73.2 345.0 3.7 1205.9 KG-R1 (N=4N{=}4 runs) 256 64 4478 17.5 142.2 340.0 2.0 612.1
E.3 Ablation studies
All ablations in Table 3 evaluate KG-R1 with Qwen-2.5-3B-it trained on WebQSP and CWQ using a maximum turn budget of H=5H{=}5. We report the full ablation table in Table 3 (training curves in Figures 89).
Turn reward. We vary the turn-level reward by setting the weights to $w_{\text{fmt}}{=}0$, $w_{\text{kg}}{=}0$, and $w_{\text{ans}}{=}0$ (default: all $0.5$).
Turn-wise advantage. Instead of computing the turn-wise group advantage $A_{t}^{(n)}$ in Sec. 3.3, we compute a trajectory-wise group advantage
$A_{t}^{(n)}\;=\;\frac{G^{(n)}-\bar{G}}{\sigma_{G}+\epsilon},\qquad G^{(n)}\;=\;\frac{1}{T}\sum_{t}r_{t}^{\text{turn},(n)}\;+\;R^{\text{global},(n)},$
and use it for token-level credit assignment.
Retrieval reward. We ablate the retrieval component by setting the weight $w_{\text{ret}}{=}0$ (default: $1$).
RL algorithm: GRPO vs. PPO. We replace GRPO with vanilla PPO (VERL (Sheng et al., 2024) defaults), and set the turn-reward mixture weight $w_{\text{turn}}{=}0$ (default: $0.5$). Advantage estimation is performed by a learned value critic (Qwen-2.5-3B).
Hierarchical relation retrieval. We change the KG-server retrieval format from a flat list to a hierarchical format that mirrors “domain.type.property” (see Table LABEL:tab:hierarchical_relation_format).
Base LLM parameter size (7B). We swap the backbone from Qwen-2.5-3B-it to Qwen-2.5-7B-it while keeping data, rewards, and budgets fixed.
E.3.1 Training curves
Figure 8: Training curves of ablation studies WebQSP, reporting F1 score across training steps.
Figure 9: Training curves of ablation studies for CWQ, reporting F1 score across training steps.
E.3.2 Hiearchical retireval format
Hierarchical retrieval format groups Freebase relations by their dot-notation structure (e.g., domain → type → property), presenting them as a compact tree instead of a flat comma-separated list. In the example(see Table E.3.2), this formatting preserves the same information while reducing verbosity—79 tokens vs. 156 tokens ($\approx$49% fewer tokens).
Table 10: Hierarchical relation retrieval formatting used in KG-R1 to organize Freebase relations.
Flat Format Hierarchical Format
<kg-query> get_tail_relations(”m.09c7w0”) </kg-query> <information> Tail relations for ”m.09c7w0”: location.country.first_level_divisions, location.location.containedby, location.location.contains, people.person.nationality, people.person.place_of_birth, government.government.government_for </information> Result: 156 tokens <kg-query> get_tail_relations(”m.09c7w0”) </kg-query> <information> Tail relations for ”m.09c7w0”: location country: first_level_divisions location: containedby, contains people person: nationality, place_of_birth government government: government_for </information> Result: 79 tokens (49.4% savings)
E.3.3 PPO reward hacking
Table 11: Policy model reward hacking under PPO. The actor model generates fake KG information that critic model cannot distinguish from real data. Blue denotes responses generated by the KG-R1 agent.
Normal KG Retrieval Fake Information Generated by Actor Policy in PPO
Question: Who directed the movie Inception? <kg-query> get_head_entities(”film.film.directed_by”, ”Inception”) </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan (m.04f9z) </information> Answer: Christopher Nolan directed Inception. Question: Who directed the movie Inception? <kg-query> get_head_entities(”film.film.directed_by”, ”Inception”) </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan, Barack Obama, Elon Musk, Albert Einstein, Michael Jordan </information> Answer: Inception was directed by Christopher Nolan, with creative input from Barack Obama, Elon Musk, Albert Einstein, and Michael Jordan as co-directors.
Appendix F Examples of Responses
KG-R1 Response Example — WebQSP Question: What are Victoria Beckham’s kids’ names? Initial entities : ’Victoria Beckham’  <think>
I need to find out the names of Victoria Beckham’s kids. I will start by finding any relations to her children.
</think>
<kg-query>
$\mathtt{get\_tail\_relations}(\text{"Victoria Beckham"})$
</kg-query>
  <information> Tail relations for entity ”Victoria Beckham”:
base.popstra.celebrity.breakup, award.award_nominee.award_nominations, …, people.person.children, …
</information>
  <think>
The relation people.person.children is relevant. I will use it to get the names of her children.
</think>
<kg-query>
$\mathtt{get\_tail\_entities}(\text{"Victoria Beckham"},\text{"people.person.children"})$
</kg-query>
  <information> Tail entities for relation ”people.person.children”:
Brooklyn Joseph Beckham, Cruz David Beckham, Harper Seven Beckham, Romeo James Beckham
</information>
  <think>
The names of Victoria Beckham’s children are found. I will verify by checking each child’s parent relation.
</think>
<kg-query>
$\mathtt{get\_head\_relations}(\text{"Brooklyn Joseph Beckham"})$
</kg-query>
  <information> Head relations for ”Brooklyn Joseph Beckham”:
base.godparents.godparent.godchildren, people.person.children, people.sibling_relationship.sibling
</information>
  <think>
Confirming Brooklyn’s parents using people.person.children.
</think>
<kg-query>
$\mathtt{get\_head\_entities}(\text{"Brooklyn Joseph Beckham"},\text{"people.person.children"})$
</kg-query>
  <information> Head entities for relation with tail ”Brooklyn Joseph Beckham”:
David Beckham, Victoria Beckham
</information>
  <think> Brooklyn, Cruz, Harper, and Romeo all confirm to have David and Victoria Beckham as parents. </think> <answer>
Brooklyn Joseph Beckham, Cruz David Beckham, Harper Seven Beckham, Romeo James Beckham
</answer>
   [ground_truth] {‘Brooklyn Joseph Beckham’, ‘Romeo James Beckham’, ‘Cruz David Beckham’, ‘Harper Seven Beckham’}
Figure 10: Example KG-R1 response on WebQSP, showing multi-step reasoning and verification for a person-children query. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example - CWQ Question: What movie, written by Mike Jefferies, starred Nick Cannon?
Initial entities: ”Mike Jefferies”, ”Nick Cannon”
  <think>
Let’s start by finding a movie written by Mike Jefferies. We can use the relation that associates writers with their movies.
</think>
<kg-query>
$\mathtt{get\_tail\_relations}$(”Mike Jefferies”)
</kg-query>
  <information>Tail relations for entity ”Mike Jefferies”:
film.film_story_contributor.film_story_credits, people.person.nationality, people.person.place_of_birth, film.writer.film, …
</information>
  <think>
From the initial query results, we have two possible options for Mike Jefferies’ work. Let’s start with ”film.writer.film” to find the movies written by Mike Jefferies and then see if Nick Cannon starred in any of those films.
</think>
<kg-query>
$\mathtt{get\_tail\_entities}$(”Mike Jefferies”, ”film.writer.film”)
</kg-query>
  <information>Tail entities for relation ”film.writer.film” with head ”Mike Jefferies”: Goal II: Living the Dream, Goal III: Taking on the World</information>  <think>
Now that we have two movies written by Mike Jefferies, let’s check if Nick Cannon has starred in any of these movies. We’$l$ use the relation that connects actors with the movies they star in.
</think> <kg-query>
$\mathtt{get\_head\_relations}$(”Nick Cannon”)
</kg-query>
  <information>Head relations for entity ”Nick Cannon”: base.popstra.dated.participant, film.personal_film_appearance.person, music.album.primary_release,
base.popstra.lived_with.participant, film.film.written_by, …
</information>
  <think>
From the relations where ”Goal II: Living the Dream” is involved, we can see that it is a film with a writer, which aligns with Mike Jefferies.
</think>
<answer>
Goal II: Living the Dream, School Dance, School Gyrls, Goal III: Taking on the World
</answer>
   [ground_truth] {’Goal II: Living the Dream’ }
Figure 11: Example KG-R1 response in CWQ. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example - SimpleQA Question: where is the madam satan located? (Initial entities: ”Madam Satan”)
  <think>
I need to find out where Madam Satan is located. First, I will check if it has a location or if it’s an entity that can be queried about its location.
</think>
<kg-query>
$\mathtt{get\_tail\_relations}(\text{"Madam Satan"})$
</kg-query>
  <information>Tail relations for entity ”Madam Satan”:
film.film.cinematography, common.topic.notable_types, film.film.language, film.film.starring,film.film.country, … award.ranked_item.appears_in_ranked_lists, film.film.film_art_direction_by
</information>
  <think>
None of the identified relations are directly related to location. Let me check if Madam Satan is detailed in a geographical context.
</think>
<kg-query>
$\mathtt{get\_tail\_entities}(\text{"Madam Satan"},\text{"film.film.country"})$
</kg-query>
  <information>Tail entities for relation ”film.film.country” with head ”Madam Satan”: United States of America</information>  <think>
I have found that Madam Satan, identified as a film, is located in the United States of America. However, it’s crucial to confirm the specific location or city within the country where Madam Satan was filmed.
</think>
<kg-query>
$\mathtt{get\_tail\_entities}(\text{"Madam Satan"},\text{"film.film.starring"})$
</kg-query>
  <information>Tail entities for relation ”film.film.runtime” with head ”Madam Satan”: m.03kx9f2</information>  <think>
I am still not obtaining the most helpful information about the location. Possibly Madam Satan was released, but not filmed in a specific place.
</think>
<kg-query>
$\mathtt{get\_head\_relations}(\text{"m.02tb01n"})$
</kg-query>
  <information>Head relations for entity ”m.02tb01n”:
film.film.starring, film.actor.film
</information>
  <think>
I have questionned the relation ”film.actor.film”. This relation includes the movies the actor is in. Let’s try to find a clearer detail for the location of Madam Satan.
</think>
<answer>
United States of America
</answer>
   [ground_truth] {’United States of America’ }
Figure 12: Example KG-R1 response in SimpleQA. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example — T-REx Question: What is the instance of of Solberg??
Initial entities: ”Solberg”
  <think>
Let’s start with the entity Solberg and try to find out what it represents. It seems like Solberg could be an entity that describes a geopolitical boundary, location, or characterization. The query can be adjusted from general to specific as needed.
</think>
<kg-query>
$\mathtt{get\_tail\_relations}(\text{"Solberg"})$
</kg-query>
  <information>Tail relations for entity ”Solberg”: P1889, P407, P1552, P282, P31, P460</information>  <think>
The relations provided show how Solberg is related to other entities. Let’s explore each relation to find out more about Solberg.
</think>
<kg-query>
$\mathtt{get\_tail\_entities}(\text{"Solberg"},\ \text{"P282"})$
</kg-query>
  <information>Tail entities for relation ”P282” with head ”Solberg”: Latin script</information>  <think>
The relation P282 corresponds to ”type.object.name” according to the API documentation, indicating that Solberg is a human name. The relation P407 is also noteworthy but does not give further specific information. Let’s try a broader query for more context related to Solberg.
</think>
<kg-query>
$\mathtt{get\_tail\_entities}(\text{"Solberg"},\ \text{"P31"})$
</kg-query>
  <information>Tail entities for relation ”P31” with head ”Solberg”: Wikimedia disambiguation page, family name</information>
  <think>
The retrieval of ”Soelberg”, ”Sohlberg”, and ”Sohlberg” suggests that Solberg might be associated with these names. To find an instance of Solberg, we can search for a specific entity of this name.
</think>
<answer>
Latin script, Wikimedia disambiguation page, family name, Soelberg, Sohlberg, Sohlberg
</answer>
   [ground_truth] {’family name’, ’human settlement’, ’urban area in Sweden’, ’minor locality in Sweden’, ’minor locality in Sweden’, ’Wikimedia disambiguation page’, ’human settlement’, ’Ortsteil’, ’stone setting’, ’human settlement’, ’minor locality in Sweden’, ’hill’, ’stone setting’, ’hill’, ’human settlement’ }
Figure 13: Example KG-R1 response in T-REx. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example - QALD10en Question: After whom is the Riemannian geometry named?? (Initial entities: ”Riemannian geometry”)
  <think>
I need to find out who the Riemannian geometry is named after. Let’s start by investigating the relations and entities connected to ”Riemannian geometry”.
</think>
<kg-query>
$\mathtt{get\_tail\_relations}(\text{"Riemannian geometry"})$
</kg-query>
  <information>Tail relations for entity ”Riemannian geometry”:
on focus list of Wikimedia project, named after, maintained by WikiProject, topic’s main category, instance of, subclass of
</information>
  <think>
From the relations extracted, it appears the Riemannian geometry is named after someone. I will now search for the person it’s named after.
</think>
<kg-query>
$\mathtt{get\_tail\_entities}(\text{"Riemannian geometry"},\text{"named after"})$
</kg-query>
  <information>Tail entities for relation ”named after” with head ”Riemannian geometry”: Bernhard Riemann</information>  <think>
I now have the information that the Riemannian geometry is named after Bernhard Riemann. Let’s confirm this by checking if there’s more detailed information about this relationship.
</think>
<kg-query>
$\mathtt{get\_head\_relations}(\text{"Bernhard Riemann"})$
</kg-query>
  <information>Head relations for entity ”Bernhard Riemann”:
named after
</information>
  <think>
These results are confirming that Bernhard Riemann is indeed the person the Riemannian geometry is named after. No further information is needed.
</think>
<kg-query>
$\mathtt{get\_head\_entities}(\text{"Bernhard Riemann"},\text{"named after"})$
</kg-query>
  <information>Head entities for relation ”named after” with tail ”Bernhard Riemann”: Riemannian geometry</information>  <think>
This confirms my findings; the Riemannian geometry is named after Bernhard Riemann.
</think>
<answer>
Bernhard Riemann
</answer>
   [ground_truth] {’Bernhard Riemann’}
Figure 14: Example KG-R1 response in QALD10en. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example — GrailQA Question: what video game is published by softdisk?
Initial entities: ”Near West Side”
  <think>
First, I need to find out what game is published by Softdisk. Without specific knowledge about ”Near West Side”, I will start with the game publisher for Softdisk.
</think>
<kg-query>
$\mathtt{get\_head\_relations}(\text{"Near West Side"})$
</kg-query>
  <information>Head relations for entity ”Near West Side”:
book.periodical_publisher_period.publisher, business.industry.companies, business.employment_tenure.company
</information>
  <think>
The head relation ”business.employment_tenure.company” suggests that ”Near West Side” might be a game company or related to a game. I will try to find out the publisher of a game related to ”Near West Side” by looking for games published by Softdisk.
</think>
<kg-query>
$\mathtt{get\_tail\_relations}(\text{"Near West Side"})$
</kg-query>
  <information>Tail relations for entity ”Near West Side”:
organization.organization.place_founded, organization.organization.headquarters, book.periodical_publisher.periodicals_published, business.employer.employees, cvg.cvg_publisher.games_published, business.business_operation.industry, common.topic.notable_types
</information>
  <think>
The previous query did not provide a conclusive answer. Let’s try a different approach by checking if ”Augusto Cury” is associated with any game publishers.
</think>
<answer>
Augusto Cury, Catacomb 3-D, Chagunitzu, Dangerous Dave in the Haunted Mansion, Hovertank 3D, Iron Seed, Lado Skorin, North Cambridge Academy, Roberto Torretti, Shadow Knights, Star Film Company, Tiles of the Dragon
</answer>
   [ground_truth] {’Rescue Rover’,’Catacomb 3-D’,’Shadow Knights’,”Dangerous Dave’s Risky Rescue”,’Tiles of the Dragon’, ’Catacomb’, ’Commander Keen: Keen Dreams’, ’Dave Goes Nutz!’,’Chagunitzu’,’Dangerous Dave in the Haunted Mansion’,’Hovertank 3D’,’Iron Seed’}
Figure 15: Example KG-R1 response in GrailQA. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example — MultiTQ Question: Who negotiated with Malaysia in 2005–10?  <think>
To answer this, I should check which relations connect to Malaysia that indicate negotiations, then look for counterpart entities within the 2005–10 range.
</think>
<kg-query>
$\mathtt{get\_head\_relations}(\text{"Malaysia"})$
</kg-query>
  <information> Head relations for entity ”Malaysia”:
Appeal for aid; Appeal for diplomatic cooperation; Appeal for humanitarian aid; … Engage in negotiation; Express intent to settle dispute
</information>
  <think>
”Engage in negotiation” appears among head relations. Next, retrieve the counterpart entities that negotiated with Malaysia.
</think>
<kg-query>
$\mathtt{get\_head\_entities}(\text{"Malaysia"},\text{"Engage in negotiation"})$
</kg-query>
  <information> Head entities connected by (?, Engage in negotiation, Malaysia):
Cambodia [2008-06]; China [2011-11]; Governor (Malaysia) [2007-05]; Iran [2006-05; 2006-07]; Japan [2005-05; 2005-06; 2009-01; 2011-10]; Julia Gillard [2011-07]; Mwai Kibaki [2007-04]; … Yemen [2005-02; 2008-10]
</information>
  <think>
We need the counterpart who negotiated with Malaysia specifically within 2005–10. Signals suggest Thailand-related negotiations in this window; inspect Thailand’s negotiation partners.
</think>
<kg-query>
$\mathtt{get\_tail\_relations}(\text{"Thailand"})$
</kg-query>
  <information> Tail relations for entity ”Thailand”:
Accede to demands for rights; Accuse; Appeal for aid; Appeal for de-escalation of military engagement; … Engage in negotiation
</information>
  <think>
Confirm the negotiation counterparts of Thailand to identify the specific actor matching the 2005–10 window.
</think>
<kg-query>
$\mathtt{get\_tail\_entities}(\text{"Thailand"},\text{"Engage in negotiation"})$
</kg-query>
  <information> Tail entities connected by (Thailand, Engage in negotiation, ?):
Abhisit Vejjajiva [2009-06; 2009-10; 2010-01]; Cambodia [2005-04; 2005-06; 2008-01; …; 2012-07]; China [2006-06; 2006-09; 2010-03; 2011-01]; Donald Rumsfeld [2005-06]; Koichiro Matsuura [2011-02]; … South Korea [2005-12]
</information>
  <think>
Within 2005–10, the counterpart that negotiated with Malaysia is Surakiart Sathirathai.
</think>
<answer>
Surakiart Sathirathai
</answer>
   [ground_truth] {’Surakiart Sathirathai’}
Figure 16: Example KG-R1 response on MultiTQ, illustrating all four query functions and their usage. Blue denotes responses generated by the KG-R1 agent.

### v5

@@
 Abstract
-Knowledge-graph retrieval-augmented generation (KG-RAG) couples large language models (LLMs) with structured, verifiable knowledge graphs (KGs) to reduce hallucinations and expose reasoning traces.
-However, many KG-RAG systems compose multiple LLM modules (e.g planning, reasoning, and responding), inflating inference cost and binding behavior to a specific target KG.
-To address this, we introduce KG-R1, an agentic KG retrieval-augmented generation (KG-RAG) framework through reinforcement learning (RL).
-KG-R1 utilizes a single agent that interacts with KGs as its environment, learning to retrieve at each step and incorporating the retrieved information into its reasoning and generation.
-The process is optimized through end-to-end RL.
-In controlled experiments across Knowledge-Graph Question Answering(KGQA) benchmarks, our method demonstrates both efficiency and transferability: Using Qwen-2.5-3B, KG-R1 improves answer accuracy with fewer generation tokens than prior multi-module workflow methods that use larger foundation or fine-tuned models.
-Furthermore, KG-R1 enables plug and play: after training, it maintains strong accuracy on new KGs without modification.
+Knowledge-graph retrieval-augmented generation (KG-RAG) couples large language models (LLMs) with structured, verifiable knowledge graphs (KGs) to reduce hallucination and provide reasoning traces.
+However, current KG-RAG systems often rely on fixed pipelines of multiple LLM modules (e.g., planning, reasoning, and responding), which inflate inference costs and tie performance to specific graph schemas.
+To address this, we introduce KG-R1, an agentic framework that optimizes KG-RAG through reinforcement learning (RL).
+Unlike modular workflows, KG-R1 uses a single agent that interacts with KGs as its environment, learning to retrieve information at each step and incorporating it into its reasoning and generation in a unified process.
+Across Knowledge-Graph Question Answering (KGQA) benchmarks, KG-R1 demonstrates both efficiency and transferability—using Qwen 2.5-3B, KG-R1 improves answer accuracy with fewer generation tokens than prior multi-module workflow methods that use much larger foundation or fine-tuned models.
+Furthermore, KG-R1 exhibits strong plug-and-play capability: after training, maintaining accuracy on unseen KGs without retraining.
 These properties make KG-R1 a promising KG-RAG framework for real-world deployment.
-Our code is publicly available at  https://github.com/Jinyeop3110/KG-R1.
+Our code is publicly available at  github.com/junhongmit/KG-R1/.
 1 Introduction
-22footnotetext: Work done during an internship at MIT-IBM Watson AI Lab, IBM Research.
-Figure 1: Overview of KG-R1, a multi-turn agentic framework for KG-RAG trained with reinforcement learning.
-The framework enables cost-efficient inference and demonstrates strong cross-KG transferability.
-Retrieval-augmented generation (RAG) has gained popularity as a way to enhance large language models (LLMs) with access to external knowledge, thereby reducing hallucinations and improving accuracy in knowledge-intensive tasks.
+Retrieval-augmented generation (RAG) [gao2023retrieval] has gained popularity as a way to enhance large language models (LLMs) with access to external knowledge, thereby reducing hallucination and improving accuracy in knowledge-intensive tasks.
 Recent research has extended this idea to knowledge graph retrieval-augmented generation (KG-RAG), where knowledge graphs (KGs) are leveraged as the retrieval source.
 A KG is a structured representation of knowledge in the form of entities (nodes) and their relationships (edges) that encodes factual knowledge in a graph format.
-Augmenting LLMs with KGs has proven effective not only in mitigating the knowledge bottleneck, but also in improving reasoning performance over complex multi-hop relations and enhancing adaptation to continually evolving real-world information (Sun et al., 2024; Luo et al., 2024; Wang et al., 2025c).
-These properties make KG-RAG especially promising in high-stakes domains, such as medical consultation and legal analysis (Xiong et al., 2023; Cui et al., 2024).
-As shown in Figure 1, typical KG-RAG adopts a modular workflow consisting of four primary subtasks: Retrieval to query facts from KGs, Reasoning to process the retrieved information, Reviewing to verify logical consistency, and Responding to synthesize the final answer (Baek et al., 2023; Sun et al., 2024; Luo et al., 2024; Wang et al., 2025c).
-Each distinct subtask is handled by specialized LLM modules through two main methods: (i) prompt-based with task-specific instructions, often including in-context demonstrations (Baek et al., 2023); and (ii) fine-tuned modules tailored to particular tasks (e.g., SPARQL generation (D’Abramo et al., 2025; Lee & Shin, 2024) or relation extraction (Yao et al., 2019)) on specific KGs.
+Augmenting LLMs with KGs has proven effective not only in mitigating the knowledge bottleneck but also in improving reasoning performance over complex multi-hop relations and enhancing adaptation to continually evolving real-world information [23, 16, 28].
+These properties make KG-RAG especially promising in high-stakes domains, such as medical consultation and legal analysis [33, 6].
+As shown in Figure˜1, a typical KG-RAG adopts a modular workflow consisting of four subtasks: retrieval to query facts from KGs, reasoning to process retrieved information, reviewing to verify logical consistency, and responding to synthesize the answer [1, 23, 16, 28].
+Each subtask is handled by specialized LLM modules through two main methods: (i) prompt-based modules with task-specific instructions, often including in-context demonstrations [1]; and (ii) fine-tuned modules tailored to specific tasks (e.g., SPARQL generation [7, 15] or relation extraction [30]) on specific KGs.
+Figure 1: Overview of KG-R1, a multi-turn agentic KG-RAG framework. The framework enables cost-efficient inference and demonstrates strong cross-KG transferability.
+Despite improved reasoning accuracy, the real-world deployment of such workflows faces two key challenges: achieving high reasoning effectiveness while maintaining efficiency, and generalization to new or updated KGs.
+First, prompt-based methods that repeatedly call large foundation LLMs accumulate inference across stages and drive up latency, token usage, and compute (e.g., ToG [23] and ReKnoS [28]; see left panel of Figure˜2).
+Second, these prompted or fine-tuned modules are typically tuned to a particular KG’s domain and schema (entity types, relations, constraints), often via curated in-context examples or KG-specific fine-tuning.
+As a result, performance does not transfer reliably when the domain shifts, the schema changes, or the system is deployed on a new KG (e.g., RoG [16]; see right panel of Figure˜2).
 Figure 2: Prior multi-module methods are costly and do not transfer well across KGs.
-Left: mean end-to-end generated tokens per query on WebQSP (Yih et al., 2016).
-Right: average $\mathrm{Hit@1}$ over five out-of-training KGQA datasets (See Sec.
-4.2).
+Left: mean end-to-end generated tokens per query on WebQSP [31].
+Right: average Hits@1 over five out-of-training KGQA datasets (see Section˜4.4).
 KG-R1 achieves both low token cost and strong cross-KG transferability.
-Despite improved reasoning accuracy, the real-world deployment of such workflows faces two key challenges, high computational cost and lack of generalization to new or updated KGs.
-First, prompt-based methods that repeatedly call large foundation LLMs accumulate inference across stages and drive up latency, token usage, and compute (e.g., ToG and ReKnoS; see left panel of Fig.
-2).
-Second, these prompted or fine-tuned modules are typically tuned to a particular KG’s domain and schema (entity types, relations, constraints), often via curated in-context examples or KG-specific fine-tuning.
-As a result, performance does not transfer reliably when the domain shifts, the schema changes, or the system is deployed on a new KG (e.g., RoG; see right panel of Fig.
-2).
-To tackle these challenges, we introduce KG-R1, an agentic KG-RAG system that employs end-to-end multi-turn reinforcement learning (Jin et al., 2025; Zeng et al., 2025; DeepSeek-AI, 2025).
-As shown in Figure 1, the architecture of KG-R1 has two components: a single LLM agent and a KG retrieval server (as an environment).
-The KG retrieval server hosts the knowledge graph along with a set of retrieval actions.
+To tackle these challenges, we introduce KG-R1, an agentic KG-RAG system that employs end-to-end multi-turn reinforcement learning [12, 32, 8] to learn effective and transferable KG interaction policies.
+As shown in Figure˜1, the architecture of KG-R1 has two components: a single LLM agent and a KG query server.
+The KG query server hosts the knowledge graph along with a set of retrieval actions.
 The LLM agent iteratively performs cycles of short reasoning followed by retrieval actions over multiple turns, with each decision informed by knowledge obtained from the KG retrieval server, and generates a final answer.
-Figure 2 demonstrates that KG-R1 achieves both high efficiency and strong cross-KG transferability simultaneously using a 3B-parameter model, outperforming prior methods.
+By learning a unified interaction policy, KG-R1 avoids redundant multi-module coordination and allocates computation adaptively across reasoning steps, leading to both improved effectiveness and reduced token usage.
+Figure˜2 demonstrates that KG-R1 achieves both high efficiency and strong cross-KG transferability simultaneously using a 3B-parameter model, outperforming prior methods.
 Our key contributions are summarized as follows:
@@
 KG-R1 Framework.
-We introduce an agentic KG-RAG system (Section 3) that replaces multi-module pipelines with a single agent for KG-RAG, running against a lightweight KG server.
+We introduce an agentic KG-RAG system (Section 3) that replaces multi-module pipelines with a single agent for KG-RAG, running on a lightweight KG server.
 The agent alternates between reasoning and retrieval actions over multiple turns, with the end-to-end trajectory optimized by RL using both turn-wise and outcome-based reward signals.
@@
 2.
-Efficient inference.
+Effective and Efficient Inference.
 By consolidating reasoning and retrieval into a single-agent, near-continuous workflow, KG-R1 achieves competitive reasoning accuracy with a small-parameter model while reducing token usage.
@@
 Plug-and-play Transferability across Diverse KGs.
-KG-R1 easily transfers to diverse KGs and maintains strong KG-RAG performance (Section 4.2).
-The trained KG-R1 agent generalizes to new KGs without modification—backend KGs can be swapped without changing prompts or hyper-parameters, and without fine-tuning.
-This enables zero-shot transfer to unseen knowledge graphs.
+KG-R1 can be easily transferred to diverse KGs and maintains strong KG-RAG performance (Section 4.4).
+The trained KG-R1 agent generalizes to new KGs without retraining: backend KGs can be swapped without changing prompts, hyperparameters, or fine-tuning.
+This generalizability enables zero-shot transfer to unseen KGs.
 2 Related Work
-KG-RAG.
-Knowledge Graph Retrieval-Augmented Generation (KG-RAG) augments LLMs with structured knowledge graphs (KGs) to improve factuality and compositional reasoning.
-Early work grounds LLMs’ generation by translating natural-language questions into executable graph queries (e.g., SPARQL/Cypher), retrieving relevant subgraphs or answers, and feeding them back to the model (Ouyang et al., 2022; Izacard et al., 2023; Lee & Shin, 2024).
-More recent approaches adopt a modular LLM pipeline over KGs, interleaving natural-language reasoning with multi-stage planning, path search, and execution, where each stage uses a prompted or fine-tuned LLM (Luo et al., 2024; Sun et al., 2024; Wang et al., 2025c).
+Knowledge Graph Retrieval-Augmented Generation (KG-RAG).
+KG-RAG augments LLMs with structured KGs to improve quality and compositional reasoning.
+Early work grounds LLMs’ generation by translating natural-language questions into executable graph queries (e.g., SPARQL/Cypher), retrieving relevant subgraphs or answers, and feeding them back to the model [18, 15].
+More recent approaches adopt a modular LLM pipeline over KGs, interleaving natural language reasoning with multi-stage planning, path search, and execution, where each stage uses a prompted or fine-tuned LLM [16, 23, 4, 28].
 Despite these advances, most systems rely on hand-engineered modules or prompt heuristics tied to a specific KG schema, which induce cost inefficiency and limit generalization.
 These challenges motivate our KG-R1 framework: a single-agent KG-RAG approach that improves efficiency and transfers well to new KGs.
-Multi-turn RL for LLM.
-Reinforcement learning (RL) has become central to equipping LLMs with step-by-step (chain-of-thought) reasoning behavior (OpenAI et al., 2024; DeepSeek-AI, 2025).
-RL-enhanced models yield substantial gains in math and coding (Le et al., 2022; Chervonyi et al., 2025), and broader complex reasoning tasks.
-More recently, RL has been applied to agentic LLMs that invoke external tools (e.g., bash terminals and APIs) or interact with knowledge bases, improving tool use (Qin et al., 2024) and retrieval-augmented generation (RAG) (Jin et al., 2025; Wang et al., 2025a) to facilitate the tool use or RAG.
-Building on these advances, our work, KG-R1 adopts end-to-end RL as the core optimization algorithm for agentic KG-RAG framework.
+Multi-turn RL for LLMs.
+Reinforcement learning (RL) has become central to equipping LLMs with step-by-step (chain-of-thought) reasoning behavior [17, 8].
+RL-enhanced models yield substantial gains in math and coding [14, 5], and broader complex reasoning tasks.
+More recently, RL has been applied to agentic LLMs that invoke external tools (e.g., bash terminals and APIs) or interact with knowledge bases, improving tool use [20] and retrieval-augmented generation (RAG) [12, 26] to facilitate tool use or RAG.
+Building on these advances, our work, KG-R1 adopts end-to-end RL as the core optimization algorithm for an agentic KG-RAG framework.
 3 KG-R1: An Agentic KG-RAG Framework
 3.1 Problem Definition
-KG–RAG spans many applications, including conversational assistants (Chaudhuri et al., 2021), recommendation systems (Wang et al., 2025b), and open-domain QA (Zhu et al., 2025).
-In this work, we instantiate and evaluate our approach on Knowledge Graph Question Answering (KGQA), which provides a grounded testbed for KG-RAG: ground-truth answers are tied to a fixed KG, evaluation is verifiable and intermediate graph reasoning is interpretable.
+KG-RAG spans many applications, including conversational assistants [3], recommendation systems [27], and open-domain QA [35].
+In this work, we instantiate and evaluate our approach on Knowledge Graph Question Answering (KGQA), which provides a grounded testbed for KG-RAG: ground-truth answers are tied to a fixed KG, evaluation is verifiable, and intermediate graph reasoning is interpretable.
 We now formalize knowledge graphs and KGQA tasks.
@@
 Consider a dataset $D=\{(q,G,A_{q})\}$, where $q$ is a natural language question, $G$ is a KG, and $A_{q}\subseteq\mathcal{E}$ is the ground-truth answer set paired with the question.
-For each gold answer $e^{*}\in A_{q}$, there exist one or more reasoning paths in $G$ that connect anchor entities mentioned in $q$ to $e^{*}$.
-A reasoning path is a sequence of labeled edges $r_{1},\dots,r_{\ell}\in\mathcal{R}$ instantiated by entities $e_{0},\dots,e_{\ell}$ such that $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell},$ where $e_{0}$ is an anchor entity mentioned in $q$ and $e_{\ell}=e^{*}\in A_{q}$, and $\mathcal{Z}(q,G,A_{q})$ denote the set of all valid reasoning paths for question $q$ over $G$ with respect to ground-truth answers $A_{q}$.
+For each ground truth answer $e^{*}\in A_{q}$, there exists one or more reasoning paths in $G$ that connect anchor entities mentioned in $q$ to $e^{*}$.
+A reasoning path is a sequence of labeled edges $r_{1},\dots,r_{\ell}\in\mathcal{R}$ instantiated by entities $e_{0},\dots,e_{\ell}$ such that $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell},$ where $e_{0}$ is an anchor entity mentioned in $q$ and $e_{\ell}=e^{*}\in A_{q}$, and $\mathcal{Z}(q,G,A_{q})$ denotes the set of all valid reasoning paths for question $q$ over $G$ with respect to ground-truth answers $A_{q}$.
 In solving KGQA, given $q$ and $G$, a model attempts to discover a subset of valid reasoning paths from $\mathcal{Z}(q,G,A_{q})$ and predicts $\hat{A}_{q}$ based on the terminal entities of the discovered paths.
-The model’s performance is evaluated by comparing $\hat{A}_{q}$ with the ground-truth $A_{q}$.
+The model’s performance is evaluated by comparing $\hat{A}_{q}$ with the ground truth $A_{q}$.
 As an example, to answer the question “What is the capital of the U.S.
 state whose largest city is Chicago?”, a valid reasoning path is
-$\text{Chicago}\xrightarrow{\texttt{located\_in\_state}}\text{Illinois}\xrightarrow{\texttt{capital}}\text{Springfield}$
+$\text{Chicago}\xrightarrow{\texttt{located\_in\_state}}\text{Illinois}\xrightarrow{\texttt{capital\_of}}\text{Springfield}.$
 This path belongs to $\mathcal{Z}(q,G,A_{q})$ and leads to the correct prediction $\hat{A}_{q}=\{\text{Springfield}\}$.
 3.2 KG-R1 Framework
-Figure 3: KG-R1 framework: a single LLM agent undergoes multi-turn generation–execution loop with a schema-agnostic KG retrieval server and responds with the final answer.
-KG-R1 casts KG-RAG as a multi-turn interaction with a KG interface (KG retrieval server).
+Figure 3: KG-R1 framework: a single LLM agent undergoes a multi-turn generation-execution loop with a schema-agnostic KG retrieval server and responds with the final answer.
+KG-R1 casts KG-RAG as a multi-turn interaction with a KG interface (KG query server).
 We prioritize two design principles.
 First, we design a single-agent architecture that simplifies deployment and enables efficient, low-cost inference.
-Second, we build a schema-agnostic KG retrieval server that avoids KG-specific assumptions and remains portable across varied KGs.
+Second, we build a schema-agnostic KG retrieval server that avoids KG-specific assumptions and remains portable across different KGs.
 Given a KGQA dataset $D=\{(q,G,A_{q})\}$, we set up a KG retrieval server and a KG-R1 agent.
-KG retrieval server. The server hosts the knowledge graph $G$ and provides a set of retrieval actions $a\in\mathcal{U}_{\text{ret}}$ that enable graph traversal through 1-hop operations:
+KG Query Server. The server hosts the knowledge graph $G$ and provides a set of retrieval actions $a\in\mathcal{U}_{\text{ret}}$ that enable graph traversal through 1-hop operations:
 […]
@@
 Forward traversal along $Z$ is implemented by $\mathtt{get\_tail\_entities}(e_{i-1},r_{i})$ for $i=1,\dots,\ell$ and backward traversal is implemented by $\mathtt{get\_head\_entities}(r_{i},e_{i})$.
-Notably, this choice of the retrieval action set guarantees completeness and schema-free transferability, as formalized by propositions 3.1 and 3.2, whose proofs are provided in Appendix A.1.
-Proposition 3.1 (Retrieval Action Set Completeness).
-For any reasoning path $Z:\ e_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{\ell}}e_{\ell}$ in $G$, there exists an action sequence in $\mathcal{U}_{\text{ret}}$ of length at most $\ell{+}1$ whose output includes $e_{\ell}$.
-Proposition 3.2 (Schema-Free Transferability).
-The semantics of $\mathcal{U}_{\text{ret}}$ depend only on directed triples $(e,r,e^{\prime})$, so any fixed client policy transfers across heterogeneous directed KGs by replacing $G$ without redefining $\mathcal{U}_{\text{ret}}$.
+This action design guarantees completeness of path realization, as formalized in Remark 3, and provides a schema-agnostic interaction interface that enables transfer across KGs.
+Full formal statements and additional properties are provided in Appendix B.
+{restatable}
+[Retrieval Action Set Completeness]remarkcompleteness For any reasoning path $Z:\ e_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{\ell}}e_{\ell}$ in $G$, there exists an action sequence in $\mathcal{U}_{\text{ret}}$ of length at most $\ell{+}1$ whose output includes $e_{\ell}$.
 KG-R1 Agent.
-We model a single LLM agent that interacts with a KG retrieval server in a multi-turn setting.
+We model a single LLM agent that interacts with a KG query server in a multi-turn setting.
 The agent undergoes initialization followed by a loop of generation and execution (see Figure 3).
-At initialization, a lightweight base LLM is configured with an instruction prompt $p$ (see the box below for the prompt template) containing general reasoning instructions, the user question $q$, and the KG retrieval server instructions (Table D.2).
+At initialization, a lightweight base LLM is configured with an instruction prompt $p$ (see the box below for the prompt template) containing general reasoning instructions, the user question $q$, and the KG retrieval server instructions (Table E.2).
 At each turn $t\leq H$, where $H$ is the maximum turn limit, the agent first undergoes the generation phase where it produces a response $m_{t}$ comprising two parts: (1) an internal reasoning wrapped in <think>...</think>, and (2) an action, which is either a KG retrieval action wrapped in <kg-query>...</kg-query>, or a final answer wrapped in <answer>...</answer>.
@@
 If $a_{t}=\operatorname{Answer}$, the interaction terminates: the content inside <answer>...</answer> is extracted and post-processed (normalization, deduplication, and entity resolution) to produce the predicted answer set $\hat{A}_{q}$.
-Given interaction trajectory $\tau=(p,\,(m_{1},o_{1}),\ldots,(m_{t},o_{t}))$ with $t\leq H$, the KG-R1 agent’s policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ is defined over token sequences $m_{t}$ and governs how textual responses are generated.
+Given an interaction trajectory $\tau=(p,\,(m_{1},o_{1}),\ldots,(m_{t},o_{t}))$ with $t\leq H$, the KG-R1 agent’s policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ is defined over token sequences $m_{t}$ and governs how textual responses are generated.
 3.3 KG-R1 Training with RL
-Our goal is to find the KG-R1 agent’s policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ that effectively retreive reasoning paths on $G$ via multi-turn interaction and generate the accurate answer $\hat{A}_{q}=A_{q}$.
-To this end, we optimize the base LLM with reinforcement learning, using a GRPO-style objective (DeepSeek-AI, 2025) in a multi-turn setting (Qian et al., 2025; Jin et al., 2025; Zeng et al., 2025).
-Our reward function combines action-level and outcome-based signals.
-The overview of the training procedure refers to Algorithm 1 (Appendix).
+Our goal is to learn a policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ for the KG-R1 agent that interacts with the knowledge graph $G$ over multiple turns to incrementally retrieve informative reasoning paths and ultimately produce the correct answer $\hat{A}_{q}=A_{q}$.
+We optimize this policy using reinforcement learning with a GRPO-style objective [8], following recent advances in multi-turn agent optimization [19, 12, 32].
+An overview of the training procedure is provided in Algorithm 1 in the Appendix˜E.
 Reward Objective.
 To effectively train the KG-R1 agent, we combine verifiable turn rewards with outcome-level global rewards.
-Turn rewards ($r^{\text{turn}}_{t}$) provide local signals at each step as the sum of three components: (i) format validity $v_{\mathrm{fmt}}(m_{t})$ checks that $m_{t}$ contains both reasoning and a well-formed action $a_{t}\in\mathcal{A}$; (ii) KG query $v_{\mathrm{kg}}(a_{t},o_{t+1})$ checks that executing $a_{t}$ yields meaningful, schema-valid retrieval in $o_{t+1}$; and (iii) answer $v_{\mathrm{ans}}(m_{T})$ checks the final answer’s format/consistency on the final turn.
-the turn rewards are computed as follows with weight $w_{fmt},w_{kg},w_{ans}$:
+Turn rewards ($r^{\text{turn}}_{t}$) provide local signals at each turn as the sum of three components: (i) format validity $v_{\mathrm{fmt}}(m_{t})$ checks that $m_{t}$ contains both reasoning and a well-formed action $a_{t}\in\mathcal{A}$; (ii) KG query $v_{\mathrm{kg}}(a_{t},o_{t+1})$ checks that executing $a_{t}$ yields meaningful, schema-valid retrieval in $o_{t+1}$; and (iii) answer consistency $v_{\mathrm{ans}}(m_{T})$ checks the final answer’s format/consistency on the final turn.
+The turn rewards are computed as follows with weights $w_{fmt}$, $w_{kg}$, and $w_{ans}$:
 $r^{\mathrm{turn}}_{t}=w_{\mathrm{fmt}}\,v_{\mathrm{fmt}}(m_{t})+w_{\mathrm{kg}}\,v_{\mathrm{kg}}(a_{t},o_{t+1})+w_{\mathrm{ans}}\,v_{\mathrm{ans}}(m_{T}).$ (1)
-Global rewards summarize trajectory-level outcomes as the sum of the following: (i) a final-answer F1 score over the predicted set $\hat{A}_{q}$ vs.
-ground-truth answer set $A_{q}$; and (ii) a retrieval score $v_{ret}$ that is $1$ if any gold entity appears anywhere in the retrieved information along the executed reasoning path, and 0 otherwise.
-$R^{\mathrm{global}}=w_{\mathrm{F1}}\cdot\mathrm{F1}\!\bigl(\hat{A}_{q},A_{q}\bigr)\;+\;w_{\mathrm{ret}}\cdot v_{ret}$ (2)
+Global rewards summarize trajectory-level outcomes as the sum of the following: (i) a binary answer correctness signal $\mathrm{H1}$ that is $1$ if the predicted answer $\hat{A}_{q}$ is included in the ground-truth answer set $A_{q}$, and 0 otherwise; and (ii) a retrieval score $v_{ret}$ that is $1$ if any ground-truth entity appears anywhere in the retrieved information along the executed reasoning path, and 0 otherwise.
+$R^{\mathrm{global}}=w_{\mathrm{H1}}\cdot\mathrm{H1}\!\bigl(\hat{A}_{q},A_{q}\bigr)\;+\;w_{\mathrm{ret}}\cdot v_{ret}$ (2)
 Group-relative Turn-level Credit Assignment and Optimization.
-To convert rewards into token-level credit, we use a group-relative, turn-level credit assignment inspired by Zeng et al.
-(2025).
+To assign credit across multi-turn trajectories, we adopt a group-relative, turn-level credit assignment strategy [32].
 We collect $N$ rollouts per query $q$.
-For each episode $\tau^{(n)}$, we compute turn rewards $r^{\mathrm{turn},(n)}_{t}$ and a global reward $R^{\mathrm{global},(n)}$, and then form turn-specific returns with $\lambda$ as the normalization factor:
+For each trajectory $\tau^{(n)}$, we compute turn rewards $r^{\mathrm{turn},(n)}_{t}$ and a global reward $R^{\mathrm{global},(n)}$, and combine them into a per-turn return with $\lambda$ as the normalization factor:
 $G_{t}^{(n)}\;=\;r^{\mathrm{turn},(n)}_{t}\;+\;\lambda\,R^{\mathrm{global},(n)}.$ (3)
@@
 The turn-level advantage $A_{t}^{n}$ is calculated using a single group baseline $\bar{G}$ that averages over $\mathcal{S}$:
-$A_{t}^{(n)}=\frac{G_{t}^{(n)}-\bar{G}}{\sigma_{G}+\varepsilon}\quad\text{where}\quad\bar{G}=\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}G_{t}^{(n)},\ \sigma_{G}=\sqrt{\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}\!\big(G_{t}^{(n)}-\bar{G}\big)^{2}}.$
-where $\varepsilon$ is a small constant for numerical stability.
-RL update. Using turn-level credit $A_{t}^{(n)}$, we optimize the agent’s policy $\pi_{\theta}$ with a GRPO-style objective $\mathcal{J}$ :
-$\mathcal{J}(\theta)=\mathbb{E}\!\Bigg[\sum_{n,t,i}\min\!\Big(\rho_{t,i}^{(n)}\,\widetilde{A}_{t,i}^{(n)},\ \operatorname{clip}\big(\rho_{t,i}^{(n)},\,1-\epsilon,\,1+\epsilon\big)\,\widetilde{A}_{t,i}^{(n)}\Big)\;-\;\beta\,\mathrm{KL}\!\big(\pi_{\theta}(\cdot\mid h_{t,i}^{(n)})\,\|\,\pi_{0}(\cdot\mid h_{t,i}^{(n)})\big)\Bigg]$
-with $\widetilde{A}_{t,i}^{(n)}=m_{t,i}^{(n)}A_{t}^{(n)}$.
-$\pi_{\theta}$ is the current policy, and $\pi_{\theta_{\text{old}}}$ is the behavior policy used for sampling; $\pi_{0}$ is a fixed reference policy used only for KL regularization (weight $\beta$).
-$\rho_{t,i}^{(n)}=\exp\!\big(\log\pi_{\theta}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})-\log\pi_{\theta_{\text{old}}}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})\big)$ is the importance ratio that corrects off-policy sampling at token $i$.
-We maximize $\mathcal{J}(\theta)$ by gradient ascent.
-The proposed group-based credit assignment per turn produces stabilized and effective signals in the multi-turn interaction decisions.
+$A_{t}^{(n)}=\frac{G_{t}^{(n)}-\bar{G}}{\sigma_{G}+\varepsilon}\quad\text{where}\quad\bar{G}=\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}G_{t}^{(n)},\ \sigma_{G}=\sqrt{\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}\!\big(G_{t}^{(n)}-\bar{G}\big)^{2}},$
+where $\varepsilon$ is a constant for numerical stability.
+Intuitively, this measures each turn’s contribution relative to the group average, reinforcing above-average decisions and discouraging below-average ones.
+RL update. Using the turn-level advantage $A_{t}^{(n)}$, we optimize the agent’s policy $\pi_{\theta}$ with a GRPO-style clipped objective $\mathcal{J}$:
+$\mathcal{J}(\theta)=\mathbb{E}\!\Bigg[\sum_{n,t,i}\Big(\min\!\Big(\rho_{t,i}^{(n)}\widetilde{A}_{t,i}^{(n)},\ \operatorname{clip}\big(\rho_{t,i}^{(n)},1-\epsilon,1+\epsilon\big)\,\widetilde{A}_{t,i}^{(n)}\Big)-\beta\mathrm{KL}\!\big(\pi_{\theta}(\cdot\mid h_{t,i}^{(n)})\|\pi_{0}(\cdot\mid h_{t,i}^{(n)})\big)\Big)\Bigg],$
+where $\widetilde{A}_{t,i}^{(n)}=m_{t,i}^{(n)}A_{t}^{(n)}$ applies the same turn-level advantage to all tokens within a turn via a mask $m_{t,i}^{(n)}$, $\operatorname{clip}$ bounds updates to improve stability, $n$, $t$, and $i$ index trajectories, turns, and tokens, respectively.
+The current policy is $\pi_{\theta}$, the behavior policy $\pi_{\theta_{\text{old}}}$ is used for sampling, and $\pi_{0}$ is a fixed reference policy used only for KL regularization (weight $\beta$).
+The importance ratio $\rho_{t,i}^{(n)}=\exp\!\big(\log\pi_{\theta}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})-\log\pi_{\theta_{\text{old}}}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})\big)$ corrects off-policy sampling at each token $i$.
+We maximize $\mathcal{J}(\theta)$ via gradient ascent.
+The proposed combination of structured rewards and group-relative credit assignment yields stable and informative learning signals for multi-turn decision-making, leading to improved reasoning quality and final answer correctness.
 We provide comprehensive ablation studies of each component of our reward design in Section 4.5.
 3.4 Cross-KG Transfer: Plug-and-Play
-We propose a plug-and-play approach for KG-R1 that enables KG–RAG to operate across different knowledge graphs without retraining.
+We propose a plug-and-play approach for KG-R1 that enables KG-RAG to operate across different knowledge graphs without retraining.
 Let $\pi_{\theta\mid\mathcal{D}}$ denote the policy of the KG-R1 agent trained on a KGQA dataset $\mathcal{D}=\{(q,G,A_{q})\}$.
@@
 4 Experiments
-We assess KG-R1 along two axes.
-(1) Performance and Efficiency on Trained KGs: We train KG-R1 on the training splits of KGQA benchmarks and measure answer quality and inference cost on their held-out test splits.
-(2) Cross-KG transferability: We test how well KG-R1 carries over to KGs out of training.
-In both parts, we compare against established baselines to support the KG-R1 results.
-4.1 Performance and efficiency on Trained KGs
-Models. We use Qwen2.5-3B-it (Qwen et al., 2025) as our base model. For the other baseline methods, we use GPT-4o-mini, GPT-3.5, and Llama2-7B-it, following the setups used in prior work.
+We evaluate KG-R1 through two research questions: RQ1 (Effectiveness and Efficiency): Can KG-R1 learn an effective and efficient KG-RAG policy when trained on a target KG?; RQ2 (Cross-KG Transferability): Does KG-R1 generalize to unseen KGs without additional training?
+For both research questions, we compare against established LLM-only and LLM+KG baselines to assess effectiveness, efficiency, and transferability.
+4.1 Experimental Setup
+Models.
+We use Qwen2.5-3B-it [29] as our base model.
+For the other baseline methods, we use Qwen2.5-7B-it, Qwen2.5-72B-it, Qwen3-235B-A22B-Instruct-2507 [yang2025qwen3], and Llama2-7B-it (only for the fine-tuned planning in RoG [16]), following the setups used in prior work.
 Datasets.
-We mainly evaluate KG-R1 in-domain on (1) WebQSP (Yih et al., 2016)—natural-language questions over Freebase, mostly 1–2 hop, using the official 3,098/1,639 train/test QA split; and (2) ComplexWebQuestions (CWQ) (Talmor & Berant, 2018)—compositional, multi-hop questions over Freebase with a 24,649/3,531 train/test QA split.
-For scalability, we extract 2-hop subgraphs for each question as in RoG (Luo et al., 2024).
-See Appendix B for detailed sources.
+We mainly evaluate KG-R1 in-domain on (1) WebQSP [31]: natural-language questions over Freebase, mostly 1–2 hop questions, using the official 3,098/1,639 train/test QA split; and (2) ComplexWebQuestions (CWQ) [24]: compositional, multi-hop questions over Freebase with a 24,649/3,531 train/test QA split.
+For scalability, we extract 2-hop subgraphs for each question as in RoG [16].
+See Appendix C for detailed sources.
+We evaluate the plug-and-play approach on diverse KGQA benchmarks spanning two categories of KGs: (1) Freebase type—SimpleQA [2] and GrailQA [11], which share the Freebase KG but differ in question complexity and reasoning path distributions; and (2) Wikidata type—T-REx [9] and QALD-10en [25] for cross-schema generalization.
+Appendix C provides a summary of dataset sources, schema, and evaluation splits.
 Metrics.
-Following prior work (Luo et al., 2024; Sun et al., 2024) to evaluate KGQA performance, we use F1 score to consider multiple valid predictions over answer sets and Hit@1 (1 if the gold answer appears in the single search, 0 otherwise).
-For efficiency, we report generated tokens (completion-only), measured end-to-end per query, aggregated across all turns ($1\ldots H$).
+Following prior works [16, 23, 4, park2026prograg] to evaluate KGQA performance, we use Hits@1 (1 if the ground-truth answer appears in the single search, and 0 otherwise).
+For efficiency, we report total tokens and generated tokens (completion-only), measured end-to-end per query, aggregated across all turns.
 This serves as a proxy for inference time and compute cost.
 For fair comparison, all token counts are computed with the Qwen2.5 tokenizer.
-We also report the number of modules to compare workflow complexity, and we analyze single-query latency and batched throughput on a single GPU node.
+We also report the number of modules to compare workflow complexity, and we analyze single-query latency and batched throughput on a single NVIDIA H100 GPU.
 Baselines.
-We compare three classes of approaches: (1) LLM-only methods that do not access an external KG—Vanilla and Chain-of-Thought (CoT) prompting; (2) prior KG-augmented pipelines: RoG (Luo et al., 2024), which uses a fine-tuned modular workflow with LLaMA2-7B-it as the backbone, and ToG (Sun et al., 2024) and ReKnoS (Wang et al., 2025c), which use a prompt-based modular workflow with GPT-3.5; and (3) our method KG-R1.
-For LLM-only baselines, we employ LLM-as-Judge (Zheng et al., 2023) using gpt-4o-mini to semantically match predicted answers to ground truth.
-Full details for baselines are provided in Appendix C.
-N-run beam search with KG-R1.
-Prior works often incorporated k-beam search (Lan & Jiang, 2020; Sun et al., 2024; Luo et al., 2024), retrieving k reasoning paths for enhanced search range on KGs and reporting improved performance.
-Following this idea, we propose KG-R1 N-run beam search : we run $N$ independent generation per question and collect predicted answer sets $\{\hat{A}_{q}^{(i)}\}_{i=1}^{N}$.
-We then take the union $\hat{A}_{q,\text{N-runs}}\;=\;\bigcup_{i=1}^{N}\hat{A}_{q}^{(i)},$ and compute F1 and Hit@1 on this union.
-This simple “$N$-beam” setup broadens search without extra orchestration, and cost scales roughly linearly with $N$ since runs can be executed in parallel.
+We compare three classes of approaches: (1) LLM-only methods that do not access an external KG: IO [brown2020gpt3] and Chain-of-Thought (CoT) [wei2022chain] prompting; (2) prior KG-augmented pipelines: RoG [16], which uses a fine-tuned modular workflow with LLaMA2-7B-it as the planner and we use the Qwen family as reasoner, and ToG [23] and ReKnoS [28], which use a prompt-based modular workflow; and (3) our method KG-R1.
+For LLM-only baselines, we employ LLM-as-Judge [34] using gpt-4o-mini [openai2024gpt4omini] to semantically match predicted answers to ground truth.
+We report the average and standard deviation over 5 runs for all models and baselines.
+Full details for the baselines are provided in Appendix D.
 Implementation details.
-We use VERL (Sheng et al., 2024) and vLLM-backed decoding (Kwon et al., 2023) for implementing the KG-R1 agent.
+We use VERL [22] and vLLM-backed decoding [13] for implementing the KG-R1 agent.
 Training uses distributed optimization with gradient checkpointing and offloading.
-Unless stated, we set $H=5$ turns and N=16N{=}16 rollouts per query.
-Additional implementation details (e.g., hyper-parameters) are provided in Appendix D.4.
-4.1.1 KG-R1 Training Stability and Effectiveness
-Figure 4 shows steady F1 improvement that plateaus after convergence during KG-R1 training on WebQSP and CWQ.
-The validation F1 scores track the training curve, indicating the agent learns generalization.
-Results are reproducible across three random seeds (Figure 6, Appendix), showing strong convergence with low variance.
-Overall, KG-R1 training yields stable, reproducible gains in end-to-end KG-RAG performance.
-Figure 4: F1 score over KG-R1 training on WebQSP and CWQ for Qwen2.5-3B-it. Training (blue) and validation (red).
-4.1.2 Inference Accuracy and Efficiency Results
-Table 1: Performance and efficiency comparison of KGQA methods on WebQSP and CWQ datasets. The max turn number set to $H=5$. All scores are percentages.
-Method Model Modules WebQSP CWQ Performance Efficiency Performance Efficiency F1 Hit@1 Total Gen F1 Hit@1 Total Gen LLM-only Methods Vanilla Qwen2.5-3B-it 1 29.4 46.6 95 30 16.6 21.1 104 42 Vanilla Qwen2.5-7B-it 1 33.2 50.9 95 67 20.7 25.7 104 92 Vanilla LLaMA2-7B-it 1 37.4 54.5 114 255 20.7 24.8 123 255 COT Qwen2.5-3B-it 1 30.6 47.6 131 192 17.3 21.4 140 216 COT Qwen2.5-7B-it 1 35.3 53.5 131 194 22.5 27.1 140 212 COT LLaMA2-7B-it 1 33.8 51.6 165 255 19.0 23.1 174 255 LLM+KG Methods RoG LLaMA2-7B-it 2 70.8 85.7 1.2K 295 56.2 62.6 1.1K 266 ToG GPT-3.5 4 72.8 76.2 3.5K 529 52.9 57.1 3.7K 520 ToG 2.0 GPT-3.5 5 74.5 77.8 39.K 605 65.8 68.9 3.8K 650 ReKnoS GPT-4o-mini 3 73.7 81.1 3.1K 617 64.7 66.8 4.1K 752 KG-R1 (Our Methods) KG-R1 (1 run) Qwen2.5-3B-it 1 77.5 84.7 3.2K 302 70.9 73.8 3.3K 377 KG-R1 (3 runs) Qwen2.5-3B-it 1 85.8 91.7 9.7K 906 81.0 83.9 10.0K 1.1K
-A1.
+Unless stated otherwise, we set $H=5$ as the maximum number of turns and $N=16$ rollouts per query.
+Additional implementation details (e.g., hyperparameters) are provided in Appendix E.4.
+Figure 4: Hits@1 over KG-R1 training on WebQSP and CWQ for Qwen2.5-3B-it.
+4.2 Training Stability and Effectiveness
+Figure 4 shows steady improvements in Hits@1 that gradually plateau as training converges for KG-R1 on both WebQSP and CWQ.
+The validation Hits@1 closely follows the training curve, indicating that the learned policy generalizes well to unseen data in the validation set without significant overfitting.
+Overall, KG-R1 training achieves stable and consistent gains in end-to-end KG-RAG performance.
+4.3 RQ1: Effectiveness and Efficiency Results
+Table 1: Performance and efficiency comparison of KGQA methods on WebQSP and CWQ datasets. The maximum turn number is set to $H=5$. All scores are percentages.
+Datasets $\rightarrow$ WebQSP CWQ
+Methods Modules Hits@1 (%) Total Tokens Gen Tokens Hits@1 (%) Total Tokens Gen Tokens
+Qwen 2.5-7B IO 1 21.37 $\pm$ 0.07 151.59 $\pm$ 0.03 9.29 $\pm$ 0.03 16.63 $\pm$ 0.01 159.76 $\pm$ 0.01 8.57 $\pm$ 0.01
+CoT 1 17.27 $\pm$ 0.09 260.43 $\pm$ 0.02 49.13 $\pm$ 0.02 14.95 $\pm$ 0.02 292.46 $\pm$ 0.01 72.27 $\pm$ 0.01
+SC 1 19.26 $\pm$ 0.19 2098.01 $\pm$ 0.92 407.60 $\pm$ 0.92 17.63 $\pm$ 0.17 2364.62 $\pm$ 1.24 603.13 $\pm$ 1.24
+RoG 2 79.89 $\pm$ 0.35 814.61 $\pm$ 0.70 246.95 $\pm$ 0.67 46.32 $\pm$ 0.49 1219.72 $\pm$ 0.61 259.46 $\pm$ 0.58
+ToG 4 65.00 $\pm$ 4.63 6455.04 $\pm$ 753.60 941.52 $\pm$ 109.27 32.25 $\pm$ 0.43 8396.80 $\pm$ 87.06 1358.54 $\pm$ 17.24
+PoG 5 64.81 $\pm$ 0.84 4474.30 $\pm$ 91.92 241.14 $\pm$ 5.17 32.00 $\pm$ 2.53 8293.67 $\pm$ 1454.76 369.84 $\pm$ 45.59
+ReknoS 3 61.22 $\pm$ 5.46 2822.34 $\pm$ 209.87 558.43 $\pm$ 45.52 53.45 $\pm$ 3.03 2908.55 $\pm$ 222.47 602.08 $\pm$ 39.33
+Qwen 2.5-72B IO 1 24.66 $\pm$ 0.07 153.50 $\pm$ 0.02 11.20 $\pm$ 0.02 22.20 $\pm$ 0.02 162.04 $\pm$ 0.03 10.85 $\pm$ 0.03
+CoT 1 21.84 $\pm$ 0.04 271.46 $\pm$ 0.02 60.16 $\pm$ 0.02 22.92 $\pm$ 0.04 296.81 $\pm$ 0.04 76.62 $\pm$ 0.04
+SC 1 22.62 $\pm$ 0.07 2177.75 $\pm$ 1.23 487.34 $\pm$ 1.23 24.76 $\pm$ 0.27 2385.73 $\pm$ 0.51 624.24 $\pm$ 0.51
+RoG 2 81.63 $\pm$ 3.33 812.32 $\pm$ 7.81 244.53 $\pm$ 7.79 45.05 $\pm$ 2.87 1205.48 $\pm$ 2.51 245.34 $\pm$ 2.51
+ToG 4 74.47 $\pm$ 0.25 6810.27 $\pm$ 39.73 942.91 $\pm$ 5.80 43.18 $\pm$ 1.02 9024.77 $\pm$ 271.68 1358.32 $\pm$ 41.02
+PoG 5 72.88 $\pm$ 6.57 4759.10 $\pm$ 210.29 316.25 $\pm$ 16.85 43.62 $\pm$ 1.38 7969.21 $\pm$ 1631.58 503.29 $\pm$ 116.35
+ReknoS 3 75.23 $\pm$ 9.25 2472.90 $\pm$ 115.94 488.44 $\pm$ 39.87 60.88 $\pm$ 2.54 2278.71 $\pm$ 196.08 548.52 $\pm$ 27.79
+Qwen 3-235B IO 1 24.01 $\pm$ 0.18 136.05 $\pm$ 0.06 14.75 $\pm$ 0.06 20.22 $\pm$ 0.09 147.36 $\pm$ 0.16 17.17 $\pm$ 0.16
+CoT 1 26.94 $\pm$ 0.22 274.66 $\pm$ 0.32 84.36 $\pm$ 0.32 24.38 $\pm$ 0.33 317.42 $\pm$ 0.31 118.23 $\pm$ 0.31
+SC 1 27.60 $\pm$ 0.15 2200.20 $\pm$ 1.26 677.79 $\pm$ 1.26 26.13 $\pm$ 0.23 2551.71 $\pm$ 0.78 958.22 $\pm$ 0.78
+RoG 2 81.68 $\pm$ 0.24 960.58 $\pm$ 5.90 254.31 $\pm$ 0.11 48.94 $\pm$ 0.39 1217.51 $\pm$ 10.26 252.45 $\pm$ 5.18
+ToG 4 75.67 $\pm$ 0.37 6831.66 $\pm$ 28.39 996.12 $\pm$ 5.21 39.82 $\pm$ 0.73 5389.00 $\pm$ 33.01 802.97 $\pm$ 9.84
+PoG 5 79.13 $\pm$ 1.89 6818.26 $\pm$ 908.76 418.26 $\pm$ 49.36 44.69 $\pm$ 0.27 10177.37 $\pm$ 1859.27 548.77 $\pm$ 82.56
+ReknoS 3 84.73 $\pm$ 1.78 2795.53 $\pm$ 139.20 592.38 $\pm$ 4.18 65.63 $\pm$ 4.86 2677.92 $\pm$ 281.97 684.38 $\pm$ 14.06
+KG-R1 (Ours) Qwen2.5-3B 1 82.75 $\pm$ 0.04 3237.17 $\pm$ 30.34 308.67 $\pm$ 0.56 65.31 $\pm$ 0.14 3205.95 $\pm$ 42.34 311.14 $\pm$ 0.52
 KG-R1 demonstrates strong accuracy on trained KGs.
-Table 1 reports accuracy and efficiency.
-On both WebQSP and CWQ, a single run (N=1N{=}1) of KG-R1 achieves competitive F1 and Hit@1 compared with non-KG and prior KG methods, including those using larger foundation models.
-With three runs with beam search, KG-R1 surpasses prior LLM+KG systems that also use $k$-beam search by a clear margin (WebQSP: F1 $77.5{\to}85.8$, Hit@1 $84.7{\to}91.7$; CWQ: F1 $70.9{\to}81.0$, Hit@1 $73.8{\to}83.9$).
-Overall, these results show that a lightweight KG-R1 agent delivers strong KG-RAG performance relative to previous baselines.
-A2.
-KG-R1 achieves efficient inference.
-Table 1 reports efficiency (generation tokens and number of modules).
-For a single run (N=1N{=}1), KG-R1 produces $\sim$60 generation tokens per turn (300–370 tokens at H=5H{=}5), substantially below prior prompt-based LLM+KG systems (e.g., ToG $\sim$520 and ReKnoS $\sim$600–650) and comparable to the fine-tuned method RoG ($\sim$250–300).
-This small generation-token budget, coupled with a lightweight model, implies improved throughput and latency.
-Under N=1N{=}1, the total tokens are comparable to prior methods, indicating a similar memory (KV-cache) footprint.
-With three-run beam search (N=3N{=}3), token usage scales nearly linearly, exposing a tunable accuracy–efficiency trade-off via $N$.
-A3.
+Table 1 reports Hits@1 and token efficiency on WebQSP and CWQ.
+Despite being trained from a lightweight Qwen2.5-3B base model, KG-R1 achieves the strongest overall performance among all compared methods on WebQSP, reaching a Hits@1 of 82.75, outperforming RoG built on Qwen2.5-7B (79.89), Qwen2.5-72B (81.63), and even Qwen3-235B (81.68).
+It also exceeds or is comparable to stronger multi-module baselines such as ToG (75.67), PoG (79.13), and ReKnoS (84.73) while using a substantially smaller backbone.
+On the more compositional CWQ, KG-R1 achieves a Hits@1 of 65.31, improving over RoG by a large margin across all scales (46.32–48.94), and surpassing ReKnoS built on Qwen2.5-7B (53.45), Qwen2.5-72B (60.88), and Qwen3-235B (55.63).
+Notably, KG-R1 remains highly competitive with ReKnoS built on Qwen3-235B (65.31 vs.
+65.63 on CWQ), despite using a Qwen2.5-3B backbone that is nearly two orders of magnitude smaller.
+This highlights that reinforcement learning over structured KG actions can substantially improve reasoning effectiveness beyond what is achievable by model scaling alone.
+[…]
+Single-agent design reduces system complexity.
+Beyond token efficiency, KG-R1 simplifies the KG-RAG pipeline into a single-agent architecture.
+Prior methods typically require 2–5 modules (e.g., retrieval, planning, verification, and answering), whereas KG-R1 performs retrieval and reasoning within a unified policy using only one module.
+This reduces deployment complexity and avoids error propagation across stages.
+Despite this simplification, KG-R1 remains competitive or superior to multi-module systems across both datasets, showing that structured RL can effectively consolidate KG interaction and reasoning into a lightweight unified agent.
 Latency and throughput analysis.
 On a single NVIDIA H100 GPU, we measure (i) single-query end-to-end latency and (ii) batched throughput.
-The single-query latency averages $6.4\pm 1.5$ s per question.
-Batched throughput reaches $3.7$ samples per second at batch size 64.
+The single-query latency averages $6.4\pm 1.5$ s per question, and the batched throughput reaches $3.7$ samples per second with a batch size of 64.
 The results suggest the feasibility of KG-R1’s single-node deployment.
-See Appendix E.2 for the full results.
-4.2 Cross-KG Transfer via Plug-and-Play
-In this subsection, we evaluate the plug-and-play approach (Sec. 3.4) for zero-shot generalization of our KG-R1 agent on knowledge graphs outside its training distribution.
-Datasets We evaluate the plug-and-play approach on diverse KGQA benchmarks spanning three categories of KGs.
-(1) Freebase type: SimpleQA (Bordes et al., 2015) and GrailQA (Gu et al., 2021), which share the Freebase KG but differ in question complexity and reasoning path distributions.
-(2) Wikidata type: T-REx (Elsahar et al., 2018) and QALD-10en (Usbeck et al., 2023) for cross-schema generalization.
-(3) Temporal reasoning benchmark: MultiTQ (Chen et al., 2023), which requires reasoning over the ICEWS temporal KG (Boschee et al., 2015) at day, month, and year granularities.
-Appendix B provides a summary of dataset sources, schema, and evaluation splits.
-Baselines For LLM-only methods, we evaluate (i) Vanilla and (ii) Chain-of-Thought (CoT) prompting with Qwen2.5–3B-it.
-For LLM+KG methods, we report results for RoG, ToG, and ReKnoS without modifying their modules (i.e prompts and models).
-In addition, for each KGQA benchmark, we include one or two recent state-of-the-art KG-specific methods (trained/tuned on the target dataset with task-specific pipelines) and use their published test-set scores.
-Full baseline details are provided in Appendix C.3.
-Table 2: Zero-shot cross-KG transferability of KG-R1.
+See Appendix F.1 for the full results.
+4.4 RQ2: Cross-KG Transferability
+Table 2: Zero-shot cross-KG transferability of KG-R1 (Hits@1, %).
 Agents are trained on WebQSP or CWQ and evaluated on new benchmarks by swapping only the KG-backend server (no policy retraining).
-AVG.
-is averaged across QA samples.
-The max turn number set to $H=5$.
-All scores are percentages.
-Freebase-based Wikidata-based Temporal
-Method Trained KG SimpleQA GrailQA T-REx QALD-10en MultiTQ AVG.
-F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1 F1 / Hit@1
-LLM-only Methods
-Vanilla Qwen 2.5 3B-IT 13.7 / 13.7 15.9 / 15.9 24.4 / 24.4 23.8 / 23.8 2.2 / 5.4 19.4 / 19.8
-CoT Qwen 2.5 3B-IT 10.9 / 10.9 18.2 / 18.2 22.0 / 22.0 25.9 / 25.9 1.9 / 3.9 18.0 / 18.2
-LLM+KG Methods
-KG-specific Baselines
-$79.6^{1}/80.6^{1}$
-$76.1^{2}/77.1^{2}$
-$84.4^{3}/79.1^{3}$
-$64.9^{4}/59.6^{4}$
-— / $85.1^{5}$
-$49.8^{6}/50.8^{6}$
-$62.2^{7}/63.2^{7}$
-— / $79.7^{8}$
-— / $38.0^{9}$
-RoG 13.5 / 43.5 6.4 / 23.7 8.1 / 36.4 11.4 / 44.3 3.6 / 8.7 8.1 / 34
-ToG 50.2 / 53.6 63.9 / 68.7 79.3 / 76.4 48.7 / 50.2 25.1 / 27.9 66.2 / 65.9
-ReKnoS 50.7 / 52.9 65.8 / 71.9 81.4 / 78.9 50.6 / 53.8 28.5 / 31.1 68.2 / 68.2
-KG-R1 (Our Methods)
-KG-R1 (1 run) WebQSP 59.1 / 59.1 32.8 / 38.5 80.5 / 84.5 51.9 / 53.4 21.6 / 31.4 64.0 / 68.3
-KG-R1 (1 run) CWQ 64.6 / 64.7 42.8 / 50.2 81.3 / 85.6 55.9 / 57.7 27.1 / 38.9 67.2 / 72.1
-KG-R1 (3 runs) WebQSP 69.1 / 69.1 44.6 / 52.1 84.6 / 88.8 61.3 / 62.8 33.5 / 45.7 70.9 / 75.8
-KG-R1 (3 runs) CWQ 73.1 / 73.1 52.8 / 61.0 86.8 / 91.5 63.9 / 65.5 36.2 / 48.4 74.1 / 79.4
-1SPARKLE 2GETT-QA 3SG-KBQA 4ARG-KBQA 5ATLAS 6COT-SPARQL 7DFSL-MQ 8Prog-TQA 9ARI ; see Table 7 for references.
-A4.
-KG-R1 transfers across diverse types of KGs.
-Table 2 reports zero-shot, plug-and-play results.
-For comparability, we report AVG., the F1/Hits@1 averaged across the five KGQA datasets, weighted by QA counts.
-KG-R1 (1 run) substantially outperforms LLM-only baselines using the same base model—Vanilla (19.4/19.8 F1/Hits@1) and CoT (18.0/18.2)—reaching 64.0/68.3 when trained on WebQSP and 67.2/72.1 when trained on CWQ.
-The gains hold not only on Freebase-based QA but also on Wikidata and Temporal benchmarks, indicating that reinforcement-learned KG-RAG transfer across diverse KGs rather than narrow, in-domain improvements.
-A5.
-KG-R1 achieves accurate inference comparable to LLM+KG baselines.
-Against LLM+KG baselines, KG-R1 (N=1) is slightly better on average: ToG (66.2/65.9) and ReKnoS (68.2/68.2).
-Notably, KG-R1, with a 3B model, delivers higher Hits@1 than both ToG and ReKnoS that uses strong foundation models.
-Increasing to N=3 runs boosts performance to 74.1/70.4 (CWQ-trained), making KG-R1 comparable to KG-specific baselines on average.
-We observe that RoG collapses (0/0 average) revealing brittle, schema-specific planning with fine-tuned model that does not transfer.
-Overall, KG-R1 attains strong transferability among plug-and-play methods with a lightweight model, supporting its practicality for real-world KG-RAG deployment across diverse KGs.
-A6.
-Training data on transferability.
-KG-R1 models trained on CWQ consistently outperform WebQSP-trained variants across all transfer targets, averaging 1–5% higher F1 and Hits@1.
-Given that CWQ provides a larger training set with greater multi-hop complexity than WebQSP, this suggests that exposure to more complex and diverse reasoning patterns yields superior generalization.
+AVG is averaged across QA samples.
+The maximum turn number is set to $H=5$.
+Freebase-based Wikidata-based
+[…]
+Qwen 2.5-7B IO 10.30 $\pm$ 0.10 8.07 $\pm$ 0.06 22.79 $\pm$ 0.05 17.43 $\pm$ 0.01 18.84
+CoT 7.07 $\pm$ 0.06 7.73 $\pm$ 0.12 17.38 $\pm$ 0.04 21.05 $\pm$ 0.06 14.82
+SC 8.70 $\pm$ 0.20 8.60 $\pm$ 0.44 20.01 $\pm$ 0.12 20.32 $\pm$ 0.46 16.93
+RoG 12.97 $\pm$ 0.35 17.67 $\pm$ 0.15 22.37 $\pm$ 0.27 29.33 $\pm$ 0.46 20.76
+ToG 52.37 $\pm$ 0.80 54.73 $\pm$ 4.62 58.55 $\pm$ 1.11 50.95 $\pm$ 0.92 56.84
+PoG 66.38 $\pm$ 0.39 69.88 $\pm$ 0.84 75.69 $\pm$ 0.08 46.07 $\pm$ 0.25 72.28
+ReknoS 56.43 $\pm$ 1.22 69.58 $\pm$ 0.35 78.17 $\pm$ 0.03 37.94 $\pm$ 1.21 72.21
+Qwen 2.5-72B IO 14.33 $\pm$ 0.15 11.63 $\pm$ 0.12 26.16 $\pm$ 0.02 17.45 $\pm$ 0.06 22.17
+CoT 11.07 $\pm$ 0.12 10.53 $\pm$ 0.06 24.48 $\pm$ 0.02 20.52 $\pm$ 0.17 20.57
+SC 11.60 $\pm$ 0.44 11.43 $\pm$ 0.25 26.01 $\pm$ 0.09 21.62 $\pm$ 0.60 21.86
+RoG 20.40 $\pm$ 0.26 22.63 $\pm$ 4.51 32.50 $\pm$ 0.09 44.94 $\pm$ 0.69 30.07
+ToG 57.83 $\pm$ 0.29 71.53 $\pm$ 0.76 67.93 $\pm$ 0.18 59.66 $\pm$ 0.46 66.67
+PoG 65.67 $\pm$ 0.42 75.30 $\pm$ 0.99 79.52 $\pm$ 0.08 48.18 $\pm$ 3.53 75.63
+ReknoS 68.40 $\pm$ 5.15 73.67 $\pm$ 0.60 77.11 $\pm$ 0.02 39.94 $\pm$ 0.30 73.77
+Qwen 3-235B IO 11.37 $\pm$ 0.12 12.80 $\pm$ 0.10 34.61 $\pm$ 0.01 24.12 $\pm$ 0.17 27.99
+CoT 14.40 $\pm$ 0.36 14.60 $\pm$ 0.30 37.72 $\pm$ 0.18 28.43 $\pm$ 0.17 30.97
+SC 14.73 $\pm$ 0.42 15.73 $\pm$ 0.15 38.79 $\pm$ 0.27 29.53 $\pm$ 0.35 31.94
+RoG 18.93 $\pm$ 0.51 18.10 $\pm$ 0.20 37.31 $\pm$ 0.19 37.54 $\pm$ 1.20 32.19
+ToG 41.27 $\pm$ 3.61 53.20 $\pm$ 0.53 77.00 $\pm$ 0.10 57.76 $\pm$ 1.05 68.01
+PoG 67.90 $\pm$ 0.17 77.93 $\pm$ 0.45 78.75 $\pm$ 0.13 52.55 $\pm$ 1.04 75.97
+ReknoS 68.01 $\pm$ 3.37 75.67 $\pm$ 0.74 77.05 $\pm$ 0.03 39.44 $\pm$ 0.35 73.92
+Qwen2.5 KG-R1 (WebQSP trained) 69.18 $\pm$ 0.10 68.55 $\pm$ 0.07 83.38 $\pm$ 0.08 41.96 $\pm$ 0.05 77.50
+3B KG-R1 (CWQ trained) 57.31 $\pm$ 0.08 64.84 $\pm$ 0.04 80.10 $\pm$ 0.05 34.74 $\pm$ 0.12 72.82
+Zero-shot transfer across unseen KGs.
+Table 2 reports zero-shot plug-and-play results, where the learned policy is directly applied to unseen KGs by swapping only the backend graph.
+For comparability, we report AVG, the Hits@1 averaged across all evaluation QA samples.
+KG-R1 demonstrates strong cross-KG transferability across both Freebase- and Wikidata-based benchmarks.
+When trained on WebQSP, KG-R1 achieves an average Hits@1 of 77.50, substantially outperforming LLM-only baselines (IO: 18.84, CoT: 14.82, SC: 16.93) and surpassing prior KG-augmented methods, including ToG (56.84), PoG (72.28), and ReKnoS (72.21).
+The gains are consistent across all transfer benchmarks, with particularly strong improvements on T-REx (83.38 vs.
+78.17 for ReKnoS).
+When trained on CWQ, KG-R1 achieves 72.82 Hits@1, remaining clearly above LLM-only baselines and competitive with strong KG-RAG methods built on much larger foundation models.
+In particular, the WebQSP-trained KG-R1 also outperforms ToG, PoG, and ReKnoS when these methods use Qwen3-235B backbones (77.50 vs.
+68.01, 75.97, and 73.92, respectively), demonstrating that learned KG interaction policies can generalize better than simply scaling model size.
+Notably, transfer gains are especially pronounced on Wikidata-based benchmarks (T-REx and QALD-10en), where schema differences are larger.
+This suggests that the schema-agnostic action design of KG-R1 enables robust generalization across diverse KG structures.
 4.5 Ablation Studies
-We conduct ablation studies for the three core components of KG-R1: reward design (Section 3.3), RL algorithm (Section 2), and KG-server implementation (Section 3) and base LLM size.
-We train Qwen-2.5-3B-it on WebQSP and CWQ with maximum turn number $H=5$ and report final step performance(F1/Hit@1/Ret.rate in Table 3).
-The training curves are shown in Figures 8, 9.
-Table 3: Ablation studies of KG-R1 components on CWQ and WebQSP datasets.
-Method WebQSP CWQ F1 Hit@1 Ret.Rate F1 Hit@1 Ret.Rate KG-R1 Default KG-R1 77.2 84.3 76.8 66.8 69.7 51.4 Reward Ablation w/o Turn Rewards 67.1 (-13.1%) 76.0 (-9.8%) 68.1 51.9 (-22.3%) 52.9 (-24.1%) 44.7 w/o Turnwise Advantage 49.6 (-35.8%) 61.2 (-27.4%) 22.8 63.5 (-4.9%) 64.5 (-7.5%) 49.4 w/o Retrieval Reward 45.8 (-40.6%) 57.7 (-31.5%) 18.6 66.0 (-1.2%) 67.0 (-3.9%) 47.2 Other Ablations w PPO 0.0 (-100.0%) 0.0 (-100.0%) 0.0 0.0 (-100.0%) 0.0 (-100.0%) 0.0 w/o Hierarchical Rel.
-Retr.
-76.8 (-0.6%) 83.1 (-1.5%) 76.0 55.5 (-16.9%) 58.6 (-15.9%) 39.7 w Qwen-2.5-7B-it 79.6 (+3.1%) 86.6 (+2.7%) 83.4 (+8.6%) 65.6 (-1.8%) 68.7 (-1.4%) 58.8 (+14.4%)
-PPO training crashed.    Peak values before 7B model training collapse.
-A7.
-Reward function.
-First, removing turn-level rewards yields the largest drop on both WebQSP (13.1% / 9.9% in F1 score and Hit@1, respectively) and CWQ (22.3% / 24.1%), highlighting the importance of per-turn validity/effectiveness signals.
-Next, disabling the turn-wise advantage calculation produces a substantial decline on WebQSP (35.8% / 27.4%) but only a moderate drop on CWQ (4.9% / 7.5%), indicating that turn-specific learning signals matter more for certain datasets.
-Finally, removing the retrieval reward significantly impacts WebQSP (40.7% / 31.5%) but marginally affects CWQ (1.2% / 3.9%), though it substantially reduces the retrieval success rate of gold entities (Ret.Rate) on both datasets, suggesting that it is essential for motivating exploration of the KG.
-A8.
-RL algorithm (GRPO vs.
-PPO).
-Replacing GRPO with vanilla PPO (Ouyang et al., 2022) results in learning collapse (0.0 across all metrics).
-Under PPO, we observe reward hacking (see examples in Appendix E.3.3): the agent fabricates ”retrieved” content matching expected formats to earn reward.
-Since PPO’s value critic (another LLM) cannot distinguish genuine KG retrievals from hallucinations, its advantage estimates become unstable, driving training collapse.
-This underscores the importance of relative advantage methods like GRPO for multi-step reasoning.
-A9.
-Retrieval format.
-In our standard KG-R1 setup, one-hop relations retrieved by $\mathtt{get\_head\_relations}$ and $\mathtt{get\_tail\_relations}$ are appended to the agent’s context as a flat, comma-separated list (e.g., “rel1, rel2, …”).
-This can lead to excessive token consumption when many relations are retrieved.
-Because Freebase relations use a compositional format (e.g., “domain.type.property”), we tested a hierarchical relation retrieval that groups relations into a tree to reduce redundancy (see Appendix E.3.2).
-Our ablation studies show that switching from flat lists to a hierarchical format reduces performance—WebQSP shows 0.5% drop, while CWQ experiences 16.9% drop.
-These results suggest that despite its higher token usage, the flat retrieval format provides more direct access to relation information, which proves critical for training.
-A10.
+We ablate four key components of KG-R1: reward design, RL algorithm, retrieval format, and model scale.
+We train on WebQSP and CWQ with turn number $H=5$ and report final Hits@1 and retrieval success rate (Ret.
+Rate), defined as the fraction of queries where any ground-truth answer entity appears in the retrieved entities, in Table 3.
+The training curves are shown in Figure 7.
+Table 3: Ablation studies of KG-R1 components.
+Datasets $\rightarrow$ WebQSP CWQ
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Model Scale.
-The 7B model improves more quickly than the 3B model but exhibits training collapse (around 350 steps on WebQSP and 200 steps on CWQ).
-This aligns with reports that fast distribution shift during GRPO—amplified by larger models and higher update rates—drives entropy collapse and model divergence (Jin et al., 2025; Liu et al., 2025).
+Scaling the base model from Qwen2.5-3B to 7B does not improve performance (WebQSP: 82.75 vs.
+82.44; CWQ: 65.31 vs.
+65.85), indicating that KG-R1’s gains mainly come from reinforcement-learned retrieval policies rather than model size alone.
 5 Conclusion
-We propose KG-R1, a single-agent, multi-turn KG-RAG framework in which one LLM queries a lightweight KG server and is optimized via RL.
-Across KG-augmented QA benchmarks, KG-R1 attains strong accuracy while using markedly fewer tokens and lower inference cost than prior work.
-It also supports plug-and-play transfer, retaining robust performance across KGs.
+We propose KG-R1, a single-agent, multi-turn KG-RAG framework where an LLM queries a lightweight KG server and is optimized via RL.
+Across KG-augmented QA benchmarks, KG-R1 achieves strong accuracy while using markedly fewer tokens and lower inference cost than in prior work.
+It also supports plug-and-play transfer, maintaining robust performance across KGs.
 Together, these results position KG-R1 as a practical and deployable KG-RAG system for real-world use.
-Appendix
-Appendix A Theoretical Proofs
-A.1 Completeness and Transferability of the KG-Server
-Here we provide robust proofs for completeness and transferability of $\mathcal{U}_{ret}$ in our KG retrieval server.
-Preliminaries. Let $G=\{(e,r,e^{\prime})\mid e,e^{\prime}\!\in\!\mathcal{E},\,r\!\in\!\mathcal{R}\}$ be a directed KG. Define the KG-R1 action set
+Appendix A Discussion
+Societal Impact.
+KG-R1 aims to improve knowledge-grounded reasoning by enabling language models to interact with structured knowledge graphs through learned retrieval policies.
+This may benefit applications requiring factual consistency and traceable reasoning, such as question answering, scientific knowledge access, and enterprise knowledge systems.
+By reducing reliance on complex multi-module pipelines, KG-R1 may also improve the accessibility and deployment efficiency of knowledge-grounded reasoning systems.
+Potential Negative Impact.
+Like other retrieval-augmented systems, KG-R1 may propagate incorrect or outdated knowledge if the underlying knowledge graph is incomplete, noisy, or maliciously manipulated.
+In high-stakes domains such as medicine or finance, such errors could lead to misleading outputs.
+Additionally, improved knowledge-grounded reasoning may be misused to automate large-scale misinformation generation with higher factual coherence.
+Limitations.
+KG-R1 assumes access to a reasonably structured and queryable knowledge graph, which may not be available in all domains.
+Its performance depends on the quality and schema consistency of the KG, and transferability may degrade under severe schema mismatch or noisy graph structures.
+Furthermore, while KG-R1 improves inference efficiency through structured interaction, reinforcement learning training remains computationally expensive compared to standard supervised fine-tuning.
+Appendix B Completeness and Transferability of the KG-Server
+Here, we discuss the completeness and transferability of $\mathcal{U}_{ret}$ in our KG retrieval server.
+Preliminaries.
+Let $G=\{(e,r,e^{\prime})\mid e,e^{\prime}\!\in\!\mathcal{E},\,r\!\in\!\mathcal{R}\}$ be a directed KG. Define the KG-R1 action set
 $\mathcal{U}_{\text{ret}}=\left\{\begin{aligned} &\mathtt{get\_tail\_relations},\ \mathtt{get\_head\_relations},\\ &\mathtt{get\_tail\_entities},\ \mathtt{get\_head\_entities}\end{aligned}\right\}$
@@
 $Z:\ e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}e_{\ell},\quad(e_{i-1},r_{i},e_{i})\in G.$
-Proposition A.1 (Finite-Horizon Bound).
-For any path $Z$ of length $\ell$, there exists an action sequence from $\mathcal{U}_{\text{ret}}$ of length at most $\ell$ whose final output contains $e_{\ell}$.
-Proof.
-Along the path $Z$, for each step $i=1,\ldots,\ell$ the call $\mathtt{get\_tail\_entities}(e_{i-1},r_{i})$ returns a set that contains $e_{i}$.
-Therefore, after the $\ell$-th step the output contains $e_{\ell}$.
-Proposition A.2 (Completeness (Integrity)).
-Every finite reasoning path in $G$ can be realized by a finite sequence of actions from $\mathcal{U}_{\text{ret}}$ whose outputs include its terminal entity.
+[…]
 Proof.
@@
 For the inductive step, assume the claim holds for length $\ell-1$.
-By the hypothesis we reach $e_{\ell-1}$; applying $\mathtt{get\_tail\_entities}(e_{\ell-1},r_{\ell})$ then returns a set containing $e_{\ell}$.
-Proposition A.3 (Schema-Free Transferability).
-The semantics of each $a\in\mathcal{U}_{\text{ret}}$ depend only on membership of triples $(e,r,e^{\prime})$ in $G$.
-Replacing $G$ with any directed KG $G^{\prime}$ preserves action meaning and allows a fixed client policy to transfer unchanged.
-Proof.
-Each operator is defined as a set comprehension over triples in $G$, independent of schema, datatypes, or ontology rules.
-Under $G\mapsto G^{\prime}$, the same comprehensions define the actions on $G^{\prime}$, so a learned policy remains well-defined.
-Proposition A.4 (Minimality of $\mathcal{U}_{\text{ret}}$).
-No proper subset of $\mathcal{U}_{\text{ret}}$ supports both symmetric navigation and local relation discovery; any operator that only filters or composes these four is derivable.
-Proof.
-Removing $\mathtt{get\_tail\_entities}$ (resp.
-$\mathtt{get\_head\_entities}$) prevents forward (resp.
-backward) traversal.
-Removing both $\mathtt{get\_tail\_relations}$ and $\mathtt{get\_head\_relations}$ blocks local relation discovery when the relation set is unknown.
-Conversely, a conditional/filtered variant can be constructed by enumerating $\mathtt{get\_tail\_relations}(e)$ and intersecting results of $\mathtt{get\_tail\_entities}$; hence no additional primitives are required.
-Appendix B Datasets
-B.1 KGQA Datasets
-These are the KGQA datasets we used in our work.
-Freebase
-WebQSP (Yih et al., 2016) (full test; 1,639) Open-domain questions with mostly 1–2 hop reasoning over Freebase MIDs and SPARQL annotations.
-ComplexWebQuestions (CWQ) (Talmor & Berant, 2018) (full test; 3,531) Compositional multi-hop questions generated from SPARQL templates that stress longer chains and constraint composition; used to probe multi-turn retrieval quality and robustness without dataset-specific rules.
-SimpleQuestions (SimpleQA) (Bordes et al., 2015) (1,000 test) Single-relation 1-hop questions over Freebase; serves as a retrieval-fidelity and token-efficiency baseline for KG-R1.
+By the induction hypothesis, we reach $e_{\ell-1}$; applying $\mathtt{get\_tail\_entities}(e_{\ell-1},r_{\ell})$ then returns a set containing $e_{\ell}$.
+Schema-Free Transferability.
+The retrieval action interface $\mathcal{U}_{\text{ret}}$ is defined purely in terms of set operations over triples $(e,r,e^{\prime})$, independent of any schema, ontology, or datatype constraints.
+As a result, the same action interface can be applied to different knowledge graphs without modification.
+This allows a learned policy to be executed on a new KG by simply replacing the backend graph.
+However, this property guarantees only that the policy remains well-defined under KG replacement.
+The effectiveness of transfer depends on the similarity between the training and target KGs and is evaluated empirically in Section 4.4.
+Appendix C Datasets
+C.1 KGQA Datasets
+Here, we list the KGQA datasets used in this paper.
+Freebase-based Datasets
+  • WebQSP [31] (1,639 test QA pairs) Open-domain questions with mostly 1–2 hop reasoning over Freebase MIDs and SPARQL annotations.
+  • ComplexWebQuestions (CWQ) [24] (3,531 test QA pairs) Compositional multi-hop questions generated from SPARQL templates that stress longer chains and constraint composition are used to probe multi-turn retrieval quality and robustness without dataset-specific rules.
+  • SimpleQuestions (SimpleQA) [2] (1,000 sampled test QA pairs) Single-relation 1-hop questions over Freebase serve as a retrieval-fidelity and token-efficiency baseline for KG-R1.
 We randomly sample 1,000 QA from the original test split (21,687).
-GrailQA (Gu et al., 2021) (1,000 test) Diverse compositional questions emphasizing generalization under Freebase; handled with the same minimal action interface and no hand-crafted schemas.
+  • GrailQA [11] (1,000 sampled test QA pairs) Diverse compositional questions emphasizing generalization under Freebase; handled with the same minimal action interface and no hand-crafted schemas.
 We randomly sample 1,000 from the original test split (13,231).
-Wikidata
-T-REx (Elsahar et al., 2018) (5,000 test) Large-scale slot-filling–style QA grounded in Wikidata triples; used to assess scalability and coverage under a different KG schema.
+Wikidata-based Datasets
+  • T-REx [9] (5,000 test QA pairs) Large-scale slot-filling–style QA grounded in Wikidata triples; used to assess scalability and coverage under a different KG schema.
 We randomly sample 5,000 from the corpus ($\sim$2.2M).
-QALD-10en (Usbeck et al., 2023) (333 test) Manually curated, linguistically varied questions over Wikidata; useful for evaluating precision on a small but challenging set. We evaluate on 333 examples.
-Temporal KG
-MultiTQ (Chen et al., 2023) (1,000 test) Time-aware QA requiring temporal qualifiers (e.g., during, from–to); evaluates KG-R1 on temporal reasoning with time-scoped entities and relations.
-We randomly sample 1,000 from the original test split (9,000).
-B.2 Dataset Preprocessing
+  • QALD-10en [25] (333 test QA pairs) Manually curated, linguistically varied questions over Wikidata; useful for evaluating precision on a small but challenging set. We evaluate on 333 examples.
+C.2 Dataset Pre-processing
 Two-hop Subgraph Extraction Methodology.
-Following subgraph preprocessing practice RoG (Luo et al., 2024), we builds a question-specific near subgraph to shrink the search space and suppress spurious matches: starting from the linked anchor entities $e_{q}$ and gold answers $A_{q}$, it performs a breadth-first expansion over the KG $G$ up to the dataset’s maximum 2-hop radius $h$ The processed subgraph for each question is cached and used for KG retrieval server in KG-R1.
-B.3 Freebase, Wikidata, and MultiTQ schema comparison
-The following tables show the different schemas of Freebase, Wikidata, and MultiTQ.
+Following the subgraph pre-processing practice RoG [16], we build a question-specific near subgraph to shrink the search space and suppress spurious matches: starting from the linked anchor entities $e_{q}$ and gold answers $A_{q}$, it performs a breadth-first expansion over the KG $G$ up to the dataset’s maximum 2-hop radius $h$.
+The processed subgraph for each question is cached and used for the KG retrieval server in KG-R1.
+C.3 Freebase and Wikidata Schema Comparison
+The following tables show the different schemas of Freebase and Wikidata.
 Table 4: Freebase knowledge graph dataset used for KGQA evaluation. All datasets share the same underlying Freebase knowledge graph structure with 4.9M entities and 663 relations.
 Dataset Entities Relations
-Freebase - P!nk - Gender - Ice Cube - United States of America - Nemanja Mikic - broadcast.content.artist - people.person.nationality - music.artist genre - location.location.containedby - basketball.basketball.player position
+Freebase - P!nk - Gender - Ice Cube - United States of America - Nemanja Mikic - broadcast.content.artist - people.person.nationality - music.artist_genre - location.location.containedby - basketball.basketball.player_position
 Questions
-WebQSP (4,737 test questions): Q: what does jamaican people speak A: [Jamaican English, Jamaican Creole English Language] CWQ (3,531 test questions): Q: Lou Seal is the mascot for the team that last won the World Series when?
-A: [2014 World Series] SimpleQA (21,687 test questions): Q: where is the madam satan located A: [United States of America] GrailQA (13,231 test questions): Q: which play is produced by the illusion?
+WebQSP (1,639 test questions): Q: What does jamaican people speak?
+A: [Jamaican English, Jamaican Creole English Language] CWQ (3,531 test questions): Q: Lou Seal is the mascot for the team that last won the World Series when?
+A: [2014 World Series] SimpleQA (21,687 test questions): Q: Where is the madam satan located?
+A: [United States of America] GrailQA (13,231 test questions): Q: Which play is produced by the illusion?
 A: [The Illusion]
@@
 T-Rex (2.2M test questions): Q: What is the occupation of Albert Einstein? A: [Physicist] QALD-10en (250 test questions): Q: Which companies were founded by Bill Gates? A: [Microsoft]
-Table 6: Temporal knowledge graph dataset used for KGQA evaluation. MultiTQ focuses on temporal reasoning with time-aware entities and relations.
-Dataset Entities Relations
-Temporal KG - Barack Obama (2009-2017) - World War II (1939-1945) - Steve Jobs (1955-2011) - Cold War (1947-1991) - Nelson Mandela (1994-1999) - president during - occurred during - CEO from to - started in - ended in
-Questions
-MultiTQ (9,000 test questions): Q: Who was the president of the United States when the iPhone was first launched?
-A: [George W.
-Bush] Q: Which major historical event ended the same year the European Union was established?
-A: [Cold War] Q: What technology company was founded during World War II?
-A: [IBM]
-Appendix C Baselines
-C.1 LLM-only Methods
+Appendix D Baselines
+D.1 LLM-only Methods
 Setup.
-We evaluate Vanilla and CoT in a zero-shot setting without access to the KG retrieval server (i.e., no KG augmentation).
-All baselines run on Qwen-2.5-3B-IT, Qwen-2.5-7B-IT, and LLaMA-2-7B-Chat with temperature=0 and top_p=50..
-Prompts. We use the following prompt templates for vanilla and CoT baselines.
-Prompte template for Vanilla setup Answer the given question directly and concisely based on your knowledge.
-Format your answer as: Answers: [”answer1”, ”answer2”, …].
-For single answers, use: Answers: [”answer”].
+We evaluate IO, CoT, and SC in a zero-shot setting without access to the KG retrieval server (i.e., no KG augmentation).
+All baselines run on Qwen2.5-7B-it, Qwen2.5-72B-it, Qwen3-235B-A22B-Instruct-2507 with temperature=0 and top_p=50.
+Prompts. We use the following prompt templates for IO and CoT baselines (the SC baseline performs majority voting over CoT results and hence uses the same prompt as CoT).
+Prompte template for IO Prompt setup Answer the given question directly and concisely based on your knowledge.
+Format your answer as: Answers: ["answer1", "answer2", …].
+For single answers, use: Answers: ["answer"].
 Question: [Question] Answers:
@@
 IMPORTANT: Follow this exact format: 1.
-Start with ”Reasoning:” followed by your step-by-step thinking.
+Start with "Reasoning:" followed by your step-by-step thinking.
 2.
-End with ”Answers:” followed by your final answer in brackets.
+End with "Answers:" followed by your final answer in brackets.
 3.
-Do NOT put ”Answers:” before your reasoning.
-Format your answer as: Answers: [”answer1”, ”answer2”, …].
-For single answers, use: Answers: [”answer”].
+Do NOT put "Answers:" before your reasoning.
+Format your answer as: Answers: ["answer1", "answer2", …].
+For single answers, use: Answers: ["answer"].
 Question: [Question] Reasoning:
 Evaluation.
-Baseline outputs (Vanilla and CoT) differ in format from the KGQA gold answer set, so we employ an LLM-as-judge to align them.
-For each question, gpt-5-mini (OpenAI API) is given the question, the ground-truth answer set $A_{q}$, and parsed predited entities $\hat{A}_{q}$ from the base model’s output with a concise semantic-entity-matching prompt (Box below); it returns only a binary vector indicating, for each gold entity in order, whether the prediction refers to the same real-world entity at the same specificity (1 for exact/equivalent matches, e.g., “Apple Inc.” = “Apple”; 0 for overly general predictions, e.g., “Islam” vs [“Shia Islam”, “Sunni Islam”]).
+Baseline outputs (IO and CoT) differ in format from the KGQA gold answer set, so we employ an LLM-as-judge to align them.
+For each question, gpt-5-mini (OpenAI API) is given the question, the ground-truth answer set $A_{q}$, and parsed predicted entities $\hat{A}_{q}$ from the base model’s output with a concise semantic-entity-matching prompt (shown in the box below); it returns only a binary vector indicating, for each gold entity in order, whether the prediction refers to the same real-world entity at the same specificity (1 for exact/equivalent matches, e.g., “Apple Inc.” = “Apple”; 0 for overly general predictions, e.g., “Islam” vs [“Shia Islam”, “Sunni Islam”]).
 We report Pass@K (K=1,2,3,4), F1, precision, recall, and generation-token usage.
@@
 Your task is to decide, for each gold entity (in order), whether the model’s prediction refers to the same real-world entity with the same level of specificity.
-Respond only with a binary vector in the format: $[0,1,0,1]$ Rules: - Exact matches or equivalent references → output 1 (e.g., ”Apple Inc.” = ”Apple”).
-- Too general or not specific enough when specificity is required → output 0 (e.g., ”Islam” vs [”Shia Islam”, ”Sunni Islam”]).
+Respond only with a binary vector in the format: $[0,1,0,1]$ Rules: - Exact matches or equivalent references → output 1 (e.g., "Apple Inc." = "Apple").
+- Too general or not specific enough when specificity is required → output 0 (e.g., "Islam" vs ["Shia Islam", "Sunni Islam"]).
 Gold entities: [gold entity list] Predicted entities: [predicted entity list] Your Response:
-C.2 LLM+KG baselinse
-RoG (Luo et al., 2024) Reasoning on Graphs (RoG) couples LLMs with KGs via a planning–retrieval–reasoning pipeline: the LLM first proposes KG-grounded relation paths as faithful plans, uses them to retrieve valid paths from the KG, and then performs stepwise reasoning to produce an answer.
-This yields interpretable multi-hop reasoning and reduces hallucinations by constraining reasoning to KG structure.
-ToG (Sun et al., 2024) Think-on-Graph (ToG) treats the LLM as an agent that iteratively explores the KG: it performs beam search over entities/relations, expands promising paths, and alternates retrieval with reasoning until a final path/answer is selected.
+D.2 KG-augmented Baselines
+RoG [16] Reasoning on Graphs (RoG) couples LLMs with KGs via a planning–retrieval–reasoning pipeline: the LLM first proposes KG-grounded relation paths as faithful plans, uses them to retrieve valid paths from the KG, and then performs stepwise reasoning to produce an answer.
+This yields interpretable multi-hop reasoning and reduces hallucinations by constraining reasoning to the KG structure.
+ToG [23] Think-on-Graph (ToG) treats the LLM as an agent that iteratively explores the KG: it performs beam search over entities/relations, expands promising paths, and alternates retrieval with reasoning until a final path/answer is selected.
 Compared to prompt-only baselines, ToG improves deep, compositional reasoning by explicitly navigating the KG during multi-hop search.
-ReKnoS (Wang et al., 2025c) Reasoning with Knowledge Super-relations (ReKnoS) introduces super-relations that summarize and connect multiple relational paths, enabling both forward and backward reasoning while expanding the search space the LLM can traverse.
+PoG [4] Plan-on-Graph (PoG) extends agentic KG reasoning with adaptive planning and self-correction.
+It decomposes a question into sub-objectives and iteratively performs path exploration, memory updating, and reflection to dynamically adjust reasoning trajectories.
+Through guidance, memory, and reflection mechanisms, PoG enables flexible exploration breadth and backtracking over erroneous paths, improving both reasoning accuracy and efficiency compared to fixed exploration strategies.
+ReKnoS [28] Reasoning with Knowledge Super-relations (ReKnoS) introduces super-relations that summarize and connect multiple relational paths, enabling both forward and backward reasoning while expanding the search space that the LLM can traverse.
 By reasoning over these super-relations (rather than isolated edges), ReKnoS boosts retrieval success and multi-hop accuracy, especially on complex queries.
-C.3 Cross-KG generliazation KG-specific Baselines
-KG-Specific Baselines We selected one or two recent (2022–) state-of-the-art KG-specific methods, taking the authors’ reported test/dev scores as basline comparison.
-The below outlines baseline methods reported in Sec.
-4.2 by KGQA benchmarks.
-The table 7 shows the summary.
-[…]
-Table 7: KG-specific Baselines For rows where Hits@1 is not reported, we show EM when available; where neither is reported, we show ”–”. Numbers are as reported in the cited papers/leaderboards.
-[…]
-GrailQA.
-SG-KBQA (Gao et al., 2025) is a schema-guided system that conditions a LLaMA-3.1-8B backbone on schema context and generates executable logical forms; official leaderboard reports overall Test EM/F1.
-ARG-KBQA (Tian et al., 2024) retrieves logical-form exemplars via an unsupervised ranker, then prompts GPT-3.5-0613 to generate and execute candidate logical forms.
-QALD-10-en.
-COT-SPARQL (D’Abramo et al., 2025) applies chain-of-thought prompting with entity/relation hints to produce SPARQL; it reports both standard F1 and the Macro F1-QALD.
-DFSL-MQ (D’Abramo et al., 2024) performs dynamic few-shot retrieval with multi-query generation and beam selection (evaluated with CodeLlama-70B); it reports standard F1 on QALD-10-en.
-T-REx. ATLAS (Izacard et al., 2023) is a retrieval-augmented seq2seq (FiD) reader paired with a dense retriever (Contriever); it reports Accuracyon T-Rex(treated as comparable to Hits@1).
-MultiTQ (temporal KGQA).
-Prog-TQA (pro, 2024) uses in-context KoPL program generation with linking, execution, and a self-improvement loop; results are reported for LLaMA-13B/Vicuna-13B readers.
-ARI (Chen et al., 2024) (Abstract Reasoning Induction) separates planning (knowledge-agnostic) from execution (knowledge-based) with GPT-3.5-0613; it reports Accuracy (treated as comparable to Hits@1).
-Notes: COT-SPARQL also reports Macro F1-QALD = 0.6387 on QALD-10 via GERBIL, while we list its standard F1 here for consistency across rows.
-ATLAS reports Accuracy on KILT hidden tests: zsRE = 80.8, T-REx = 85.1 (we do not map these to F1/Hits@1/EM).
-SG-KBQA numbers (EM=79.140, F1=84.403) are from the official GrailQA leaderboard (Test).
-ARG-KBQA is reported on Dev set (EM=59.6, F1=64.9).
-ARI reports Accuracy on MultiTQ (treated as comparable to Hits@1).
-Appendix D KG-R1 details
-D.1 Training algorithm
+Appendix E KG-R1 details
+E.1 Training algorithm
 The following pseudocode outlines the training procedure for KG-R1.
-Input: Dataset $D=\{(q,G,A_{q})\}$, base LLM $\pi_{0}$, horizon $H$, rollouts NN
+Input: Dataset $D=\{(q,G,A_{q})\}$, base LLM $\pi_{0}$, maximum number of turns $H$, rollouts NN
 Output: Trained policy $\pi_{\theta}$
@@
       Turn: $r^{\text{turn},(n)}_{t}=w_{\text{fmt}}v_{\text{fmt}}(r_{t})+w_{\text{kg}}v_{\text{kg}}(a_{t},o_{t+1})+w_{\text{ans}}v_{\text{ans}}(r_{T})$;
-      Global: $R^{\text{global},(n)}=w_{\text{F1}}\cdot\text{F1}(\hat{A}_{q},A_{q})+w_{\text{ret}}\cdot v_{\text{ret}}$;
+      Global: $R^{\text{global},(n)}=w_{\text{H1}}\cdot\text{H1}(\hat{A}_{q},A_{q})+w_{\text{ret}}\cdot v_{\text{ret}}$;
    end foreach
@@
 Algorithm 1 KG-R1 Training with RL
-D.2 Server Instruction Prompt
-Server Instruction template consists of A,B,C. We use the same server instruction for both training and evaluation across experiments.
+E.2 Server Instruction Prompt
+The server instruction template consists of a set of query actions and descriptions. We use the same server instructions for both training and evaluation across experiments.
 Server instruction template for KG-R1 If you encounter a KG-related error, read the error message carefully and correct your query.
@@
 -$\mathtt{get\_head\_entities}$(entity, relation) : Returns entities from which the given entity is connected by the specified relation.
-D.3 Reward weights
+E.3 Reward Weights
 Our RL objective combines per-turn and episode-level signals with fixed weights.
@@
 Episode-level reward emphasizes answer correctness and retrieval coverage: $R^{\mathrm{global}}=w_{\mathrm{F1}}\cdot\mathrm{F1}(\hat{A}_{q},A_{q})+w_{\mathrm{ret}}\cdot v_{\mathrm{ret}}$, with $v_{\mathrm{ret}}\in\{0,1\}$ indicating adequate retrieval coverage.
-Unless otherwise noted, we use the following Table 8 as defaults.
+Unless otherwise noted, we use the default values listed in Table 6.
 Table 6: Reward-component weights ($w$) for KG-R1 reward function.
@@
 Notes: Only reward-component weights are shown; optimization and rollout hyperparameters are omitted.
-D.4 RL Implementation & Hyperparameter
+E.4 RL Implementation and Hyperparameters
 Learning Rates and Optimizer.
@@
 Batch Configuration and Gradient Accumulation.
-The 3B model uses a training batch size of $128$ with validation batch size of $256$, whereas the 7B model uses a training batch size of $256$ with validation batch size of $128$.
+The 3B model uses a training batch size of $128$ with a validation batch size of $256$, whereas the 7B model uses a training batch size of $256$ with a validation batch size of $128$.
 During GRPO rollout, we collect $N=16$ trajectories per prompt for the 3B model and $N=8$ for the 7B model to balance exploration with memory constraints.
@@
 Training runs for $400$ steps with checkpoints saved every $50$ steps (3B) or every $25$ steps (7B).
-D.5 KG-R1 Retreival Server details
+E.5 KG-R1 Retreival Server Details
 Setup.
-We implemented the KG-R1 retrieval server with FastAPI (Ramírez, 2018) and Uvicorn (Encode, 2018) to support (i) a schema-free 1-hop KG query API, (ii) high-throughput async batch execution, and (iii) robust validation and observability (regex action parsing, standardized <information> wrapping with auto-closure, and final-turn safeguards).
+We implemented the KG-R1 retrieval server with FastAPI [21] and Uvicorn [10] to support (i) a schema-free 1-hop KG query API, (ii) high-throughput async batch execution, and (iii) robust validation and observability (regex action parsing, standardized <information> wrapping with auto-closure, and final-turn safeguards).
 After the KG-R1 agent generates a response, the parsed <kg-query> action is sent to the server, which performs exact string matching over the per-question knowledge graph to resolve entities and relations and returns the retrieved information.
 If the call is invalid—one of: KG_SERVER_ERROR: Invalid Action; KG_FORMAT_ERROR: Missing Required Fields; KG_FORMAT_ERROR: Wrong Argument Count; KG_SAMPLE_NOT_FOUND: Sample Missing; KG_ENTITY_NOT_FOUND: Entity Not in KG; KG_RELATION_NOT_FOUND: Invalid Relation; KG_NO_RESULTS: No Relations Found; KG_NO_RESULTS: No Entities Found—the server responds with a descriptive error message (see box below).
-KG-R1 Server Error Examples KG_SERVER_ERROR: Invalid Action <error>Action ”get_entity_info” not available (use: get_head_relations, get_tail_relations, get_head_entities, get_tail_entities)</error>  KG_FORMAT_ERROR: Missing Required Fields <error>Missing required fields for get_tail_entities: relation_name</error>  KG_FORMAT_ERROR: Wrong Argument Count <error>get_tail_relations accepts only one entity argument</error>  KG_SAMPLE_NOT_FOUND: Sample Missing <error>Sample ”sample_12345” not found in KG</error>  KG_ENTITY_NOT_FOUND: Entity Not in KG <error>Entity ”Barack Obamaa” not found in KG</error>  KG_RELATION_NOT_FOUND: Invalid Relation <error>Relation ”location.capital” not found in KG</error>  KG_NO_RESULTS: No Relations Found <error>No tail relations found for entity ”Random_Entity_123” in knowledge graph</error>  KG_NO_RESULTS: No Entities Found <error>No tail entities found for relation ”film.director.film” with head ”Barack Obama” in knowledge graph</error>
+KG-R1 Server Error Examples KG_SERVER_ERROR: Invalid Action <error>Action "get_entity_info" not available (use: get_head_relations, get_tail_relations, get_head_entities, get_tail_entities)</error>  KG_FORMAT_ERROR: Missing Required Fields <error>Missing required fields for get_tail_entities: relation_name</error>  KG_FORMAT_ERROR: Wrong Argument Count <error>get_tail_relations accepts only one entity argument</error>  KG_SAMPLE_NOT_FOUND: Sample Missing <error>Sample "sample_12345" not found in KG</error>  KG_ENTITY_NOT_FOUND: Entity Not in KG <error>Entity "Barack Obamaa" not found in KG</error>  KG_RELATION_NOT_FOUND: Invalid Relation <error>Relation "location.capital" not found in KG</error>  KG_NO_RESULTS: No Relations Found <error>No tail relations found for entity "Random_Entity_123" in knowledge graph</error>  KG_NO_RESULTS: No Entities Found <error>No tail entities found for relation "film.director.film" with head "Barack Obama" in knowledge graph</error>
 Figure 5: KG-R1 error types with actual server error messages.
-Appendix E Supplementary Experimental Results
-Figure 6 shows Training dynamics of Qwen-2.5b-it are highly consistent across three random seeds, with smooth F1 improvements and minimal variance throughout 400 optimization steps.
-Both WebQSP (red) and CWQ (blue) curves show rapid early gains and stable convergence, indicating robust optimization and reproducible policy learning.
-E.1 Reproducibility of KG-R1 Training
-Figure 6: Training dynamics of Qwen 2.5b-it across 3 random seeds demonstrate reproducibility with steady F1 improvement and low variance.
-WebQSP (red) and CWQ (blue) metrics over 400 steps show stable convergence.
-E.2 Latency and Throughput Analysis
+Appendix F Supplementary Experimental Results
+F.1 Latency and Throughput Analysis
 […]
@@
 Batched Throughput We evaluate batched inference at batch size $64$ on a single NVIDIA H100 (Table 7).
-LLM-only baselines (no KG calls) achieve high throughput—$81.8$ (Vanilla) and $70.1$ (CoT) samples/s—driven by short generation (43.0/206.0 tokens per sample).
+LLM-only baselines (no KG calls) achieve high throughput—$81.8$ (IO) and $70.1$ (CoT) samples/s—driven by short generation (43.0/206.0 tokens per sample).
 KG-R1 incurs KG-server retrieval, reducing throughput but remaining practical for offline processing: the single-run setting reaches $3.7$ samples/s ($1205.9$ gen tokens/s) with $4.4$ KG calls per query, while the N=4N{=}4 runs setting trades throughput for more KG interaction ($17.5$ calls per query), yielding $2.0$ samples/s ($612.1$ gen tokens/s).
@@
 Gen Tok./s: generated tokens per second.
-Configuration Samples Batch KG Calls Calls/Sample Total (s) Gen Tok./Sample Samples/s Gen Tok./s Vanilla Baseline 256 64 0 0.0 12.4 43.0 81.8 887.7 Vanilla CoT 256 64 0 0.0 14.1 206.0 70.1 3740.1 KG-R1 (single run) 256 64 1127 4.4 73.2 345.0 3.7 1205.9 KG-R1 (N=4N{=}4 runs) 256 64 4478 17.5 142.2 340.0 2.0 612.1
-E.3 Ablation studies
+Configuration Samples Batch KG Calls Calls/Sample Total (s) Gen Tok./Sample Samples/s Gen Tok./s IO 256 64 0 0.0 12.4 43.0 81.8 887.7 CoT 256 64 0 0.0 14.1 206.0 70.1 3740.1 KG-R1 (single run) 256 64 1127 4.4 73.2 345.0 3.7 1205.9 KG-R1 (N=4N{=}4 runs) 256 64 4478 17.5 142.2 340.0 2.0 612.1
+F.2 Ablation Studies
 All ablations in Table 3 evaluate KG-R1 with Qwen-2.5-3B-it trained on WebQSP and CWQ using a maximum turn budget of H=5H{=}5.
-We report the full ablation table in Table 3 (training curves in Figures 89).
-Turn reward. We vary the turn-level reward by setting the weights to $w_{\text{fmt}}{=}0$, $w_{\text{kg}}{=}0$, and $w_{\text{ans}}{=}0$ (default: all $0.5$).
-Turn-wise advantage. Instead of computing the turn-wise group advantage $A_{t}^{(n)}$ in Sec. 3.3, we compute a trajectory-wise group advantage
+We report the full ablation table in Table 3 and training curves in Figures 7.
+Turn Reward. We vary the turn-level reward by setting the weights to $w_{\text{fmt}}{=}0$, $w_{\text{kg}}{=}0$, and $w_{\text{ans}}{=}0$ (the default is $0.5$ for all weights).
+Turn-level Advantage. Instead of computing the turn-level group advantage $A_{t}^{(n)}$ in Sec. 3.3, we compute a trajectory-wise group advantage
 $A_{t}^{(n)}\;=\;\frac{G^{(n)}-\bar{G}}{\sigma_{G}+\epsilon},\qquad G^{(n)}\;=\;\frac{1}{T}\sum_{t}r_{t}^{\text{turn},(n)}\;+\;R^{\text{global},(n)},$
 and use it for token-level credit assignment.
-Retrieval reward. We ablate the retrieval component by setting the weight $w_{\text{ret}}{=}0$ (default: $1$).
+Retrieval Reward. We ablate the retrieval component by setting the weight $w_{\text{ret}}{=}0$ (the default is $1$).
 RL Algorithm: GRPO vs.
 PPO.
-We replace GRPO with vanilla PPO (VERL (Sheng et al., 2024) defaults), and set the turn-reward mixture weight $w_{\text{turn}}{=}0$ (default: $0.5$).
-Advantage estimation is performed by a learned value critic (Qwen-2.5-3B).
+We replace GRPO with vanilla PPO (VERL [22] defaults), and set the turn-reward mixture weight $w_{\text{turn}}{=}0$ (default: $0.5$).
+Advantage estimation is performed by a learned value critic (Qwen-2.5-3B-it).
 Hierarchical Relation Retrieval.
@@
 Base LLM Parameter Size (7B). We swap the backbone from Qwen-2.5-3B-it to Qwen-2.5-7B-it while keeping data, rewards, and budgets fixed.
-E.3.1 Training curves
-Figure 8: Training curves of ablation studies WebQSP, reporting F1 score across training steps.
-Figure 9: Training curves of ablation studies for CWQ, reporting F1 score across training steps.
-E.3.2 Hiearchical retireval format
+F.2.1 Training curves
+Figure 7: Training curves of ablation studies, reporting Hits@1 across training steps.
+F.2.2 Hierachical Retrieval Format
 Hierarchical retrieval format groups Freebase relations by their dot-notation structure (e.g., domain → type → property), presenting them as a compact tree instead of a flat comma-separated list.
-In the example(see Table E.3.2), this formatting preserves the same information while reducing verbosity—79 tokens vs.
+In the example (see Table LABEL:tab:hierarchical_relation_format), this formatting preserves the same information while reducing verbosity—79 tokens vs.
 156 tokens ($\approx$49% fewer tokens).
@@
 Flat Format Hierarchical Format
-<kg-query> get_tail_relations(”m.09c7w0”) </kg-query> <information> Tail relations for ”m.09c7w0”: location.country.first_level_divisions, location.location.containedby, location.location.contains, people.person.nationality, people.person.place_of_birth, government.government.government_for </information> Result: 156 tokens <kg-query> get_tail_relations(”m.09c7w0”) </kg-query> <information> Tail relations for ”m.09c7w0”: location country: first_level_divisions location: containedby, contains people person: nationality, place_of_birth government government: government_for </information> Result: 79 tokens (49.4% savings)
-E.3.3 PPO reward hacking
-Table 11: Policy model reward hacking under PPO. The actor model generates fake KG information that critic model cannot distinguish from real data. Blue denotes responses generated by the KG-R1 agent.
+<kg-query> get_tail_relations("m.09c7w0") </kg-query> <information> Tail relations for "m.09c7w0": location.country.first_level_divisions, location.location.containedby, location.location.contains, people.person.nationality, people.person.place_of_birth, government.government.government_for </information> Result: 156 tokens <kg-query> get_tail_relations("m.09c7w0") </kg-query> <information> Tail relations for "m.09c7w0": location country: first_level_divisions location: containedby, contains people person: nationality, place_of_birth government government: government_for </information> Result: 79 tokens (49.4% savings)
+F.2.3 PPO Reward Hacking
+The following LABEL:tab:fake_kg_information shows the PPO reward hacking examples.
+Table 9: Policy model reward hacking under PPO.
+The actor model generates fake KG information that the critic model cannot distinguish from real data.
+Blue denotes responses generated by the KG-R1 agent.
 Normal KG Retrieval Fake Information Generated by Actor Policy in PPO
 Question: Who directed the movie Inception?
-<kg-query> get_head_entities(”film.film.directed_by”, ”Inception”) </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan (m.04f9z) </information> Answer: Christopher Nolan directed Inception.
+<kg-query> get_head_entities("film.film.directed_by", "Inception") </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan (m.04f9z) </information> Answer: Christopher Nolan directed Inception.
 Question: Who directed the movie Inception?
-<kg-query> get_head_entities(”film.film.directed_by”, ”Inception”) </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan, Barack Obama, Elon Musk, Albert Einstein, Michael Jordan </information> Answer: Inception was directed by Christopher Nolan, with creative input from Barack Obama, Elon Musk, Albert Einstein, and Michael Jordan as co-directors.
-Appendix F Examples of Responses
+<kg-query> get_head_entities("film.film.directed_by", "Inception") </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan, Barack Obama, Elon Musk, Albert Einstein, Michael Jordan </information> Answer: Inception was directed by Christopher Nolan, with creative input from Barack Obama, Elon Musk, Albert Einstein, and Michael Jordan as co-directors.
+Appendix G Examples of Responses
+Table 10: Example KG-R1 response on WebQSP, showing multi-step reasoning and verification for a person-children query. Blue denotes responses generated by the KG-R1 agent.
 KG-R1 Response Example — WebQSP Question: What are Victoria Beckham’s kids’ names? Initial entities : ’Victoria Beckham’  <think>
@@
 </kg-query>
-  <information> Tail relations for entity ”Victoria Beckham”:
+  <information> Tail relations for entity "Victoria Beckham":
 base.popstra.celebrity.breakup, award.award_nominee.award_nominations, …, people.person.children, …
@@
 </kg-query>
-  <information> Tail entities for relation ”people.person.children”:
+  <information> Tail entities for relation "people.person.children":
 Brooklyn Joseph Beckham, Cruz David Beckham, Harper Seven Beckham, Romeo James Beckham
@@
 </kg-query>
-  <information> Head relations for ”Brooklyn Joseph Beckham”:
+  <information> Head relations for "Brooklyn Joseph Beckham":
 base.godparents.godparent.godchildren, people.person.children, people.sibling_relationship.sibling
@@
 </kg-query>
-  <information> Head entities for relation with tail ”Brooklyn Joseph Beckham”:
+  <information> Head entities for relation with tail "Brooklyn Joseph Beckham":
 David Beckham, Victoria Beckham
@@
    [ground_truth] {‘Brooklyn Joseph Beckham’, ‘Romeo James Beckham’, ‘Cruz David Beckham’, ‘Harper Seven Beckham’}
-Figure 10: Example KG-R1 response on WebQSP, showing multi-step reasoning and verification for a person-children query. Blue denotes responses generated by the KG-R1 agent.
+Table 11: Example KG-R1 response in CWQ. Blue denotes responses generated by the KG-R1 agent.
 KG-R1 Response Example - CWQ Question: What movie, written by Mike Jefferies, starred Nick Cannon?
-Initial entities: ”Mike Jefferies”, ”Nick Cannon”
+Initial entities: "Mike Jefferies", "Nick Cannon"
   <think>
@@
 <kg-query>
-$\mathtt{get\_tail\_relations}$(”Mike Jefferies”)
+$\mathtt{get\_tail\_relations}$("Mike Jefferies")
 </kg-query>
-  <information>Tail relations for entity ”Mike Jefferies”:
+  <information>Tail relations for entity "Mike Jefferies":
 film.film_story_contributor.film_story_credits, people.person.nationality, people.person.place_of_birth, film.writer.film, …
@@
 From the initial query results, we have two possible options for Mike Jefferies’ work.
-Let’s start with ”film.writer.film” to find the movies written by Mike Jefferies and then see if Nick Cannon starred in any of those films.
+Let’s start with "film.writer.film" to find the movies written by Mike Jefferies and then see if Nick Cannon starred in any of those films.
 </think>
 <kg-query>
-$\mathtt{get\_tail\_entities}$(”Mike Jefferies”, ”film.writer.film”)
+$\mathtt{get\_tail\_entities}$("Mike Jefferies", "film.writer.film")
 </kg-query>
-  <information>Tail entities for relation ”film.writer.film” with head ”Mike Jefferies”: Goal II: Living the Dream, Goal III: Taking on the World</information>  <think>
+  <information>Tail entities for relation "film.writer.film" with head "Mike Jefferies": Goal II: Living the Dream, Goal III: Taking on the World</information>  <think>
 Now that we have two movies written by Mike Jefferies, let’s check if Nick Cannon has starred in any of these movies. We’$l$ use the relation that connects actors with the movies they star in.
 </think> <kg-query>
-$\mathtt{get\_head\_relations}$(”Nick Cannon”)
+$\mathtt{get\_head\_relations}$("Nick Cannon")
 </kg-query>
-  <information>Head relations for entity ”Nick Cannon”: base.popstra.dated.participant, film.personal_film_appearance.person, music.album.primary_release,
+  <information>Head relations for entity "Nick Cannon": base.popstra.dated.participant, film.personal_film_appearance.person, music.album.primary_release,
 base.popstra.lived_with.participant, film.film.written_by, …
@@
   <think>
-From the relations where ”Goal II: Living the Dream” is involved, we can see that it is a film with a writer, which aligns with Mike Jefferies.
+From the relations where "Goal II: Living the Dream" is involved, we can see that it is a film with a writer, which aligns with Mike Jefferies.
 </think>
@@
    [ground_truth] {’Goal II: Living the Dream’ }
-Figure 11: Example KG-R1 response in CWQ. Blue denotes responses generated by the KG-R1 agent.
-KG-R1 Response Example - SimpleQA Question: where is the madam satan located? (Initial entities: ”Madam Satan”)
+Table 12: Example KG-R1 response in SimpleQA. Blue denotes responses generated by the KG-R1 agent.
+KG-R1 Response Example - SimpleQA Question: Where is the madam satan located? (Initial entities: "Madam Satan")
   <think>
@@
 </kg-query>
-  <information>Tail relations for entity ”Madam Satan”:
+  <information>Tail relations for entity "Madam Satan":
 film.film.cinematography, common.topic.notable_types, film.film.language, film.film.starring,film.film.country, … award.ranked_item.appears_in_ranked_lists, film.film.film_art_direction_by
@@
 </kg-query>
-  <information>Tail entities for relation ”film.film.country” with head ”Madam Satan”: United States of America</information>  <think>
+  <information>Tail entities for relation "film.film.country" with head "Madam Satan": United States of America</information>  <think>
 I have found that Madam Satan, identified as a film, is located in the United States of America.
@@
 </kg-query>
-  <information>Tail entities for relation ”film.film.runtime” with head ”Madam Satan”: m.03kx9f2</information>  <think>
+  <information>Tail entities for relation "film.film.runtime" with head "Madam Satan": m.03kx9f2</information>  <think>
 I am still not obtaining the most helpful information about the location. Possibly Madam Satan was released, but not filmed in a specific place.
@@
 </kg-query>
-  <information>Head relations for entity ”m.02tb01n”:
+  <information>Head relations for entity "m.02tb01n":
 film.film.starring, film.actor.film
@@
   <think>
-I have questionned the relation ”film.actor.film”. This relation includes the movies the actor is in. Let’s try to find a clearer detail for the location of Madam Satan.
+I have questioned the relation "film.actor.film". This relation includes the movies the actor is in. Let’s try to find a clearer detail for the location of Madam Satan.
 </think>
@@
    [ground_truth] {’United States of America’ }
-Figure 12: Example KG-R1 response in SimpleQA. Blue denotes responses generated by the KG-R1 agent.
-KG-R1 Response Example — T-REx Question: What is the instance of of Solberg??
-Initial entities: ”Solberg”
+Table 13: Example KG-R1 response in T-REx. Blue denotes responses generated by the KG-R1 agent.
+KG-R1 Response Example — T-REx Question: What is the instance of Solberg??
+Initial entities: "Solberg"
   <think>
@@
 </kg-query>
-  <information>Tail relations for entity ”Solberg”: P1889, P407, P1552, P282, P31, P460</information>  <think>
+  <information>Tail relations for entity "Solberg": P1889, P407, P1552, P282, P31, P460</information>  <think>
 The relations provided show how Solberg is related to other entities. Let’s explore each relation to find out more about Solberg.
@@
 </kg-query>
-  <information>Tail entities for relation ”P282” with head ”Solberg”: Latin script</information>  <think>
-The relation P282 corresponds to ”type.object.name” according to the API documentation, indicating that Solberg is a human name.
+  <information>Tail entities for relation "P282" with head "Solberg": Latin script</information>  <think>
+The relation P282 corresponds to "type.object.name" according to the API documentation, indicating that Solberg is a human name.
 The relation P407 is also noteworthy but does not give further specific information.
@@
 </kg-query>
-  <information>Tail entities for relation ”P31” with head ”Solberg”: Wikimedia disambiguation page, family name</information>
+  <information>Tail entities for relation "P31" with head "Solberg": Wikimedia disambiguation page, family name</information>
   <think>
-The retrieval of ”Soelberg”, ”Sohlberg”, and ”Sohlberg” suggests that Solberg might be associated with these names. To find an instance of Solberg, we can search for a specific entity of this name.
+The retrieval of "Soelberg", "Sohlberg", and "Sohlberg" suggests that Solberg might be associated with these names. To find an instance of Solberg, we can search for a specific entity of this name.
 </think>
@@
    [ground_truth] {’family name’, ’human settlement’, ’urban area in Sweden’, ’minor locality in Sweden’, ’minor locality in Sweden’, ’Wikimedia disambiguation page’, ’human settlement’, ’Ortsteil’, ’stone setting’, ’human settlement’, ’minor locality in Sweden’, ’hill’, ’stone setting’, ’hill’, ’human settlement’ }
-Figure 13: Example KG-R1 response in T-REx. Blue denotes responses generated by the KG-R1 agent.
-KG-R1 Response Example - QALD10en Question: After whom is the Riemannian geometry named?? (Initial entities: ”Riemannian geometry”)
+Table 14: Example KG-R1 response in QALD10en. Blue denotes responses generated by the KG-R1 agent.
+KG-R1 Response Example - QALD10en Question: After whom is the Riemannian geometry named?? (Initial entities: "Riemannian geometry")
   <think>
-I need to find out who the Riemannian geometry is named after. Let’s start by investigating the relations and entities connected to ”Riemannian geometry”.
+I need to find out who the Riemannian geometry is named after. Let’s start by investigating the relations and entities connected to "Riemannian geometry".
 </think>
@@
 </kg-query>
-  <information>Tail relations for entity ”Riemannian geometry”:
+  <information>Tail relations for entity "Riemannian geometry":
 on focus list of Wikimedia project, named after, maintained by WikiProject, topic’s main category, instance of, subclass of
@@
 </kg-query>
-  <information>Tail entities for relation ”named after” with head ”Riemannian geometry”: Bernhard Riemann</information>  <think>
+  <information>Tail entities for relation "named after" with head "Riemannian geometry": Bernhard Riemann</information>  <think>
 I now have the information that the Riemannian geometry is named after Bernhard Riemann. Let’s confirm this by checking if there’s more detailed information about this relationship.
@@
 </kg-query>
-  <information>Head relations for entity ”Bernhard Riemann”:
+  <information>Head relations for entity "Bernhard Riemann":
 named after
@@
 </kg-query>
-  <information>Head entities for relation ”named after” with tail ”Bernhard Riemann”: Riemannian geometry</information>  <think>
+  <information>Head entities for relation "named after" with tail "Bernhard Riemann": Riemannian geometry</information>  <think>
 This confirms my findings; the Riemannian geometry is named after Bernhard Riemann.
@@
    [ground_truth] {’Bernhard Riemann’}
-Figure 14: Example KG-R1 response in QALD10en. Blue denotes responses generated by the KG-R1 agent.
+Table 15: Example KG-R1 response in GrailQA. Blue denotes responses generated by the KG-R1 agent.
 KG-R1 Response Example — GrailQA Question: What video game is published by softdisk?
-Initial entities: ”Near West Side”
+Initial entities: "Near West Side"
   <think>
-First, I need to find out what game is published by Softdisk. Without specific knowledge about ”Near West Side”, I will start with the game publisher for Softdisk.
+First, I need to find out what game is published by Softdisk. Without specific knowledge about "Near West Side", I will start with the game publisher for Softdisk.
 </think>
@@
 </kg-query>
-  <information>Head relations for entity ”Near West Side”:
+  <information>Head relations for entity "Near West Side":
 book.periodical_publisher_period.publisher, business.industry.companies, business.employment_tenure.company
@@
   <think>
-The head relation ”business.employment_tenure.company” suggests that ”Near West Side” might be a game company or related to a game.
-I will try to find out the publisher of a game related to ”Near West Side” by looking for games published by Softdisk.
+The head relation "business.employment_tenure.company" suggests that "Near West Side" might be a game company or related to a game.
+I will try to find out the publisher of a game related to "Near West Side" by looking for games published by Softdisk.
 </think>
@@
 </kg-query>
-  <information>Tail relations for entity ”Near West Side”:
+  <information>Tail relations for entity "Near West Side":
 organization.organization.place_founded, organization.organization.headquarters, book.periodical_publisher.periodicals_published, business.employer.employees, cvg.cvg_publisher.games_published, business.business_operation.industry, common.topic.notable_types
@@
   <think>
-The previous query did not provide a conclusive answer. Let’s try a different approach by checking if ”Augusto Cury” is associated with any game publishers.
+The previous query did not provide a conclusive answer. Let’s try a different approach by checking if "Augusto Cury" is associated with any game publishers.
 </think>
@@
 </answer>
-   [ground_truth] {’Rescue Rover’,’Catacomb 3-D’,’Shadow Knights’,”Dangerous Dave’s Risky Rescue”,’Tiles of the Dragon’, ’Catacomb’, ’Commander Keen: Keen Dreams’, ’Dave Goes Nutz!’,’Chagunitzu’,’Dangerous Dave in the Haunted Mansion’,’Hovertank 3D’,’Iron Seed’}
-Figure 15: Example KG-R1 response in GrailQA. Blue denotes responses generated by the KG-R1 agent.
-KG-R1 Response Example — MultiTQ Question: Who negotiated with Malaysia in 2005–10?  <think>
-To answer this, I should check which relations connect to Malaysia that indicate negotiations, then look for counterpart entities within the 2005–10 range.
-</think>
-<kg-query>
-$\mathtt{get\_head\_relations}(\text{"Malaysia"})$
-</kg-query>
-  <information> Head relations for entity ”Malaysia”:
-Appeal for aid; Appeal for diplomatic cooperation; Appeal for humanitarian aid; … Engage in negotiation; Express intent to settle dispute
-</information>
-  <think>
-”Engage in negotiation” appears among head relations. Next, retrieve the counterpart entities that negotiated with Malaysia.
-</think>
-<kg-query>
-$\mathtt{get\_head\_entities}(\text{"Malaysia"},\text{"Engage in negotiation"})$
-</kg-query>
-  <information> Head entities connected by (?, Engage in negotiation, Malaysia):
-Cambodia [2008-06]; China [2011-11]; Governor (Malaysia) [2007-05]; Iran [2006-05; 2006-07]; Japan [2005-05; 2005-06; 2009-01; 2011-10]; Julia Gillard [2011-07]; Mwai Kibaki [2007-04]; … Yemen [2005-02; 2008-10]
-</information>
-  <think>
-We need the counterpart who negotiated with Malaysia specifically within 2005–10. Signals suggest Thailand-related negotiations in this window; inspect Thailand’s negotiation partners.
-</think>
-<kg-query>
-$\mathtt{get\_tail\_relations}(\text{"Thailand"})$
-</kg-query>
-  <information> Tail relations for entity ”Thailand”:
-Accede to demands for rights; Accuse; Appeal for aid; Appeal for de-escalation of military engagement; … Engage in negotiation
-</information>
-  <think>
-Confirm the negotiation counterparts of Thailand to identify the specific actor matching the 2005–10 window.
-</think>
-<kg-query>
-$\mathtt{get\_tail\_entities}(\text{"Thailand"},\text{"Engage in negotiation"})$
-</kg-query>
-  <information> Tail entities connected by (Thailand, Engage in negotiation, ?):
-Abhisit Vejjajiva [2009-06; 2009-10; 2010-01]; Cambodia [2005-04; 2005-06; 2008-01; …; 2012-07]; China [2006-06; 2006-09; 2010-03; 2011-01]; Donald Rumsfeld [2005-06]; Koichiro Matsuura [2011-02]; … South Korea [2005-12]
-</information>
-  <think>
-Within 2005–10, the counterpart that negotiated with Malaysia is Surakiart Sathirathai.
-</think>
-<answer>
-Surakiart Sathirathai
-</answer>
-   [ground_truth] {’Surakiart Sathirathai’}
-Figure 16: Example KG-R1 response on MultiTQ, illustrating all four query functions and their usage. Blue denotes responses generated by the KG-R1 agent.
+   [ground_truth] {’Rescue Rover’,’Catacomb 3-D’,’Shadow Knights’,"Dangerous Dave’s Risky Rescue",’Tiles of the Dragon’, ’Catacomb’, ’Commander Keen: Keen Dreams’, ’Dave Goes Nutz!’,’Chagunitzu’,’Dangerous Dave in the Haunted Mansion’,’Hovertank 3D’,’Iron Seed’}
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The abstract and introduction clearly state the paper’s main claims: improving KG-RAG effectiveness, efficiency, and transferability through reinforcement learning over structured KG interactions.
+These claims are supported by experimental results in Section˜4.34.5.
+  5. Guidelines:
+    • The answer [N/A] means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A [No] or [N/A] answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: We discuss the limitations of KG-R1 in Appendix A, including its dependence on knowledge graph quality, potential degradation under schema mismatch, and the computational overhead of reinforcement learning training.
+  10. Guidelines:
+    • The answer [N/A] means that the paper has no limitation while the answer [No] means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate “Limitations” section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14.
+Justification: Yes.
+We provide formal propositions and proofs for the completeness, finite-horizon realizability, schema-free transferability, and minimality of the KG action interface in Appendix B.
+The assumptions and operator definitions are explicitly stated in the preliminaries.
+  15. Guidelines:
+    • The answer [N/A] means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We provide full methodological details, including action design, reward formulation, training procedures, hyperparameters, and evaluation protocols in the main paper and appendix.
+Code is released anonymously for reproducibility.
+  20. Guidelines:
+    • The answer [N/A] means that the paper does not include experiments.
+    • If the paper includes experiments, a [No] answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We release anonymized code, training scripts, and evaluation pipelines (link provided in the abstract). All benchmark datasets used in this work are publicly available.
+  25. Guidelines:
+    • The answer [N/A] means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so [No] is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We specify the benchmark datasets, model backbones, training horizon, reward settings, RL optimization configurations, and evaluation metrics in Section˜4 and the appendix.
+  30. Guidelines:
+    • The answer [N/A] means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We report mean and standard deviation across five runs in all main experimental tables and ablation studies. The reported error bars reflect run-to-run variation.
+  35. Guidelines:
+    • The answer [N/A] means that the paper does not include experiments.
+    • The authors should answer [Yes] if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
+    • If error bars are reported in tables or plots, the authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39.
+Justification: We provide hardware specifications, memory configurations, precision settings, and training duration details in the appendix, including GPU type and model-specific resource allocation.
+  40. Guidelines:
+    • The answer [N/A] means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44.
+Justification: We have reviewed the NeurIPS Code of Ethics and confirm that this work complies with its guidelines.
+The experiments use publicly available datasets and models without involving human subjects or sensitive personal data.
+  45. Guidelines:
+    • The answer [N/A] means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer [No], they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49.
+Justification: We discuss both positive societal impacts and potential negative impacts in Appendix A, including improved factual reasoning and risks of propagating incorrect or manipulated knowledge.
+  50. Guidelines:
+    • The answer [N/A] means that there is no societal impact of the work performed.
+    • If the authors answer [N/A] or [No], they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate Deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: This work does not release high-risk models or sensitive datasets. The proposed method builds upon existing open-weight language models and public benchmark datasets.
+  55. Guidelines:
+    • The answer [N/A] means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All datasets, models, and baseline implementations used in this work are properly cited, and their original licenses and usage terms are respected.
+  60. Guidelines:
+    • The answer [N/A] means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We release the implementation of KG-R1 with documentation and scripts for reproducing the main experimental results.
+  65. Guidelines:
+    • The answer [N/A] means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This work does not involve crowdsourcing or research with human subjects.
+  70. Guidelines:
+    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: This work does not involve human subjects and does not require IRB approval.
+  75. Guidelines:
+    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigor, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79.
+Justification: Large language models are a core component of the proposed KG-R1 framework and are fully described in the methodology section.
+Additionally, we used an LLM-based coding assistant to support parts of the implementation.
+All generated code was reviewed and validated by the authors, and all experiments were designed, executed, and verified by the authors.
+  80. Guidelines:
+    • The answer [N/A] means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy in the NeurIPS handbook for what should or should not be described.